MDI Guide

IDL Calibration and Analysis Software

Ver 1.0 M.Morrison 9-Jun-94

This Copy Printed: June 9, 1994

MDI-340025 SOI Tech Note#: 117 Some suggested reference books, material, and methods of finding more information:

• “MDI Guide: An Overview of the /mdisw System”

• TBD - what is useful to list here????

• The official MDI IDL software should have a detailed description of the routine in the header of the source file. It is possible to access this header from within IDL by using DOC LIBRARY. A sample call is: IDL> doc library,’mdipict’

If you have general comments, suggestions, or questions regarding this document or software, please send them to:

morrison@quake.stanford.edu

Contents

1 Introduction 1 1.1 Using the /mdisw System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 This Document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Telemetry Filenames, Datasets, and MicroVAX Files . . . . . . . . . . . . . . . . 1 1.4 The FITS Header for MDI Images . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 General Use Routines 5

2.1 How to View MDI Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.1 VIEW IMAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.2 SHOW DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.1.3 Access to PostScript Summary Atlas . . . . . . . . . . . . . . . . . . . . . 7

2.2 How to Read MDI Data (READ MDI) . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3 How to Make Plots versus Time (UTPLOT) . . . . . . . . . . . . . . . . . . . . . 8 2.4 How to Get General Summary Information of an Image (IMG SUMMARY) … 9 2.5 How to Direct Which Printer is Used for Hardcopies (SET PRINTER) . . . . . . 9 2.6 Manipulation of the FITS Headers . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.6.1 Extracting a Single Field (SXPAR) . . . . . . . . . . . . . . . . . . . . . . 10 2.6.2 Adding a New Field to an Existing Header (SXADDPAR) . . . . . . . . . 10 2.6.3 Building a Basic FITS Header (MK FITS HEAD) . . . . . . . . . . . . . 10 2.6.4 Reading a FITS File (RFITS2) . . . . . . . . . . . . . . . . . . . . . . . . 10 2.6.5 Writing an Image to a FITS File (WRT FITS) . . . . . . . . . . . . . . . 11

3 LFFT Routines 12 3.1 MDIPICT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.2 MDIFOCUS2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.3 LTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

3.3.1 Details on the Light Transfer Calculation . . . . . . . . . . . . . . . . . . 14 3.4 MDIISS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.5 MDILASER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.6 MDIDUST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.7 MDIMROT, MDIPROT, and MDIPDIST . . . . . . . . . . . . . . . . . . . . . . 19 3.8 IPCOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.9 SHUTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.10 TRAP CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

4 House Keeping Routines 31 4.1 How to Make Temperature and Voltage HK Plots (HKPLOT) . . . . . . . . . . . 31 4.2 GET HK INFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

5 Calibration Routines 34 5.1 Michelson Tuning Analysis (DO DISP MMAP1) . . . . . . . . . . . . . . . . . . 34 6 Command History Routines 36

6.1 HIST SEA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 6.2 HIST SUMMARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 6.3 HIST COUNT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

A Description of the Purpose of Several Control Files 38 A.1 MDION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 A.2 MDICOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 A.3 MDIPICT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 A.4 MDILTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 A.5 MDIFOCUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 A.6 MDIISS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 A.7 MDIDSOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 A.8 MDIMROT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 A.9 MDIPROT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 A.10 MDIPDIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 A.11 MDIDUST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 A.12 MDIKAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 A.13 SHUTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 A.14 MDILASER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 A.15 DMCDOOR1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

B Listing of MDI Data Product Codes (DPC) 44

C Listing of House Keeping Mnemonics 45

D Listing of FITS Fields 54

E Listing of How Many Times Some Control Files were Run (HIST COUNT output) 70

F Listing of When Control Files were Run (HIST SUMMARY output) 71

G Chronological Listing of When Control Files were Run (CAL LOG output) 87

List of Figures

1 Sample Output for MDIPICT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2 Sample Output for MDIFOCUS2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3 Sample Output for LTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4 Sample Output for MDIISS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5 Sample Output for MDILASER . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6 Sample Output for MDIDUST . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7 MDIMROT Sample - Raw Grid Positions . . . . . . . . . . . . . . . . . . . . . . 22 8 MDIMROT Sample - Positions Smoothed . . . . . . . . . . . . . . . . . . . . . . 22 9 MDIMROT Sample - Remove Stimulus Drift . . . . . . . . . . . . . . . . . . . . 22 10 MDIMROT Sample - Michelson 1 Position Shifts for Tunings . . . . . . . . . . . 24 11 MDIPROT Sample - Position Shifts for Each Polarization . . . . . . . . . . . . . 25 12 MDIPDIST Sample - Position Shift for Polarization . . . . . . . . . . . . . . . . 26 13 MDIPDIST Sample - Plate Scale Size Change for each Polarization . . . . . . . . 27 14 Sample Output for IPCOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 15 Sample Output for TRAP CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . 30

Preface

The conventions used in this manual are:

• Bold face for examples of what the user should type

• Program names are capitalized when mentioned in the text, but are not capitalized in the user input examples

• Variable names are italicized when mentioned in the text

• Dates and times are in Universal Time (UT) and have the convention “dd-mmm-yy”

• There are several different types of prompts that appear in the examples. The prompt IDL> is used for when the user has already entered and is running IDL. The prompt % is used to signify Unix commands, and the prompt $ signifies VMS commands. There are also FTP> prompts for ‘File Transfer Protocol.’

For example, you would only type .run show data for the following description. You would not type IDL>. IDL> .run show data

1 Introduction

1.1 Using the /mdisw System

In order for you to use the IDL routines and some of the other capabilities described here, you need to be using the /mdisw setup. A full description of this system is described in the document ”An Overview of the /mdisw System”. Basically all that you need to do is to add the following line to your /.cshrc file

          source /mdisw/setup/mdi_setup

1.2 This Document

A PostScript copy of this document can be found in the file:

          /mdisw/doc/guide/mdisw_idl.ps

If you have an X-terminal, you can view this document using either gs to review the PostScript file, or xdvi to review the DVI file. The DVI previewer will not show the plots or images. Aliases have been set up to make this easy for you. The commands are:

% gs idl % xdvi idl

1.3 Telemetry Filenames, Datasets, and MicroVAX Files

The raw telemetry files at LPARL are typically located in one of the following directories:

        /md60/log/sci160k
        /md65/log/sci160k
        /md66/log/sci160k
        /md54/optest

They have names of the form YYMMDD HHMMSS EE (e.g. 940420 193742 01.tfr) where “YYMMDD HHMMSS” is the year, month, day, hour, minutes seconds when the telemetry file was started. “EE” is the extension since the files are closed periodically to avoid having very large files. It starts with “01” and increments up. The extension “.tfr” is for the raw telemtry file. The “.tfd” is for a directory of what is in the file, and it is created as the raw telemetry file is created. The “.tfrx” file is made whenever the RDTFR IDL routine is used, and it is a roadmap of which data products are in the telemetry files and what the starting record number is for each data product. This allows for quick direct reading reading of the telemetry file.

Links are used to access the raw telemetry files at Stanford. You can use the command:

% ls -l ∼soi/data/ | grep tlm

to find out which telemetry files are on-line. An example of what is returned is:

lrwxrwxr-x  1 mark           33 Apr 23 21:39 tlm940423_052704 -> /scr10/mdi_data/tlm/940423_052704
lrwxrwxr-x  1 mark           33 Apr 24 22:48 tlm940423_065147 -> /scr16/mdi_data/tlm/940423_065147

The raw telemetry files are converted to FITS files at Stanford. The following convention is used:

Prefix Start  End  Start  End Description
Series Series  Date  Date
OPT 013  438  Nov-92 2-Oct-93 Optical Performance Tests
OCT 500  617 19-Oct-93 7-Nov-93 Optical Calibration Tests
NOV 618a  766 25-Nov-93 29-Mar-94 November Oven Vacuum
APR 800  1200 1-Apr-94 26-Apr-94 April Test
CAT 000  426 29-Jun-93 16-Mar-94 Calibration and Alignment Tests

The following files contain information on which telemetry files were used for each series.

-rw-rw-r--  1 rick         1813 Apr 13 16:18 /home/soi/data/aprlog
-rw-rw-r--  1 rick        28246 Apr 18 13:46 /home/soi/data/catlog
-rw-rw-r--  1 rick        14469 Apr 12 18:13 /home/soi/data/novlog
-rw-rw-r--  1 rick        10397 Apr  2 15:52 /home/soi/data/octlog
-rw-rw-r--  1 rick        29471 Mar 30 14:32 /home/soi/data/optlog
-rw-rw-r--  1 rick        43284 Apr 23 17:54 /home/soi/data/tlglog

A cross reference listing of the different datasets which are available is shown in the Appendix. This listing shows which telemetry file is used for which datasets, the number of images in the series, the date and time of the first image, the number of minutes it took to complete the series, the control file which was run, whether the telemetry and FITS files are on-line at Stanford, the backup tape and file numbers, and the CCD and oven temperature at the beginning of the series.

1.4 The FITS Header for MDI Images

The following is a sample FITS header for an MDI image.

SIMPLE  =                    T / file conforms with FITS standard; SOI v 0.7
BITPIX  =                   16
NAXIS   =                    2
NAXIS1  =                  256
NAXIS2  =                  256
DATE-OBS= '29/11/93'
TIME-OBS= '02:52:23'
IMAGE_T = '1993:11:29:02:52:23'
CONTROL = '   0   1   0   0   0'
DEPSCR_0= '   0   0'
DEPSCR_1= '   0   0'
DEPSCR_2= '   0   0'
DEPSCR_3= '   0   0'
DEPSCR_4= '   0   0'
DEPSCR_5= '   0   0'
DEPSCR_6= '   0   0'
DEPSCR_7= '   0   0'
DEPSCR_8= '   0   0'
DEPSCR_9= '   0   0'
DEPSCR_A= '   0   0'
DEPSCR_B= '   0   0'
DEPSCR_C= '   0   0'
DEPSCR_D= '   0   0'
DEPSCR_E= '   0   0'
DEPSCR_F= '   0   0'
SQID    =                65536
SQNUM   =                    0
SQFRM   =                    0
RSTRT   =                    0
FRMNUM  =                   29
HKPACK  =                23043
DEPVR   = '   0  17'
OCTVR   = '   0   0'
SEQVR   = '   0   2'
IPVR    = '   0   0'
QUEUEVR = '   0   0'
ALIGN   =                    0
C1POS   =                    0
C2POS   =                    0

MISCSTAT=                    0
LTERRS  = '   0   0   0   0'
LTSTAT  =                    0
HTRSTAT = '   0   0   0   0'
OVNTEMP = '   0   0'
OPTEMPS = ' 134 136'
CCDTMP_0= '   0 255 254   0   0 255 255   0   0 254'
CCDTMP_1= ' 255   0   0 254 255   0   0 255 255   0'
MKM1POS =                    0
MKM2POS =                    0
MKPAPOS =                    0
MKSHPATH= 'DARK'
END

There is a list of the FITS fields and their meanings in the Appendix on page 54.

2 General Use Routines

2.1 How to View MDI Data

There are a few different methods to look at MDI data. SHOW DATA and VIEW DATA are IDL routines which you can run interactively, but which require that the telemetry or FITS files be available. A Postscript Previewer can access a database of PostScript files and summary pages for all MDI calibration data is available.

2.1.1 VIEW IMAGE

The routine VIEW IMAGE prompts the user for a telemetry file name or dataset name, and which images to extract. When prompted for the file name, it is possible to answer “MENU” and to be prompted with the telemetry files or datasets which are available. A sample session is:

IDL> .run view image

The first question asked is if you wish to read some images from a file. If you have already read in the images that you are interested in using VIEW IMAGE, then you can answer no to this question. If you simply hit < CR >, it will use the answer YES.

Do you wish to read some images from a file?   [Default: YES ] ?
Do you wish to read a raw telemetry file (other option is FITS)  [Default: NO ] ?
For telemetry files, do not include the extension.   Some examples are:
      940311_232118
      /md66/log/sci160k/940311_232118
Enter MENU to display a menu of the available telemetry files:
Enter the filename (QUIT to abort)  [Default:  ]

The second question is whether it is a raw telemetry file. If you simply hit < CR > it will use the Stanford FITS datasets. The third question is the file name. If you know the file name, you can enter it. If you do not, you can enter MENU and a menu of the telemetry files or datasets which are on-line will appear. Click on the dataset that you wish to read. If you have already read a image from a file, and wish to read a different image, it is simple since the default input file name is the last file which was read (ie: you just have to hit < CR >).

Next you need to specify the images that you want to read. You are asked to specify the starting image number and the ending image number of the images to read.

Number of images available:       43
Enter the starting image to extract (starting at 0)  [Default: 0 ]
Enter the ending image to extract                    [Default: 42 ] 1

In the example above, it will read the first two images in the file. The image and the FITS header and in the variables img and head. These variables are still available even after you finish running VIEW IMAGE. The image minimum, maximum, average and standard deviation are printed to the screen. The following message is also printed.

The image cube is stored in the variable "img"
The header information is stored in the variable "head"
You can run stepper any time using the command "stepper, img, xsiz=512"
Will rebin to XSIZE=512, YSIZE=512
-------- STEPPER Options --------
There are 2 images in the array
There are 2 images selected
Enter "b" to step backwards                 "h" to call HARDCOPY
      "s" to select new start index         "z" for ZOOM
      "c" to call LOADCT                    "x" for XLOADCT
      "p" to call the IDL PROFILES routine  "l" to call light curve routine
      "g" to overlay SXT solar grid         "o" for observing region location
      "q" to quit                           "m" for movie mode
      "?" to display this help menu         "anything else" to step forward

VIEW IMAGE then then calls the Yohkoh routine STEPPER which is a general purpose routine to display an image. Some of the above options are not available for non-SXT images. The most simple options to step forward or backward are performed by pressing the space bar or “b”. A useful method for stretching an image to adjust the contrast is to use XLOADCT (the “x” option). Line and column traces can be accessed using the IDL PROFILE routine using the “p” option. “z” will allow you to zoom in on a feature. “q” will exit STEPPER.

2.1.2 SHOW DATA

SHOW DATA is an older version of a simple image reading and displaying routine. It is recommended to use VIEW IMAGE if possible.

The routine SHOW DATA is a wrapper around some basic reading and displaying routines that allows a user to display any data in the 160k tlm files and also to choose autoscaling or specifically fix the scaling. Again, the two routines (sepa- rated to have the reader and displayer de-coupled) are in one file “showdata.pro”. A sample run is:

IDL> .run showdata IDL> readata

Simply answer the questions:

1. input the filename (use the extension to read that specific file, or omit it to read the whole sequence)

2. input the first image to be read 3. input the last image to be read (=first if you want to read only one image altogether)

NOTE don’t use this method to read and display one image at a time from the same sequence: VERY inefficient; read the FULL RANGE you want. It’s the display program below that displays them one at a time...

IDL> show ;for autoscaling, convenient for first time IDL> show, lo=min, hi=max ;to set min and max IDL> show, lo=1900, hi=2500 ;to set scaling from 1900 to 2500

NOTE If you omit lo it will be set to 0, if you omit hi, it will be set to 4095. Omitting both, of course puts you to autoscaling again.

You will be told how many images You have to your disposition (if you wanted to read 5 to 10, you’ll have all 10, cause the program needs to read them all any way - except if you had specified ONLY 10, then you will have only that one) Simply answer the questions:

- which do you want to display (enter a number, or 0 for all,
  or Q to quit)

It will loop in this mode to let you review the images ’all-you-can-eat’ Type Q (case insensitive) to quit show.

You can re-run show again with or without scaling as many times as You wish. To read another set of data, invoke READATA again.

The questions are reasonably fool-proof in terms of logic, but NOT in terms of type conversion error, so type a true integer number where asked for it. If you mistype and IDL ”crashes” due to Your type error, just type:

IDL> retall

and start again (readata or show).

For the IDL initiated, the data is passed via the common statement, thus, if You want the images to be to your disposition at the MAIN level, just declare the common at the main level (exactly as is):

IDL> common pictures,a

Now You have the images in a=intarr(xdim,ydim,numberofimages) to play with.

2.1.3 Access to PostScript Summary Atlas

A set of summary PostScript pages are available for each of the datasets which were taken on the ground. These files are stored at Stanford in the directory ∼soi/data/atlas. There is a subdirectory for each of apr, nov, oct, opt and cat. A sample of how to display them to the screen with the PostScript previewer is:

% gs ∼soi/data/atlas/apr/apr999.ps

2.2 How to Read MDI Data (READ MDI)

The routine READ MDI will read the raw telemetry data files, the FITS data files, or the “F0” MicroVAX data files. The routine figures out which type of file it is based on the file name. The outputs are the image array plus a FITS header array. Some sample calls are:

IDL> read mdi, infil, images, array, header IDL> read mdi, ’/md66/log/sci160k/940401 143902’, 3, img, h IDL> read mdi, 766, [3,6,12,18], img, h IDL> infil = ’/md66/log/sci160k/940401 143902’ IDL> read mdi, infil, indgen(10), img, h IDL> read mdi, infil, indgen(10)+10, img, h

The inputs are infil which is the input file or series number and images which is a list of the images to be read starting with 0. The output is array which is a 2-D or 3-D array with the images you requested, and header which is a string FITS header of dimensions NxM where you requested M images (N is the number of lines of text in the FITS header for a single image).

In the second example, the fourth image of “940401 143902” is read. In the last two examples, the first and then the second 10 images in the file are read.

The routine has several optional keywords. If you wish to find out how many images are in a data series or a telemetry file, you use a command of the following format:

IDL> read mdi, infil, xxxx, nimages, /nimages

The result is returned in the output variable nimages.

2.3 How to Make Plots versus Time (UTPLOT)

UTPLOT plots any quantity on a time plot with hours, minutes, and seconds. It will automatically determine the x-axis labels (whether to mark ticks by HH:MM:SS or by DD-MMM-YY). The input to UTPLOT can either be a struture with .TIME and .DAY (where .TIME is milliseconds of day, and .DAY is the number of days since 1-Jan-79) or you can pass a floating point vector which has the number of seconds past a reference time. Some examples are shown below:

IDL> utplot, x, y, ref time, xrange=xrange 2.4 How to Get General Summary Information of an Image (IMG SUMMARY) 9

IDL> utplot, index, y, timerange=[index(0), index(100)] IDL> utplot, x, y, ref time, timerange=[’1-jun-91’,’5-jun-91’] IDL> utplot, x, y, ref time, psym=10

Most typical IDL plotting parameters also apply. You can overplot other quantities using OUT-PLOT. For example,

IDL> utplot, x, y, ref time IDL> outplot, x2, y2

2.4 How to Get General Summary Information of an Image (IMG SUMMARY)

It is possible to get a summary display of a single image by using the routine IMG SUMMARY. The summary page includes:

• The image displayed in grey scale for the full range and for ±1 sigma around the average.

• Histogram plots in log and linear scale for the full image and also for ±2 sigma around the average (±1 sigma if the standard deviation is large)

• Single line and column plots, and plots of the average of all lines and all columns

• The image average, standard deviation, minimum, maximum, and range for both the whole image and for the image excluding the first and last two columns and lines.

• OPTIONALLY display some comment information

• OPTIONALLY display the location and amplitude of the N brightest pixels (with an additional option to exclude excluding an M pixel margin around the edge of the image).

Some sample calls to IMG SUMMARY are:

IDL> img summary, img IDL> img summary, img, comments=[’Temperature = 22.2’, ’Pressure = 1e-5’] IDL> img summary, img, tit=tit, /hc IDL> img summary, img, brightest=8 IDL> img summary, img, brightest=8, margin=2

A sample output can be seen in Figure 1.

2.5 How to Direct Which Printer is Used for Hardcopies (SET PRINTER)

The routine SET PRINTER prompts the user with a menu to allow the user to interactively define which printer should be used for hardcopy output. The routine PPRINT should be used by all IDL routines to spawn the print commands. The following environment variables are used: $PRINTER for the printer queue name, $PRINT CMD for the explicit printer command (if defined, PPRINT actually ignores $PRINTER), and $PPRINT NODE for the computer system to use. If $PPRINT NODE is used, the user must have a ∼/.rhosts file which allows “rsh” commands into that account. A sample call is:

IDL> set printer

2.6 Manipulation of the FITS Headers

2.6.1 Extracting a Single Field (SXPAR)

To extract the information for a given field from a FITS header, you can use the routine SX-PAR. The FITS header is a string array. The routine will figure which string holds the field and what it’s data type is (it will return integer, floating or string types). Sample calls are:

IDL> xpos = sxpar(h, ’LASPOS X’) IDL> m1tune = fix(strmid(sxpar(header0, ’DEPSCR D’), 0, 4))

2.6.2 Adding a New Field to an Existing Header (SXADDPAR)

The routine SXADDPAR will allow a user to append or replace certain fields. Sample calls are:

IDL> sxaddpar, head, ’AVG OFF1’, sum arr(0) IDL> sxaddpar, head, ’FILENAM’, filnam

NOTE: The routine expects to have the ”END” field at the end of the array, and I believe that string has to be 80 characters long.

2.6.3 Building a Basic FITS Header (MK FITS HEAD)

MK FITS HEAD will build an elementry FITS header with the fields (SIMPLE, BITPIX, NAXIS, NAXIS1, NAXIS2, ...). You can optionally specify the time, date, comment or history lines. Some sample calls are:

IDL> head = mk fits head(data) IDL> head = mk fits head(data, time=’17:00:01’, date=’27-Oct-92’)

2.6.4 Reading a FITS File (RFITS2)

The routine RFITS2 will accept an array of file names, and will return a data cube with all of those images. NOTE: All images must be the same size for this option to work. The byte swapping depending on the platform is performed. Sample calls are:

IDL> imgs = rfits2(filenames, head=heads) IDL> img = rfits2(’kitt peak.mag’, head=head)

2.6.5 Writing an Image to a FITS File (WRT FITS)

The routine WRT FITS will save an image and header to a FITS file. The byte swapping depending on the platform is performed. Sample calls are:

IDL> wrt fits, outfil, head, image IDL> wrt fits, ’demo image.fits’, head, image

If the variable head is not defined, then an elementary FITS header is made using the routine MK FITS HEAD.

3 LFFT Routines

This chapter describes some of the standard data reduction routines which are available for the long form function test (LFFT) data.

3.1 MDIPICT

The routine MDIPICT produces a single page summary for each image in a series. It is normally run on data from the ”mdipict.src” control file, but it can be run on any image series. The summary includes:

• The image displayed in grey scale for the full range and for ±1 sigma around the average.

• Histogram plots in log and linear scale for the full image and also for ±2 sigma around the average (±1 sigma if the standard deviation is large)

• Single line and column plots, and plots of the average of all lines and all columns

• The image average, standard deviation, minimum, maximum, and range for both the whole image and for the image excluding the first and last two columns and lines.

• Information on the conditions and parameters for that image

• The location and amplitude of the 8 brightest pixels (excluding a 2 pixel margin around the edge of the image).

To run MDIPICT you must pass the telemetry file name. Sample calls are shown below. The second sample causes the output to go to a printer instead of the screen.

IDL> mdipict, ’/md66/log/sci160k/940401 143902’ IDL> mdipict, ’/md66/log/sci160k/940401 143902’, /hc IDL> mdipict, infil, tit=’MDIPICT Run in Vacuum, Warm’ IDL> mdipict, ’/md66/log/sci160k/940401 143902’, stimg=4, enimg=5

3.2 MDIFOCUS2

The routine MDIFOCUS2 produces a display of x and y-direction sharpness versus focus step position in full disk and high resolution pixels. It should only be run on the ”mdifocus.src” control file output which takes 73 images. The ”*” are overplotted from gaussian fits and the center for that gaussian fit is displayed for each plot. All four polarizations are overplotted on each of the four plots. Sample calls are:

IDL> mdifocus2, ’/md66/log/sci160k/940402 045113’ IDL> mdifocus2, ’/md66/log/sci160k/940402 045113’, img, sharp, norm, dark IDL> mdifocus2, ’/md66/log/sci160k/940402 045113’, /hc IDL> mdifocus2, ’/md66/log/sci160k/940402 045113’, tit=’Focus with new blocker’

In the second example, img, sharp, norm and dark are output variables holding the raw images, 3.2 MDIFOCUS2 13

Figure 1: Sample Output for MDIPICT the sharpness and normalization data arrays, and the dark image used. You can supply a title that is to appear on the hardcopy by supplying the keyword title.

3.3 LTC

The routine LTC calculates the light transfer function of the camera. The gain constant in electrons per data number (e/DN) is calculated along with the read noise for the camera. Some sample calls are:

IDL> ltc IDL> ltc, ’/md66/log/sci160k/940401 142048’, 120, 60000 IDL> ltc, ’/md66/log/sci160k/940401 142048’, 120, 60000, dev=’PS’

The values of 120 and 60000 are the minimum and maximum exposure times in ticks for the data set.

3.3.1 Details on the Light Transfer Calculation

1. SIGNAL: The signal level is the mean signal level in the image pair taken at a given exposure. The mean signal level is computed as follows; Each image is divided into several subimages (usually 256) of 16X16 pixels,positioned to cover most of the CCD area. The corresponding subimages from the two images are then averaged. Mean signal level is then calculated for each subimage. These mean signal levels are then averaged to give the mean signal level for the image pair. Dark current is NOT subtracted prior to the above calculation. Standard deviation is calculated from the 256 mean signal levels for each subimage.

2. NOISE: Average of the noise levels of all subimages in the image pair. The noise level of a subimage is the standard deviation of the signal level in that subimage. Standard deiation is calculated from the 256 noise levels for each subimage.

3. SHOT NOISE: Shot noise is the average shot noise level of all subimages. The shot noise of a subimage is computed by differencing the corresponding subimages in the image pair and then taking the standard deviation of the difference. The Standard deviation is calculated from 256 shot noise values for the subimages.

4. LINEARITY: Linearity is calculated by subtracting the ratio of mean signal at the middle exposure with the mean signal at each exposure, from 1 and multiplying the result by 100.

L_{k} = (1 − \frac{I_{m}}{I_{k}}) × 100

where I_{m} = Intensity (Dn/Exp) at the midpoint in exposures and I_{k} = Intesity at (k)th exposure. The linearity curves are calculated for all the sub-areas.

5. DARK CURRENT: Dark Current is determined from the two image pairs taken at 30 milliseconds and at 10 second exposures. The image pairs at each exposure are averaged to yield the average dark current at the two exposures. The Dark Current Rate is calculated 3.3 LTC 15

Figure 2: Sample Output for MDIFOCUS2 16 3 LFFT ROUTINES

Figure 3: Sample Output for LTC by,

DC_{rate} = \frac{(DC_{high}−DC_{low})}{(EXP_{high}−EXP_{low})}
DC_{offset} = DC_{low} − (DC_{rate} × EXP_{low})

From these quantities, the dark current at each exposure is calculated using the expression,

DC_{n} = DC_{offset} + (DC_{rate} × EXP_{n})

6. GAIN: The Gain is determined from the plot of Shot Noise vs Signal. It is calculated by deriving the expression for a straight line through the point representing the shot noise at the exposure number 7, with slope of 0.5. The Signal level corresponding to a shot noise of 1.0 is the Gain of the CCD.

7. FULL WELL:

F ullwell = MaxSignalLevel × Gain

8. READ NOISE:

ReadNoise = Readnoiseatlowestexposure × Gain

3.4 MDIISS

When the analysis program MDIISS is run, it displays a representative image and the areas it has selected for sharpness and intensity normalization calculations (HR only). The user hits < CR > to move on to the final result: the plot of sharpness curves, in the x- & y-direction, as a function of input jitter (amplitude in the x-axis, frequency with different linestyles in each panel). The upper panels are for the sine-wave ouputs, the lower ones are plotting the random input. The sharpness values are normalized to the un-distrubed case: i.e 1.0 means as sharp as the un-jittered image. A sample call is:

IDL> mdiiss, ’/md54/log/sci160k/940425 213204’, /hc

You can supply a title that is to appear on the hardcopy by supplying the keyword title. It is very involved to correct for MDIISS images being out of order, and the results will

simply not be meaningful at all unless the images are precisely as expected.

3.5 MDILASER

The routine MDILASER will take the data from the “mdilaser.src” control file and produces wavelength uniformity maps for the two Michelsons. The raw images which were used are displayed along with the Michelson maps and the sine curves which were fit. The transmission image and smaller contrast images are also displayed. A variety of information on the average signals and Michelson positions for the maximum transmission are printed on the summary 18 3 LFFT ROUTINES

Figure 4: Sample Output for MDIISS page. Some sample calls are:

IDL> mdilaser, ’/md66/log/sci160k/940402 033500’ IDL> mdilaser, infil, /hc

Nominal results should have two nice sine curves with good fits overplotted with dashed lines as shown in Figure 5. The M1 image should have a dark area in the lower right, and the M2 image should have a bright area in the right edge. Look for approximately 97 mA for the M2 wavelength co-tuning for when M1 wavelength is 50 mA.

3.6 MDIDUST

The routine MDIDUST will take the data from the “mdilaser.src” control file and produces images showing the dust particles on the different optical surfaces. The “mdilaser.src” data set takes CAL mode images using the laser as a source with nine different pointing angles. By shifting the images by different amounts, artifacts in the different surfaces are enhanced. This procedure can be used for both the high resolution and full disk images. Some sample calls are:

IDL> mdidust, ’/md66/log/sci160k/940402 033500’ IDL> mdidust, ’/md66/log/sci160k/940402 033500’, /hc IDL> mdidust, ’/md66/log/sci160k/940402 033500’, /fd IDL> mdidust, ’/md66/log/sci160k/940402 033500’, /fd, /obsmode

3.7 MDIMROT, MDIPROT, and MDIPDIST

The routines MDIMROT, MDIPROT, and MDIPDIST analyze the results from the respective control files. Each of them uses the routine FIT GCROSS. FIT GCROSS takes a grid image and locates all of the grid crossings. It then fits two gaussians to the X and Y averages to find the location of the grid crossing to within a few hundreds of a pixel. The results are saved in a text file, and the next time that one of these routines is run for the same set, the calculations are not performed. Instead, the routine reads the results from the text file and simply plots them. You can use the keyword /force calc to force the regeneration of the text file which contains the results of the grid locations.

NOTE: The stimulus image drifts with time (due to thermal changes and disturbances in the laboratory). The long term drift can be removed by smoothing the data. There is still some high frequency jitter in the stimulous evident by the lack of repeatablility of results for the same michelson tunings. The first few plots in these programs shows the raw, smoothed and differenced positions (see Figures 7, 8, and 9).

The MDIMROT experiment takes an image through 10 different M1 tunings and repeats those 10 images for a number of repeats (usually 10 or 15 repeats). It then does the same thing for 10 different M2 tunings. The analysis looks to see how much the grid location is shifted for each of the tunings. The results (see Figure 10) should show some distinct clustering for 20 3 LFFT ROUTINES

Figure 5: Sample Output for MDILASER 3.7 MDIMROT, MDIPROT, and MDIPDIST 21

Figure 6: Sample Output for MDIDUST 22 3 LFFT ROUTINES

Figure 7: MDIMROT Sample - Raw Grid Positions Figure 8: MDIMROT Sample - Positions Smoothed Figure 9: MDIMROT Sample - Remove Stimulus Drift M1 tunings, especially when comparing tuning 0 with tuning 6. The amplitude of the variation between average tuning positions (lower right plot) should be less than ± 0.2 pixels, and the scatter for each position should be less than ± 0.10 pixels. A sample call is:

IDL> mdimrot, ’/md54/log/sci160k/940426 050146’

The MDIPROT experiment is similar to MDIMROT, but there are four different polarizations instead. The results (see Figure ??) should show less than ± 0.05 pixel variation in the average polarization positions, and the scatter for each position should be less than ± 0.10 pixels. It is possible to run MDIPROT once with the ISS on and once with it off. If this is the case, then it is important to use the keyword option /qiss when calling MDIPROT.

IDL> mdiprot, ’/md54/log/sci160k/940426 060248’ IDL> mdiprot, ’/md65/log/sci160k/940505 104648’, /qiss

MDIPDIST takes full disk images of the grid stimulus through the four polarizations. There are typically 10 or 15 repeats of the four polarization set. The grid crossing is located for each grid and averaged. These average grid locations can be compared for different polarizations (the long term stimulous drift is effectively removed because the polarization images are interleaved).

The difference in the absolute pixel grid location between two polarizations is shown in Figure 12. This result shows the shift in the image between two polarizations, and the amplitude should be approximately 0.1 pixels.

Another plot can show the change in magnification or plate scale size. By chosing four grid locations which form a square, we can examine how the box size changes between the different polarizations. This is done by ratioing two polarization results. Figure 13 shows the results for ”left” versus ”right” polarization. A sample calling sequence is:

IDL> mdipdist, ’/md54/log/sci160k/940426 061816’, /hc

3.8 IPCOM

The routine IPCOM takes the data from the “ipcom.src” control file and produces a plot of the square root, reciprocal, and velocity tables. In addition to the plots, the routine also (a) checks the telemetry board compression and packing of the images, and (b) checks the image addition and subraction which can be performed by the DMA board. Sample calls are:

IDL> ipcom, ’/md66/log/sci160k/940417 033526’ IDL> ipcom, ’/md66/log/sci160k/940417 033526’, /hc 24 3 LFFT ROUTINES

Figure 10: MDIMROT Sample - Michelson 1 Position Shifts for Tunings 3.8 IPCOM 25

Figure 11: MDIPROT Sample - Position Shifts for Each Polarization 26 3 LFFT ROUTINES

Figure 12: MDIPDIST Sample - Position Shift for Polarization 3.8 IPCOM 27

Figure 13: MDIPDIST Sample - Plate Scale Size Change for each Polarization 28 3 LFFT ROUTINES

Figure 14: Sample Output for IPCOM

3.9 SHUTEST

This section still needs to be written

3.10 TRAP CHECK

It is recommended that the routine TRAP CHECK be run on the camera commision data (part of the ipcom control file). This routine monitors the size of the traps in the CCD and also checks for the development of new traps. Some sample calling sequences are:

IDL> trap check, ’/md65/log/sci160k/940503 110612’, /camera com IDL> trap check, ’/md65/log/sci160k/940506 120808’, /mdipict IDL> trap check, infil, [5,24] IDL> trap check, infil, fixed bad=[450, 362]

The total integrated signal for column 362 (column 661 for the B amplifier) should be approximately 1200 DN. If it is more than 1400 DN, the program manager should be alerted.

30 3 LFFT ROUTINES Figure 15: Sample Output for TRAP CHECK

4 House Keeping Routines

A list of the mnemonics, their meanings, and which words they come from is provided in the Appendix on page 45.

4.1 How to Make Temperature and Voltage HK Plots (HKPLOT)

It is possible to make plots of temperatures and voltages from the MDI housekeeping data by using the routine HKPLOT. NOTE: This routine will only run on SGI machines since the IDL routine spawns a call to the C executable READ HIST. It is also necessary to have defined the environment variable $MDI HISTORY DIR to point to the directory which has the .HIS files for the time period you need. A sample specification is:

% setenv $MDI HISTORY DIR /md54/log/history

It is necessary to know the names of the mnemonics which you wish to have plotted. In the second calling sequence is shown below, the output is sent to a tektronics plotting device.

IDL> hkplot IDL> hkplot, dev=’tek’

A sample session is shown below. You must specify the starting time and the duration you wish plotted.

HKPLOT - History files are assumed to be in directory: /md54/log/history
Enter date (DD-MMM-YY i.e. 01-oct-93) : 17-mar-94
Enter time (HH:MM:SS | HH:MM): 20:00
Enter duration (HH:MM:SS or HH:MM ): 10:00
Enter input file name or hit return for interactive mode
:
file name=
Interactive Mode
type 'plot <mnemonic> [y_low y_high | a]<CR>' for each mnemonic to plot
type 'end' to terminate input mode
Max of 4 mnemonics, Example:
    plot m+5cm      (default to auto y scale)
    plot m+5ip a    ( use auto y scale )
    plot m+5ip 5 30 ( specify y scale )
    plot m+5ip p    ( use previous y scale )
    oplot mtsqopt4  ( overplot filter oven temperature )
    plotave 2 mtopts4 mtopts5 20 30 ( plot ave oven temp)
    end

: plot mtopts5
: oplot mtopts6
: end
Enter output file name to save plot commands
  output file will go to directory: /md54/tmp
  or just hit return if no save desired
:
Using LVT file: /md54/log/history/READ_HIST.DBS
Select Hard Copy Option (a/0/1/2/3/p/<CR>):
Plot again? (<CR>=YES,y=YES,n=NO):

In the example above, the answers:

   plot mtopts5
   oplot mtopts6
   end

where supplied to plot the mnemonic mtopts5 and to over plot mtopts6. The Y-range plotted could have been specified to be from 10 to 45 by using a command syntax of the form:

   plot mtopts5 10 45
   oplot mtopts6
   end

4.2 GET HK INFO

The IDL routine GET HK INFO directly accesses the database which holds the MDI and spacecraft house keeping packets (not the command .HIS files). It uses the same mnemonics as HKPLOT and RTPLOT. You can either get the value of the mnemonic for a single time, a list of times, or a range of times. Some sample calls are:

IDL> info = get hk info(times, mnem) IDL> print, get hk info(’16-mar-94 09:00’, ’mtopts5’) IDL> info = get hk info([’16-mar-94 5:00’, ’16-mar-94 14:00’], /range, ’mtopts5’) IDL> info = get hk info(timerange, /range, mnem, timarr=timarr)

The input times can be a single time or a list of times. If the /range option is going to be used, then times must be the start and stop times (a two element array). mnem can be a list of mnemonics if the output are all of the same data type (some outputs are floating point and others are string type to say something like “ON” or “OFF”). The output timarr is the time array for each of the values returned. You can plot these results using UTPLOT with a command similar to the following:

IDL> val = get hk info([’16-mar-94 5:00’, ’16-mar-94 14:00’], /range, $ IDL> [’mtopts5’, ’mtopts6’], timarr=timarr) IDL> utplot, timarr, val(*,0) IDL> outplot, timarr, val(*,1)

5 Calibration Routines

5.1 Michelson Tuning Analysis (DO DISP MMAP1)

The routine DO DISP MMAP1 allows the user to perform the Michelson tuning analysis on most datasets (you can use MDILASER on the nominal laser tuning datasets). It is a program, so you type the following to run it:

IDL> .run do disp mmap1

A sample session is shown below.

------------------------------------------------------------
For LFFT sets, the following parameters are needed:
    "Bogus" image offset:                                   1
    # Images per position:                                 27
    Dark Image # within the "one position" series           6
    Obs Image # within the "one position" series            0
    Image # of first of 20 Michelson tuning                 7
There are 9 positions available (answer 0-8)
------------------------------------------------------------
Do you wish to select the latest telemetry file  [Default: YES ] ?n
********************************************************************************
Absolute image numbers start counting at zero (first image = image #0)
Telemetry File Name (including directory):       $MDI_SCI160K_LOG_DIR/940312_052339
Directory for the telemetry files:               /disk6/log/sci160k
"Bogus" image offset:                            0
# Images per position:                           22
Dark Image # within the "one position" series    1
Obs Image # within the "one position" series     0
Image # of first of 20 Michelson tuning          2
Automatically save Michelson maps                Yes
Enter position(s) to run   [Default: 0 ]
% Compiled module: TYPE_CONV.
Enter the repeat to run   [Default: 0 ]
Ready to run DISP_MMAP now?   [Default: YES ] ?
Automatically generate a hardcopy                Yes
********************************************************************************
Do you wish to change any of the above parameters?   [Default: N ] ?
--------------------------------------------------------------------------------
To run all 9 positions of "mdilaser" enter:    0,1,2,3,4,5,6,7,8
Enter position(s) to run   [Default: 0 ]

Enter the repeat to run   [Default: 0 ]
Ready to run DISP_MMAP now?   [Default: YES ] ?

If you answer “No” to the question about changing any of the above parameters, you can adjust which images are used for the tuning or the telemetry file name.

6 Command History Routines

Text files have been created with contain a list of the control files and MDI commands which have been executed. The files which are created are in the directory $MDI CAL TXT (which is normally /mdisw/dbase/cal/summary). There is a file per month with names of the form cmd hist.9405 for May-94. A list of the .HIS files which were used to create the summary list is shown in the beginning of each file.

A sample portion of the full listing is:

940504_095350 <         mbltyofs !LI2                          4-MAY-94 09:49:55
940504_095350       mbcmtp3 0 0 4000                           4-MAY-94 09:51:16
940504_095350       mbltclsl                                   4-MAY-94 09:51:41
940504_095350       mbcmtp3 0 0 4000                           4-MAY-94 09:51:58
940504_095350    >> mdipict                                    4-MAY-94 09:52:42
940504_095350           mbcmsumm 1 1                           4-MAY-94 09:52:44
940504_095350           mbiplbic 0 3                           4-MAY-94 09:52:45
940504_095350           mbcmhdr1 0x4040 0x4040 0 1 0           4-MAY-94 09:52:46
940504_095350           mbcmaadc 3                             4-MAY-94 09:52:47
940504_095350           lobsmode                               4-MAY-94 09:53:15
940504_095350       pconf 0x0100 0x0100 0x0100 0x0100 0x0100   4-MAY-94 09:53:18
940504_095350       pconf     13     82    125     55     89   4-MAY-94 09:53:19
940504_095350           mbcmtp1r !GI7 !GI0 !GI0                4-MAY-94 09:53:51
940504_095350           mbcmtp3r !GI7 !GI0 !GI0                4-MAY-94 09:55:43
940504_095350           lcalmode                               4-MAY-94 09:57:34
940504_095350       pconf 0x0100 0x0100 0x0100 0x0100 0x0100   4-MAY-94 09:57:37
940504_095350       pconf     13     82    125     19    125   4-MAY-94 09:57:38
940504_095350           mbcmtp1r !GI7 !GI0 !GI0                4-MAY-94 09:58:05
940504_095350           mbcmtp3r !GI7 !GI0 !GI0                4-MAY-94 09:59:56
940504_095350           mbcmdrkr !GI7 !GI0 !GI0                4-MAY-94 10:01:48
940504_095350           mbcmbadc 3                             4-MAY-94 10:03:40
940504_095350           mbcmdrkr !GI7 !GI0 !GI0                4-MAY-94 10:03:41

The symbol < signifies that this is the first command which is associated with the telemetry file, and the symbol >> signifies that it is a command which a user executed (used the “perf” command). It is simple to “grep” the file to search for certain conditions.

The alias cmdgrep has been created to search all of the files for a given string. A sample use of that symbol is:

% cmdgrep mdipict

to find all instances when mdipict was run.

6.1 HIST SEA

HIST SEA is the routine which creates the command history summary files. It is a program and a sample call is:

IDL> .run hist sea

6.2 HIST SUMMARY

The routine HIST SUMMARY will search through all of the $MDI CAL TXT/cmd hist.9* files for every instance that a control file is run. A sample of the default HIST SUMMARY results are shown in the appendix on page 71. Some sample calls are:

IDL> hist summary IDL> hist summary, /default IDL> hist summary, [’mdipict’, ’mdiltc’, ’mdifocus’], outfil=’test.hist’

6.3 HIST COUNT

The routine HIST COUNT searches through all of the $MDI CAL TXT/cmd hist.9* files and counts the number of times a control file is executed for each month. A sample listing is shown in the appendix. Some sample calls are:

IDL> hist count IDL> hist count, /default IDL> hist count, [’mdipict’, ’mdiltc’, ’mdifocus’], outfil=’count.hist’

A Description of the Purpose of Several Control Files

This appendix describes some of the details on a few of the more common control files. For more details you can look at the long form function test (LFFT) procedure or at the source file of the control file.

A.1 MDION

MDION opens a log file if one is not already open. Turns on MDI, except for heaters and power to the front door, the alignment mechanism, and the limb tracker; and resets subsystems to known states. The operator selects the redundant or primary MDI power converters in response to a prompt. After several prompts, a picture is taken (in obs mode) to demonstrate that the instrument and EGSE are functioning nominally. If hung up, manual commands may be required and additional pictures/images should can be taken.

A.2 MDICOM

MDICOM calls a series of control files in order to exercise (commission) the majority of the MDI subsystems. Contains a few operator options for subsystems like the Front Door and Limb Tracker whose use depends on the test configuration. The test flow is as given below.

1. A set of instrument temperatures are recorded. Then, structure heater #3 (fixed power) and the primary oven controller (Set Point 2) are turned on to begin warming things up early in the test flow so that temperature changes will be observable later.

2. The five internal mechanisms (MTM1, MTM2, PAW, CAL1, and CAL2) are simultaneously moved to four different locations and the resultant encoder values are put in the log file.

3. Power to the alignment mechanism is turned on. MDI is pointed to a new location by moving both motors and the potentiometer values are put in the log file. MDI is then returned to the initial pointing.

4. Camera and shutter functionality is demonstrated by taking six images and down linking them. There is no light because the door is closed. In addition, the measured shutter times are put in the log file. The six images are:

Image#0: 1024x1024 1x1 summed dark image, offset 3 (Amp A)

Image#1: 1024x1024 1x1 summed dark image, offset 3 (Amp B)

Image#2: 1024x1024 2x2 summed dark image, offset 3 (Amp A)

Image#3: 1024x1024 2x2 summed dark image, offset 3 (Amp B)

Image#4: 1024x1024 1x1 summed dark image, offset 15 (Amp A)

Image#5: 1024x1024 1x1 summed dark image, offset 15 (Amp B)

This set of images exercises both the A and B amplifiers, the 2x2 camera summing mode, the minimum and maximum offset levels (thus one of the images is put into a state where all data bits are 0’s and one to where nearly all are 1’s), and the ability to output data through the IP.

5. If the stimulus telescope is to be used and the front door is to be opened, the procedure will prompt the user to open the front door and turns on the limb tracker in the nominal stimulus telescope mode. MDICOM then activates LLTCNTR which takes continuous FD Obs mode images, summed 4x4 in the IP to make the repetition rate acceptable. At a prompt the operator manually moves the stimulus telescope until the image is well centered on the CCD. Continuing on, LLTCNTR iteratively adjusts the X and Y Offsets until the X and Y Average Errors are between 2.2 and 2.4.

6. The limb tracker system is exercised, first with the loop open. If there is no stimulus telescope this only consists of changing the X and Y offsets and noting changes in the X and Y offset readings, and changing the PZT A, B, and C offsets and noting changes in the PZT positions (multiply by 20 to obtain volts). If the stimulus is on, changing X and Y offsets will affect the average error signals as will changing the PZT positions. If the stimulus is on, LLTCOM continues by moving the alignment legs in a small square ( 8”) and records the corresponding changes in the error signals with the loop open. The loop is then closed and both the PZT offsets are changed and the PZT positions should now remain unchanged. Next, the square pattern of pointing is repeated and the error signals should now remain unchanged while the PZT positions do vary to compensate for the off-pointing. Finally, the PA wheel is moved to verify that the DEP software is automatically adjusting the ISS to handle this interruption of the beam. The loop is then opened and power is removed from the alignment mechanism and the ISS.

7. Provided nothing is in the way of opening the door and the cleanliness conditions are acceptable, power to the front door is turned on and if it is open already it is closed. Next the door is opened and closed with motor 1 and then with motor 2. The time it takes for the door to open and close are recorded as an indication of performance. Power is then turned off to avoid accidentally moving the door at an undesired time.

8. The prime oven controller is stepped through five set points. Time is not taken to see corresponding temperature changes. The oven power is then turned off.

9. The backup oven power is turned on and five set points are exercised. At each set point, the INTSIM current is recorded. The backup oven power is then turned off.

10. The optics package structure heater power is turned on and six settings are exercised. At each setting, the INTSIM current is recorded. The structure heater power is then turned off.

11. The CCD main/decontamination heater is turned on, the new (higher) INTSIM current is recorded as is the temperature. Then the CCD heaters is turned off and a full set of temperatures are recorded.

12. The DEP is exercised. The state of the Sequence Registers is put in the log file, the registers are changed, and the new states are put in the log file. An MTM Lookup Table is loaded and the load verified. An optional oven monitoring activity is accomplished if the operator feels conditions are such that it will be meaningful. Next, a sequence is run for about three minutes which takes full disk, 1024x1024 pixel images; sums pixels 4x4 in the IP; and outputs 256x256, uncompressed images via the high rate telemetry channel. After that, a sequence is loaded via telecommands and run for another three minutes. The high rate data is not saved to disk.

13. The IP is exercised.

• The first part of the IP commissioning verifies the capability to load and read back the instruction queue, to load tables into memory, and to transfer data using the double buffer.

• The control file will next perform an IP memory test. Memories are tested in groups of three. There are twenty IP memory pages, so there will be seven tests (one memory is tested twice).

• The control next does a check of the telemetry board by loading the IP with a test pattern and downlinking that pattern in raw, compressed, and packed form.

• The control file next verifies that tables have been loaded to the IP correctly. In the process of doing so, the control file executes macro programs from the queue, thus verifying DEP queue control.

• The control next verifies the DMA controller interface to the camera. An image is taken with the camera and loaded into 4 (0-3) memories. A second image is loaded into memory 1, subtracted from memory 2, and added to memory 3. All 4 images are then downlinked.

• Next, the control file loads the queue with instructions that execute the dynamics program and then dumps the queue. The control file asks if the dump of the queue is okay.

• Lastly, fully integrated DEP, camera, and IP functionality is demonstrated. The control file will invoke a DEP sequence that executes one full cycle of the dynamics program. Full disk velocity and square root compressed continuum intensity are downlinked on the high rate channel. A pair of images are displayed. The sequence is allowed to repeat three more times. Following the third repeat, the intermediate sums used to generate the third pair of images are downlinked and will later be used to verify successful execution of the velocity algorithm.

14. If the operator chooses, the Degraded Motor Controller (DMC) is exercised by the DEP. Each mechanism relay is set to the DMC state and the encoder values are recorded. Then, each internal motor except the shutter is moved to a few positions and the encoders are recorded. The motors are moved back and the encoders are recorded again. Lastly, the DMC relays are put back to normal.

15. The DMC capability of the alignment mechanism is verified by switching the appropriate relay and moving the legs over a considerable distance. Then, the DMC relays are put back to normal.

16. “perf mdihoc” to exercise the new/controllable CCD heater. Presently the control file begins by loading a patch to the DEP, but this will not be needed when a future version of the DEP software is resident. It exercises each of the three new heaters individually and then the combination of all three. At each step the operator records the new INTSIM current and after the last step the variation (increase) in CCD header/strap temperature

is noted.

A.3 MDIPICT

MDIPICT is used in order to carefully examine CCD cosmetics and/or MDI contamination, as well as to enhance the evaluation of the CCD camera system. This control file takes the following images:

Image#0: 1024x1024 FD Obs Mode image (Amp A)

Image#1: 1024x1024 HR Obs Mode image (Amp A)

Image#2: 1024x1024 FD Cal Mode image (Amp A)

Image#3: 1024x1024 HR Cal Mode image (Amp A)

Image#4: 1024x1024 dark image (Amp A)

Image#5: 1024x1024 dark image (Amp B)

Image#6: 1024x1024 HR Cal Mode image (Amp B)

A.4 MDILTC

MDILTC produces two classical light transfer curves by taking 12, HR Cal, exposure pairs with each (A & B) amplifier chain. Only the central 256x256 pixels of data are down linked through the IP in order to keep the required test time ( 8 min) acceptable.

A.5 MDIFOCUS

MDIFOCUS steps through all 9 focus positions, first FD and then in HR mode; and also takes a software dark. Repeats the sequence for each PAW position. Only the central 256x256 pixels of data are down linked through the IP in order to keep the required test time ( 15 min) acceptable.

A.6 MDIISS

MDIISS collects open loop and closed loop HR images for a matrix of 3 different input jitter frequencies and 3 different amplitudes. These are 1, 10, and 45 Hz and 100, 400, and 1000 volts (peak-to-peak) on the jitter mirror PZT which moves at 45 degrees to the MDI axes thus imparting approximately +.2, +.7, and +1.7 arc sec of jitter. Also collects open and closed loop images for 3 different amplitudes of random input jitter if an appropriate noise generator is available. These are band width limited at 256 Hz and have a peak-to-peak amplitude of 100, 400, and 1000 V. Only the central 256x256 pixels of data are down linked through the IP in order to keep the required test time ( 15 min) acceptable.

A.7 MDIDSOS

MDIDSOS exercises the degraded science operating system with the prime MDI converters powered by the spacecraft bus B. Demonstrates a dummy block command and the ability to do a memory load and dump. Demonstrates a dummy command and loads a simple DSOS program. This program produces a stable telemetry image of the loaded program in HK1 and sets the prime oven setpoint to 2, the structure heater setpoint to 128, and the new/controllable CCD heater setpoint to 1.

A.8 MDIMROT

MDIMROT looks for image movement with the stepping of a single tuning waveplate, first MTM1 and then MTM2. Each motor is stepped through 10 positions and the sequence for each motor is repeated a least 10 (selectable, with a default of 15) times. Only the central 256x256 pixels of data (HR Obs) are down linked through the IP in order to keep the required test time (57 min) acceptable. This is done with the ISS locked. It is also done using the B amplifier chain of the CCD so the image is reversed from ”normal”.

A.9 MDIPROT

MDIPROT examines image motion associated with changing positions of the PAW by taking a HR image in each of the 4 locations and then repeating the activity at least 10 (selectable, with a default of 15) times. This is done first with the ISS loop open and then with it closed. Only the central 256x256 pixels of data (HR Obs) are down linked through the IP in order to keep the required test time (24 min) acceptable.

A.10 MDIPDIST

opens the ISS loop and examines image distortions as a function of changing the 4 PAW positions. Does at least 3 (selectable, with a default of 5) repeats of the 4 positions taking FD 1024x1024 images. This requires about 44 minutes.

A.11 MDIDUST

MDIDUST collects a series of images using a sub-aperture mask in order to look for cosmetic non-perfections (”dust”) within the MDI optical train.

A.12 MDIKAL

MDIKAL is used to verify functionality of the Keep Alive Power. A pattern of information is loaded into IP table memory and then MDI power converter #1 is turned off. The converter is turned back on and it is verified that the information is still in the IP table memory.

A.13 SHUTEST

SHUTEST takes 100 each 1X and 3X exposures and puts the measured shutter times into the log file. Requires about 25 minutes.

A.14 MDILASER

The laser fiber is fed into the stimulus telescope at its telecentric focus thereby providing a point-source with a spectral delta function. MDILASER uses the alignment mechanism (legs) to point MDI in the 8 cardinal directions (radius user-selectable) plus center. At each pointing (= incident angle) a series of images in CAL mode are taken: a short Michelson de-tune, a strategic selection of images to search for contaminants, as well as a dark and an OBS mode image to identify the particular pointing.

A.15 DMCDOOR1

DMCDOOR1 verifies the functionality of the front door under DMC control. It is not part of any SFFT or LFFT because it can potentially damage the door mechanism.

44 B LISTING OF MDI DATA PRODUCT CODES (DPC)

B Listing of MDI Data Product Codes (DPC)

;FILE:
;       $MDI_CAL_INFO/dpc_table.txt
;HISTORY:
;       Created 18-Apr-94 by M.Morrison
;        5-May-94 (MDM) - Adjusted format slightly.  Added Mnemonic column
;
; For Summing:
;       0 means non-image, 1 is full resolution, 4 is 4x4 IP summing
; For NX:
;       -1 means variable length
;
; Data Product    NX      NY    Summing         Mnemonic                        Comments
;   Code
'FFFFFF01'XL     256     256    4       256x256_sum_5fast       # 4 x 4 summed image, 5 images taken fast and stored in IP
'FFFFFF02'XL     256     256    4       256x256_sum             # 4 x 4 summed image
'FFFFFF03'XL     256     256    1       256x256_ext             # 256 x 256 image extracted from 1024x1024
'FFFFFF04'XL    1024    1024    1       1024x1024               # 1024 x 1024 image
'FFFFFF15'XL     128     128    1       128x128_ext             # 128 x 128 image extracted from 1024x1024
'FFFFFF40'XL    1024    1024    1       1024_velocity           # 1024X1024 Velocity
'FFFFFF41'XL    1024    1024    1       1024_continuum          # 1024X1024 square root compressed continuum
'FFFFFF42'XL    1040    1033    1       1024_raw_i1-i3          # i1-i3 raw
'FFFFFF43'XL    1040    1033    1       1024_raw_i2-i4          # i2-i4 raw
'FFFFFF44'XL    1040    1033    1       1024_raw_continuum      # raw continuum
'FFFFFF82'XL    1024    1024    1       mathur_test             # D.Mathur test IP data product
'FFFFFF84'XL    1024    1024    1       1024x1024_comp          # 1024 x 1024 telemetry compressed image from IP
'FFFFFF50'XL    1040    1033    1       ipdmatst_raw            # ipdmatst - raw image
'FFFFFF51'XL    1040    1033    1       ipdmatst_sum            # ipdmatst - DMA addition of two raw images (1st - 2nd)
'FFFFFF52'XL    1040    1033    1       ipdmatst_diff           # ipdmatst - DMA difference of two raw images
'FFFFFF70'XL    '4000'XL  1     0       iptbltst_sqrt           # iptbltst - SQRT table
'FFFFFF71'XL    '8000'XL  1     0       iptbltst_reciprocal     # iptbltst - reciprocal table
'FFFFFF72'XL    '8000'XL  1     0       iptbltst_velocity       # iptbltst - velocity table
'FFFFFF73'XL    -1      1       0       crop_table              # crop table
'FFFFFF60'XL    1040    1033    0       ipmemtst_raw            # ipmemtst - FILL pattern raw
'FFFFFF61'XL    1040    1033    0       ipmemtst_compressed     # ipmemtst - FILL pattern compressed
'FFFFFF62'XL    1040    1033    0       ipmemtst_packed         # ipmemtst - FILL pattern packed
'10000000'XL    -1      1       0       unknown                 # ??

C Listing of House Keeping Mnemonics

PR_HK_INFO.PRO   Ver 1.0
mkpkfmid         Packet Format ID                  MDIHK    0  ff
mkconfr          Configuration Register            MDIHK    1  ff
m+5ae            App. Elec  +5 volts; Cnv2         MDIHK    2  ff
m+15ae           App. Elec  +15 volts; Cnv2        MDIHK    3  ff
m-15ae           App. Elec  -15 volts; Cnv2        MDIHK    4  ff
m+75lt           Limb Tracker +75 volts; Cnv2      MDIHK    5  ff
m+15htr          Heater +15 volts; Cnv2            MDIHK    6  ff
m+15mtr          Motor +15 volts; Cnv2             MDIHK    7  ff
m+5ip            IP AND TLM +5 volts; Cnv1         MDIHK    8  ff
m+5cm            CCD Camera +5 volts; Cnv1         MDIHK    9  ff
m+15cm           CCD Camera +15 volts; Cnv1        MDIHK   10  ff
m-15cm           CCD Camera -15 volts; Cnv1        MDIHK   11  ff
m+30cm           CCD Camera +30 volts;Cnv1         MDIHK   12  ff
mkpwae           Application I/F Power (K2-2)      MDIHK   12  80
mkpwaer          Application I/F Power RAW         MDIHK   13  80
mkpwhpo          Prime Oven Heater Power    (K17)  MDIHK   13  40
mkpwhop          Optics Package heater (K18)       MDIHK   13  20
mkpwlt           Limb Tracker Power   (K2-3)       MDIHK   13  10
mkpwsh           Shutter Power  (K2-4)             MDIHK   13  08
mkpwm1           MTM 1 Power    (K2-5)             MDIHK   13  04
mkpwm2           MTM 2 Power    (K2-6)             MDIHK   13  02
mkpwc1           Cal 1 Power    (K3-1)             MDIHK   13  01
mkpwc2           Cal 2 Power    (K3-2)             MDIHK   14  80
mkpwpa           PAW Power      (K3-3)             MDIHK   14  40
mkpwfd           Door Power     (K3-4)             MDIHK   14  20
mkpwam           Alignment  Power  (K3-5)          MDIHK   14  10
mkpwdm           DMC Power      (K3-6)             MDIHK   14  04
mkpwip           Image Processor Power   (K1-2)    MDIHK   14  02
mkpwcm           CCD Camera Power     (K1-1)       MDIHK   14  01
mktmsel          IP TLM 1-2 Power (K1-6)           MDIHK   15  80
mkpwhro          B/U Oven Heater Pwr  (K29)        MDIHK   15  40
mkpwhcm          CCD Heater Power           (K30)  MDIHK   15  20
mkltloop         ISS Loop Open/Closed              MDIHK   15  10
mkfdopn1         Door Open Status 1                MDIHK   15  02
mkfdopn2         Door Open Status 2                MDIHK   15  08
mkfdmtr1         Door Motor 1 Status               MDIHK   15  01
mkfdmtr2         Door Motor 2 Status               MDIHK   15  04
mkfdtime         Door Move Time                    MDIHK   16  ff

mkfdfunc         Door Function                     MDIHK   17  38
mkfdstat         Door Operation Status             MDIHK   17  04
mkfdmove         Door Moving Indicator             MDIHK   17  02
muhopctl         Optics package Heater Controller  MDIHK   18  ff
muhpobrg         Filter Oven Bridge Out            MDIHK   19  ff
muhpolop         Filter Oven Loop Out              MDIHK   20  ff
muhroctl         Filter Oven Redundant Controller  MDIHK   21  ff
mtopts1          Front Window Mounting Ring        MDIHK   22  ff
mtopts2          Primary Lens                      MDIHK   23  ff
mtopts3          Secondary Lens                    MDIHK   24  ff
mtopts4          Filter Oven                       MDIHK   25  ff
mtopts5          Filter Oven                       MDIHK   26  ff
mtopts6          CCD Strap/Header (P)              MDIHK   27  ff
mtopts7          CCD Electronics                   MDIHK   28  ff
mtopts8          OP Lid Near CCD Radiator (P)      MDIHK   29  ff
mtopts9          Rear Leg/Struct Interface         MDIHK   30  ff
mtopts10         PA Calib Wheel                    MDIHK   31  ff
mtopts11         Outside MLI                       MDIHK   32  ff
mtopts12         Wire Harness/OP Struct            MDIHK   33  ff
mtopts13         OP Struct Inside Top              MDIHK   34  ff
mtopts14         OP Struct Oven Ctrl               MDIHK   35  ff
mtopts15         OP Struct Inside Front            MDIHK   36  ff
mtopts16         Beam Dist. Housing                MDIHK   37  ff
mtepts1          Outside Structure                 MDIHK   38  ff
mtepts2          Redundant Converter 2             MDIHK   39  ff
mtepts3          Primary Converter 2               MDIHK   40  ff
mtepts4          Outside Side Radiator             MDIHK   41  ff
mxhporef         Temperature Sensor Reference      MDIHK   42  ff
mkam1cws         AM1 Clockwise Switch              MDIHK   43  40
mkam1ccs         AM1 Counterclockwise Switch       MDIHK   43  80
mxam1pos         Leg Potentiometer Wiper 1         MDIHK   43  03
mkam2cws         AM2 Clockwise Switch              MDIHK   45  40
mkam2ccs         AM2 Counterclockwise Switch       MDIHK   45  80
mxam2pos         Leg Potentiometer Wiper 2         MDIHK   45  03
miaemtr          Motor Current Monitor             MDIHK   47  ff
mktccnt          Command counter                   MDIHK   48  ff
mktcnvcb         Non-valid Command Blocks          MDIHK   49  f0
mktcvcb          Valid Command Blocks              MDIHK   49  0f
mktclnvc         MLA of Latest NVCB                MDIHK   50  ff
mktcnver         Error code of latest NVCB         MDIHK   52  ff
mktciist         Intra-Instrument Status           MDIHK   53  80
mktciim          Intra-Instrument Master           MDIHK   53  70

mktciix          Intra-Instrument X Coordinate     MDIHK   53  03
mktciie          Intra-Instrument Event            MDIHK   55  70
mktciiy          Intra-Instrument Y Coordinate     MDIHK   55  03
mkpapos          Polarization Analyzer Wheel       MDIHK   57  ff
mkc1pos          Cal/focus wheel 1                 MDIHK   58  ff
mkc2pos          Cal/focus wheel 2                 MDIHK   59  ff
mkm1pos          Michelson Tuning Motor 1          MDIHK   60  ff
mkm2pos          Michelson Tuning Motor 2          MDIHK   61  ff
multdio1         Limb Tracker Photodiode 1         MDIHK   62  ff
multdio2         Limb Tracker Photodiode 2         MDIHK   63  ff
multdio3         Limb Tracker Photodiode 3         MDIHK   64  ff
multdio4         Limb Tracker Photodiode 4         MDIHK   65  ff
multpzt1         Limb Tracker PZT1                 MDIHK   66  ff
multpzt2         Limb Tracker PZT2                 MDIHK   67  ff
multpzt3         Limb Tracker PZT3                 MDIHK   68  ff
multxmax         Limb Tracker X Max Error          MDIHK   69  ff
multymax         Limb Tracker Y Max Error          MDIHK   70  ff
multxave         Limb Tracker X Average Error      MDIHK   71  ff
multyave         Limb Tracker Y AverageError       MDIHK   72  ff
mkltxofs         Limb Tracker X offset             MDIHK   73  ff
mkltyofs         Limb Tracker Y offset             MDIHK   74  ff
mkltxgn          Limb Tracker X gain               MDIHK   75  ff
mkltygn          Limb Tracker Y gain               MDIHK   76  ff
mkltaos          Limb Tracker A offset             MDIHK   77  ff
mkltbos          Limb Tracker B offset             MDIHK   78  ff
mkltcos          Limb Tracker C offset             MDIHK   79  ff
mkltstat         Limb Tracker Status Register      MDIHK   80  ff
mkovnmax         Filter Oven Max Temp              MDIHK   81  ff
mkhpospt         Prime Oven Set Point              MDIHK   82  e0
mkhpomon         Prime Oven Monitoring             MDIHK   82  10
mkhpomtw         Prime Oven Monitoring Time        MDIHK   82  0f
mkhrodbh         Backup Oven Dead Band High        MDIHK   83  ff
mkhrodbl         Backup Oven Dead Band Low         MDIHK   84  ff
mkhrocbh         Backup Oven Control Band High     MDIHK   85  ff
mkhrocbl         Backup Oven Control Band Low      MDIHK   86  ff
mkhrospt         Backup Oven Set Point             MDIHK   87  ff
mkhrocpt         Backup Oven Control Point         MDIHK   88  ff
mkhcmspt         CCD Heater Set Point              MDIHK   89  e0
mkhromon         Backup Oven Monitoring            MDIHK   89  01
mkhoph2          Optics Package Heater 2 On/Off    MDIHK   90  80
mkhoph3          Optics Package Heater 3 On/Off    MDIHK   90  40
mkhopspt         Optics Package Heater Set Point   MDIHK   90  3f

mkshnexp         Shutter Nominal Exposure          MDIHK   91  ff
mkpadir          PAW Default Direction             MDIHK   93  08
mkm1dir          MTM1 Default Direction            MDIHK   93  04
mkm2dir          MTM2 Default Direction            MDIHK   93  02
mkshpath         Shutter Default Beam Path         MDIHK   93  01
mkdpmerr         Correctable DEP memory errors     MDIHK   94  ff
mkipmerr         Correctable IP memory errors      MDIHK   95  ff
mkdpm0os         Offset for MTM0                   MDIHK   96  ff
mkdpm1os         Offset for MTM1                   MDIHK   97  ff
mkam1cmd         AM1 Commanded position            MDIHK   98  ff
mkam2cmd         AM2 Commanded position            MDIHK  100  ff
mkshenc          Shutter Encoder Position          MDIHK  102  ff
mkund01          Undefined Byte                    MDIHK  103  ff
mkund02          Undefined Byte                    MDIHK  104  ff
mksqintc         Interval Counter                  MDIHK  105  ff
mksqfqrt         Ref Time occurred in packet       MDIHK  107  80
mksqfqrp         Frame Queue Pointer at Ref Time   MDIHK  107  60
mksqfqpp         Frame Queue Pointer for Packet    MDIHK  107  18
mksqfqne         Number of Frame Queue Entrie      MDIHK  107  07
mksqrt1          Relative Time of Frame 1          MDIHK  108  ff
mksqsho1         Shutter Open Time Frame 1         MDIHK  110  ff
mksqshc1         Shutter Close Time Frame 1        MDIHK  112  ff
mksqpaa1         PAW Actual Position Frame 1       MDIHK  114  ff
mksqm1a1         MTM1 Actual Position Frame 1      MDIHK  115  ff
mksqm2a1         MTM2 Actual Position Frame 1      MDIHK  116  ff
mksqc1a1         CAL1 Actual Position Frame 1      MDIHK  117  ff
mksqc2a1         CAL2 Actual Position Frame 1      MDIHK  118  ff
mksqadr1         Sequence Address of Frame 1       MDIHK  119  ff
mksqun11         Undefined                         MDIHK  121  ff
mksqun21         Undefined                         MDIHK  122  ff
mksqun31         Undefined                         MDIHK  123  ff
mksqun41         Undefined                         MDIHK  124  ff
mksqun51         Undefined                         MDIHK  125  ff
mksqun61         Undefined                         MDIHK  126  ff
mksqun71         Undefined                         MDIHK  127  ff
mksqrt2          Relative Time of Frame 2          MDIHK  128  ff
mksqsho2         Shutter Open Time Frame 2         MDIHK  130  ff
mksqshc2         Shutter Close Time Frame 2        MDIHK  132  ff
mksqpaa2         PAW Actual Position Frame 2       MDIHK  134  ff
mksqm1a2         MTM1 Actual Position Frame 2      MDIHK  135  ff
mksqm2a2         MTM2 Actual Position Frame 2      MDIHK  136  ff
mksqc1a2         CAL1 Actual Position Frame 2      MDIHK  137  ff

mksqc2a2         CAL2 Actual Position Frame 2      MDIHK  138  ff
mksqadr2         Sequence Address of Frame 2       MDIHK  139  ff
mksqun12         Undefined                         MDIHK  141  ff
mksqun22         Undefined                         MDIHK  142  ff
mksqun32         Undefined                         MDIHK  143  ff
mksqun42         Undefined                         MDIHK  144  ff
mksqun52         Undefined                         MDIHK  145  ff
mksqun62         Undefined                         MDIHK  146  ff
mksqun72         Undefined                         MDIHK  147  ff
mksqrt3          Relative Time of Frame 3          MDIHK  148  ff
mksqsho3         Shutter Open Time Frame 3         MDIHK  150  ff
mksqshc3         Shutter Close Time Frame 3        MDIHK  152  ff
mksqpaa3         PAW Actual Position Frame 3       MDIHK  154  ff
mksqm1a3         MTM1 Actual Position Frame 3      MDIHK  155  ff
mksqm2a3         MTM2 Actual Position Frame 3      MDIHK  156  ff
mksqc1a3         CAL1 Actual Position Frame 3      MDIHK  157  ff
mksqc2a3         CAL2 Actual Position Frame 3      MDIHK  158  ff
mksqadr3         Sequence Address of Frame 3       MDIHK  159  ff
mksqun13         Undefined                         MDIHK  161  ff
mksqun23         Undefined                         MDIHK  162  ff
mksqun33         Undefined                         MDIHK  163  ff
mksqun43         Undefined                         MDIHK  164  ff
mksqun53         Undefined                         MDIHK  165  ff
mksqun63         Undefined                         MDIHK  166  ff
mksqun73         Undefined                         MDIHK  167  ff
mksqrt4          Relative Time of Frame 4          MDIHK  168  ff
mksqsho4         Shutter Open Time Frame 4         MDIHK  170  ff
mksqshc4         Shutter Close Time Frame 4        MDIHK  172  ff
mksqpaa4         PAW Actual Position Frame 4       MDIHK  174  ff
mksqm1a4         MTM1 Actual Position Frame 4      MDIHK  175  ff
mksqm2a4         MTM2 Actual Position Frame 4      MDIHK  176  ff
mksqc1a4         CAL1 Actual Position Frame 4      MDIHK  177  ff
mksqc2a4         CAL2 Actual Position Frame 4      MDIHK  178  ff
mksqadr4         Sequence Address of Frame 4       MDIHK  179  ff
mksqun14         Undefined                         MDIHK  181  ff
mksqun24         Undefined                         MDIHK  182  ff
mksqun34         Undefined                         MDIHK  183  ff
mksqun44         Undefined                         MDIHK  184  ff
mksqun54         Undefined                         MDIHK  185  ff
mksqun64         Undefined                         MDIHK  186  ff
mksqun74         Undefined                         MDIHK  187  ff
mksqrg0          Sequence Register 0               MDIHK  188  ff

mksqrg1          Sequence Register 1               MDIHK  190  ff
mksqrg2          Sequence Register 2               MDIHK  192  ff
mksqrg3          Sequence Register 3               MDIHK  194  ff
mksqrg4          Sequence Register 4               MDIHK  196  ff
mksqrg5          Sequence Register 5               MDIHK  198  ff
mksqrg6          Sequence Register 6               MDIHK  200  ff
mksqrg7          Sequence Register 7               MDIHK  202  ff
mksqrg8          Sequence Register 8               MDIHK  204  ff
mksqrg9          Sequence Register 9               MDIHK  206  ff
mksqrga          Sequence Register A               MDIHK  208  ff
mksqrgb          Sequence Register B               MDIHK  210  ff
mksqrgc          Sequence Register C               MDIHK  212  ff
mksqrgd          Sequence Register D               MDIHK  214  ff
mksqrge          Sequence Register E               MDIHK  216  ff
mksqrgf          Sequence Register F               MDIHK  218  ff
mksqid           Sequence Identifier               MDIHK  220  ff
mksqnum          Sequence Number                   MDIHK  224  ff
mksqfrm          Frame Number in Sequence          MDIHK  226  ff
mksqref0         LOBT Reference 0                  MDIHK  228  ff
mksqref1         LOBT Reference 1                  MDIHK  229  ff
mksqref2         LOBT Reference 2                  MDIHK  230  ff
mksqref3         LOBT Reference 3                  MDIHK  231  ff
mksqref4         LOBT Reference 4                  MDIHK  232  ff
mksqref5         LOBT Reference 5                  MDIHK  233  ff
mkdprstn         DEP Restart Number                MDIHK  234  ff
mkdpfrmn         Frame Number                      MDIHK  236  ff
mkdphkpn         Packet Number                     MDIHK  240  ff
mkdpswvr         DEP Software Version              MDIHK  244  ff
mkdpswpn         DEP Software Patch                MDIHK  245  ff
mkdpocvr         Optics Config Table Version       MDIHK  246  ff
mkdpolvr         OCT Lookup Table Version          MDIHK  247  ff
mkdpsqvr         DEP Sequence List Version         MDIHK  248  ff
mkdpsqpn         DEP Sequence List Patch           MDIHK  249  ff
mkipfwvr         IP Firmware Version               MDIHK  250  ff
mkipfwpn         IP Firmware Patch                 MDIHK  251  ff
mkipiqvr         IP Instruction Queue Version      MDIHK  252  ff
mkipiqpn         IP Instruction Patch              MDIHK  253  ff
mksqam1          AM1 Position and Status           MDIHK  254  ff
mksqam2          AM2 Position and Status           MDIHK  256  ff
mksqc1           C1 Position                       MDIHK  258  ff
mksqc2           C2 Position                       MDIHK  259  ff
mksqmsta         Misc Status (= byte 21)           MDIHK  260  ff

mksqltxy         LT Errors (= bytes 73-76)         MDIHK  261  ff
mksqltst         Limb Tarcker Status               MDIHK  265  ff
mksqhtrs         Heater Status (=bytes 24 - 27)    MDIHK  266  ff
mtsqopt1         Front Window Temp                 MDIHK  270  ff
mtsqopt4         Oven Temperature                  MDIHK  271  ff
mtsqopt5         Oven Temperature                  MDIHK  272  ff
mtsqopt6         CCD Temp                          MDIHK  273  ff
mkiptid1         IP Table ID #1                    MDIHK  274  ff
mkiptvr1         IP Table Version #1               MDIHK  276  ff
mkiptal1         IP Table Low address #1           MDIHK  278  ff
mkiptah1         IP Table High Address #1          MDIHK  280  ff
mkiptne1         IP Table Number of Entries #1     MDIHK  282  ff
mkiptid2         IP Table ID #2                    MDIHK  284  ff
mkiptvr2         IP Table Version #2               MDIHK  286  ff
mkiptal2         IP Table Low address #2           MDIHK  288  ff
mkiptah2         IP Table High Address #2          MDIHK  290  ff
mkiptne2         IP Table Number of Entries #2     MDIHK  292  ff
mkipdmps         IP Dump Error Status              MDIHK  294  ff
mkipdmp1         IP Dump Word 1                    MDIHK  296  ff
mkipdmp2         IP Dump Word 2                    MDIHK  298  ff
mkipdmp3         IP Dump Word 3                    MDIHK  300  ff
mkipdmp4         IP Dump Word 4                    MDIHK  302  ff
mkipdmp5         IP Dump Word 5                    MDIHK  304  ff
mkipecnt         IP Error Count                    MDIHK  306  ff
mkdpun1          Undefined DEP Status Byte         MDIHK  307  ff
mkdpun2          Undefined DEP Status Byte         MDIHK  308  ff
mkdpspec         DEP Special Error Code            MDIHK  309  ff
mkdpspe1         DEP Special Error Byte 1          MDIHK  310  ff
mkdpspe2         DEP Special Error Byte 2          MDIHK  311  ff
mkdpspe3         DEP Special Error Byte 3          MDIHK  312  ff
mkdpspe4         DEP Special Error Byte 4          MDIHK  313  ff
mkdpamsk         AXIS Mask                         MDIHK  314  ff
mkdppxst         Picture State Machine Status      MDIHK  316  ff
mkipstat         Image Processor Status Byte       MDIHK  317  ff
mkiperst         Image Processor Error Status      MDIHK  318  ff
mkipbank         Image Processor Bank              MDIHK  320  ff
mkdpmd           DEP Mode                          MDIHK  321  ff
mkdpl9st         DEP Level 9 Status                MDIHK  322  ff
mkdpl9er         DEP Level 9 Error Bits            MDIHK  324  ff
mkdpsqeb         DEP Sequence Error Bits           MDIHK  326  ff
mkdpsqst         DEP Sequence Status               MDIHK  328  ff
mkdpsqpt         DEP Sequence Pointer              MDIHK  330  ff

mkdpltst         DEP Limb Tracker State            MDIHK  332  ff
mkdpun3          Undefined DEP Status Byte         MDIHK  333  ff
mkdpdfsw         Opto Coupler Reg Default          MDIHK  334  ff
mkdpflag         DEP Misc Control Bits             MDIHK  336  ff
mspc1n           Primary Converter 1 Status (N)     SCHK   12  20
mspc1r           Primary Converter 1 Status (R)     SCHK   12  20
msrc1n           Redundant Converter 1 Status (N)   SCHK   12  40
msrc1r           Redundant Converter 1 Status (R)   SCHK   12  40
mspc2n           Primary Converter 2 Status (N)     SCHK   12  80
mspc2r           Primary Converter 2 Status (R)     SCHK   12  80
msrc2n           Redundant Converter 2 Status (N)   SCHK   13  20
msrc2r           Redundant Converter 2 Status (R)   SCHK   13  20
mspwdpn          DEP Power Status (N)               SCHK   13  40
mspwdpr          DEP Power Status (R)               SCHK   13  40
mspwaen          App Electronics Power Status (N)   SCHK   13  80
mspwaer          App Electronics Power Status (N)   SCHK   13  80
mspwipn          IP Power Status (N)                SCHK   12  01
mspwipr          IP Power Status (R)                SCHK   12  01
mspwdsn          DSOS Power Status (N)              SCHK   12  02
mspwdsr          DSOS Power Status (R)              SCHK   12  02
mstmseln         TLM Interface Select (N)           SCHK   12  04
mstmselr         TLM Interface Select (R)           SCHK   12  04
msfdopn          Front Door Open (N)                SCHK   12  08
msfdopr          Front Door Open (R)                SCHK   12  08
msfdcln          Front Door Closed (N)              SCHK   13  01
msfdclr          Front Door Closed (R)              SCHK   13  01
msltaokn         LT PZT A In Range (N)              SCHK   13  02
msltaokr         LT PZT A In Range (R)              SCHK   13  02
msdsokn          DSOS On-Line (N)                   SCHK   13  04
msdsokr          DSOS On-Line (R)                   SCHK   13  04
msltbokn         LT PZT B In Range (N)              SCHK   13  08
msltbokr         LT PZT B In Range (R)              SCHK   13  08
msltcokn         LT PZT C In Range (N)              SCHK   13  10
msltcokr         LT PZT C In Range (R)              SCHK   13  10
m+5pc1n          Converter 1 +5 Volts               SCHK    1  ff
m+5pc1r          Converter 1 +5 Volts               SCHK    1  ff
m+15pc1n         Converter 1 +15 Volts              SCHK    2  ff
m+15pc1r         Converter 1 +15 Volts              SCHK    2  ff
m-15pc1n         Converter 1 -15 Volts              SCHK    3  ff
m-15pc1r         Converter 1 -15 Volts              SCHK    3  ff
m+5pc2n          Converter 2 +5 Volts               SCHK    4  ff
m+5pc2r          Converter 2 +5 Volts               SCHK    4  ff

m+15pc2n         Converter 2 +15 Volts              SCHK    5  ff
m+15pc2r         Converter 2 +15 Volts              SCHK    5  ff
m-15pc2n         Converter 2 -15 Volts              SCHK    6  ff
m-15pc2r         Converter 2 -15 Volts              SCHK    6  ff
qtopts1n         OP Radiator                        SCHK    7  ff
qtopts1r         OP Radiator                        SCHK    7  ff
qtopts2n         OP Structure                       SCHK    8  ff
qtopts2r         OP Structure                       SCHK    8  ff
qtopts3n         Filter Oven                        SCHK    9  ff
qtopts3r         Filter Oven                        SCHK    9  ff
qtepts1n         Primary Converter 1                SCHK   10  ff
qtepts1r         Primary Converter 1                SCHK   10  ff
qtopts4n         CCD Radiator                       SCHK   11  ff
qtopts4r         CCD Radiator                       SCHK   11  ff

54 D LISTING OF FITS FIELDS

D Listing of FITS Fields

The following is a copy of the document ∼soi/doc/TechNotes/92.085 which describes the fields in the FITS headers.

SOI FITS Keyword List
R. S. Bogart & J. Suryanarayanan
1992.IX.10
rev. 1992.IX.24
rev. 1994 Jun 7 JKS
SOI Technical Note 92-085
This is a prototype list of standard FITS keywords to be used in the
SOI project.  The list is not intended to be exclusive, but represents
the full set of project reserved keywords.
Designations for value indicators are as follows:
char:   character strings delimited by single quotes as described in
        the FITS standard 5.3.1.
bool:   characters T or F as described in 5.3.2.
int:    integer representations as described in 5.3.3.
real:   floating point representations as described in 5.3.4.  The standard
        is not clear on whether the ASCII character string "NaN" is an
        acceptable representation for a floating point NaN.  We assume it is.
1.  Keywords reserved by the FITS standard
[blank]
        Any number may appear.  Column 9 may not contain "=".
AUTHOR  = char
BITPIX  = int
        Mandatory in all headers.  Must be second keyword in header.
BLANK   = int
        May be used only if BITPIX value is positive (integer data).
BLOCKED = bool
        If present, value must apparently be T.  Deprecated.
BSCALE  = real
BUNIT   = char
BZERO   = real
CDELTn  = real          [n = 1, 999]
COMMENT
        No value.  Any number may appear.  Column 9 may not contain "=".
CONVENTN= string
        Convention used to describe how the binary tables are used to store data.
        This keyword will appear in the binary table extension header. The value
        of this keyword will be TDIM.
CRPIXn  = real          [n = 1, 999]
CROTAn  = real          [n = 1, 999]
CRVALn  = real          [n = 1, 999]
CTYPEn  = char          [n = 1, 999]
DATAMAX = real
DATAMIN = real
DATE    = 'DD/MM/YY'
        Date on which the Header Data Unit was created.  DD, MM, and YY are
        two-digit integers (with leading 0's) giving the day of the month,
        month number, and year mod 100, respectively.
DATE-OBS= 'DD/MM/YY'
        UT date on which the observations were made.  Values as described

under DATE keyword.
END
Mandatory in all headers.  Must be last keyword in header.
EPOCH   = char
Deprecated
EQUINOX = char
EXTEND  = bool
Mandatory in primary header if data set contains extensions, in
which case it must immediately follow the last NAXISn keyword card
image (or NAXIS if NAXIS = 0).
EXTLEVEL= int
EXTNAME = char
EXTVER  = int
GCOUNT  = int
Mandatory in extension headers, in which case it must immediately
follow the PCOUNT keyword.
GROUPS  = bool
HISTORY
No value.  Any number may appear.  Column 9 may not contain "=".
INSTRUME= char
'MDI     ' and 'Michelson Doppler Imager' are appropriate values.
NAXIS   = int
Mandatory in all headers.  Must be third keyword in header.
NAXISn  = int           [n = 1, 999]
Mandatory in all headers up to n = NAXIS of the header.  Must follow
NAXIS keyword in order (NAXIS1, NAXIS2, ...) unless NAXIS = 0.  It is
not clear whether additional NAXISn keywords may be present, but the
implication is that the number of NAXISn keywords should be exactly
equal to the value of NAXIS.
OBJECT  = char
OBSERVER= char
ORIGIN  = char
PCOUNT  = int
Mandatory in extension headers.
PSCALn  = real          [n = 1, 999]
PTYPEn  = char          [n = 1, 999]
PZEROn  = real          [n = 1, 999]
REFERENC= char
SIMPLE  = bool
Mandatory in primary header; not permitted in extension headers.
The value must be false if the file does not conform to the
FITS standard.  Otherwise it should be true.  Must be first keyword
in header.
TBCOLn  = int           [n = 1, 999]
Mandatory in ASCII Tables Extension headers, in which case the number
of these keywords must be equal to the value of TFIELDS and they must
appear in order, analogously to NAXISn.
TDELTn  = string
Extension of CDELTn keywords to binary tables.
TDESCn  = string
Extension of CTYPEn keywords to binary tables. Has the format
TDESCn='(desc1,desc2,. .).' where each element in the list corresponds
to one of the dimensions in the associated TDIMnnn keyword.
TDIMn   = string    [n = 1, 999]
The value is a character string containing the dimensions of a
multidimensional array stored in column nnn in a binary table, with
the format TDIMn='(dim1,dim2,. .).
TDISn   = ???       [n = 1, 999]
TDMINn  = real          [n = 1, 999]
Equivalent of DATAMIN keyword for a column in an ASCII or binary table.

TDMAXn  = real          [n = 1, 999]
Equivalent of DATAMAX keyword for a column in an ASCII or binary table.
TELESCOP= char
TFIELDS = int
Mandatory in ASCII Tables Extension headers, in which case it must
immediately follow the GCOUNT keyword.
TFORMn  = char          [n = 1, 999]
Mandatory in ASCII Tables Extension headers, in which case the number
of these keywords must be equal to the value of TFIELDS and they must
appear in order, analogously to NAXISn.  The first TFORMn keyword in
an ASCII Tables Extension header must immediately follow the last
TBCOLn keyword.
THEAP   = ???
TNULLn  = char          [n = 1, 999]
Reserved only for ASCII Tables extension data.  Should be treated as
reserved.
TROTAn  = char
Extension of CROTAn keywords to binary tables. Has the format
TROTAn='(rota1,rota2,. .).' where each element in the list corresponds to
one of the dimensions in the associated TDIMnnn keyword.
TRPIXn  = char
Extension of CRPIXn keywords to binary tables.
TRVALn  = char
Extension of CRVALn keywords to binary tables.
TSCALn  = real          [n = 1, 999]
Reserved only for ASCII Tables extension data.  Should be treated as
reserved.
TTYPEAn = char          [n = 1, 999]
Reserved only for ASCII Tables extension data.  Should be treated as
reserved.
TUNITn  = char          [n = 1, 999]
Reserved only for ASCII Tables extension data.  Should be treated as
reserved.
TZEROn  = real          [n = 1, 999]
Reserved only for ASCII Tables extension data.  Should be treated as
reserved.
XTENSION= char
Mandatory in extension header; not permitted in primary headers.
Must be first keyword in header.
2.  Keywords specifically associated with SOHO
ANGLE   = real
Angle of rotation of the vertical axis of the instrument field-of-view
relative to solar north.
AIT_END = real
On-board value for the time of the end of data acquisition.
AIT_TIME= real
On-board value for the starting time of data acquisition.
CMP_NAME= character
Name of observing campaign.
CMP_NO  = integer
Unique identifier for an observing campaign.
COMMENTS= character
Either \Y" or \N", depending on whether or not an entry in a database is
commented or not. For use in the online catalogs only.
COMM_NO = integer
Comment number.  For use in the online catalogs only.
COUNT   = integer
Number of repeated observing sequences within an observing program.

DATE_END= char
Ending date of data acquisition, in UT. Shall be in the same format
as DATE_OBS.
DATE_MOD= date
Date that an entry in a catalog was last modified. For use in the online
catalogs only.
DATE_OBS= char
Starting date of data acquisition, in UT. Shall be in the format
YYYY/MM/DD (Note the distinction between this and the standard FITS keyword     DATE-OBS. The standard
an underscore.)
DETECTOR= char
Detector used to acquire the data.
EXPCOUNT= integer
Number of exposures within an observing sequence.
EXP_IND = integer
Index of an exposure within an observing sequence, from one to the value
of EXPCOUNT.
EXPTIME = real
Exposure time in seconds, to millisecond accuracy.
FILENAME= char
Name of data file. The use of this keyword will not be construed to mean
that the FITS header will need to be changed if the file is copied from
one computer to another, or to another media such as tape.
FILEORIG= char
Original filename as collected.
HEIGHT  = real
Maximum height of the instrument field-of-view in the direction along the
vertical axis as used in keyword ANGLE.
INS_ROLL= real
Instrument roll angle relative to the spacecraft, i.e.  the angle between
the instrument Y-axis and the spacecraft Y-axis, measured in a
counter-clockwise direction from the spacecraft Y-axis toward the
spacecraft Z-axis.
INSTITUT= character
Name of an intrument team or institution taking part in an observing
campaign.
INS_Y   = real
Pointing of instrument along the instrument Y-axis.
INS_Y0  = real
Pointing of instrument origin along the spacecraft Y-axis.
INS_Z   = real
Pointing of instrument along the instrument Z-axis.
INS_Z0  = real
Pointing of instrument origin along the spacecraft Z-axis.
OBJ_ID  = char
Object identifier, e.g. active region number.
OBS_MODE= char
Observing mode.
OBS_PROG= char
The observing program.
OBS_SEQ = char
The name of the observing sequence used.
PROG_ID = integer
Observing program identification number. PROG_ID is used to asociate
individual observations together as being part of the same observing
program. In particular, an observing program that is not contiguous
in time, for instance an observation made at the same time each day,
or once every solar rotation, could be identified by the same value
of PROG_ID.
PROG_NUM= integer

Unique identifier for the observing program. Each observing program is
assigned a unique identifier number each time it's run, even if it's a
continuation of an earlier program.
SC_ROLL = real
Spacecraft roll angle relative to solar coordinates.
SC_Y0   = real
Spacecraft pointing along the solar Y-axis as defined in the Science
Operations Plan.
SC_Z0   = real
Spacecraft pointing along the solar Z-axis as defined in the Science
Operations Plan.
SCI_OBJ = char
The science objective.
SEQ_FROM= integer
First value of SEQ_NUM in a series of observing sequences.
SEQ_IND = integer
Index of an observing sequence within an observing program, from one to
the value of COUNT.
SEQ_NUM = integer
Unique number assigned to an execution of an observing sequence.
Depending on how an instrument team organizes their data, this number
may represent an individual observing sequence or a repeated series of
the same observing sequence.
SEQONLIN= character
Either \Y" or \N", denoting whether or not the data from an observing
sequence is online or not.
SEQ_TO  = integer
Last value of SEQ_NUM in a series of observing sequences.
SEQVALID= character
Either \Y" or \N", denoting whether or not the data from an observing
sequence has scientific value.
SFDUADID= char
Standard Formatted Data Unit (SFDU) Authority and Description Identifier
(ADID) for a data object.  Corresponds to bytes 1{4 and 9{12 of the SFDU
label associated with the data object (file). Associating this keyword
with the data maintains this information even if the SFDU labels are
separated from it.
SOLAR_P0= real
The solar P0 angle.
SOLAR_B0= real
The solar B0 angle.
TIME_END= char
Ending time of data acquisition, in UT. Shall be in the same format as
TIME_OBS.
TIME_MOD= time
Time that an entry in a catalog was last modified.  For use in the online
catalogs only.
TIME_OBS= char
Starting time of data acquisition, in UT. Shall be in the format
HH:MM:SS.SSS
WAVELNTH= real
The  wavelength  of  observation,  in  nm.    Also, TWAVEnnn for binary
tables.  When used for observations that cover a range of wavelengths,
then this keyword represents the wavelength of interest, not necessarily
the central wavelength.
WAVEMIN = real
Minimum wavelength of observation, in nm.  Also, TWMIN for binary tables.
WAVEMAX = real
Maximum wavelength of observation, in nm.  Also, TWMAX for binary tables.
WIDTH   = real

        Maximum width of the instrument field-of-view in the direction
        perpendicular to the vertical axis as used in keyword ANGLE.
YCEN    = real
        Center of the instrument field-of-view along the solar Y-axis.
ZCEN    = real
        Center of the instrument field-of-view along the solar Z-axis.
3.  Keywords corresponding to MDI housekeeping mnemonics
Note: A few of Lockheed's proposed Keywords had to be changed to confirm to
FITS standard and some others were changed for consistency.  For that purpose
"+" in the keyword name has been changed to "VP" (Voltage Plus) and "-" in the
keyword name has been changed to "VM" (Volt Minus).
MFLTXOFS
        Limb Tracker X offset
MFLTYOFS
        Limb Tracker Y offset; differs from proposed mnemonic MFLTXOFS which
        is not unique
MFLTAOS
        Limb Tracker A offset
MFLTBOS
        Limb Tracker B offset
MFLTCOS
        Limb Tracker C offset
MFLTSTAT
        Limb Tracker Status Register
MFLTXGN
        Limb Tracker X gain
MFLTYGN
        Limb Tracker Y gain
MIAEMTR
        Motor Current Monitor
MKC1POS = int
        Calibration/focus wheel 1
        Range: 0 - 143
MKC2POS = int
        Calibration/focus wheel 2
        Range: 0 - 143
MKDPMERR
        Correctable DEP memory errors
MKDPPATN
        DEP Patch Number
MKDPSWVR
        DEP Software Version
MKFDMTR1
        Door Motor 1 Status
MKFDMTR2
        Door Motor 2 Status
MKFDOPN1
        Door Open Status 1
MKFDOPN2
        Door Open Status 2
MKHOPH2
        Optics Package Heater 2 On/Off
MKHOPH3
        Optics Package Heater 3 On/Off
MKHOPSPT
        Optics Package Heater Set Point
MKHPOMON

Prime Oven Monitoring
MKHPOMTW
Prime Oven Monitoring Time
MKHPOSPT
Prime Oven Set Point
MKHROCBH
Backup Oven Control Band High
MKHROCBL
Backup Oven Control Band Low
MKHROCPT
Backup Oven Control Point
MKHRODBH
Backup Oven Dead Band High
MKHROMON
Backup Oven Monitoring
MKHRODBL
Backup Oven Dead Band Low
MKHROSPT
Backup Oven Set Point
MKIPFLAG
IP Flags
MKIPMAC
IP Macro Load Number
MKIPMASK
IP Mask Set
MKIPMERR
Correctable IP memory errors
MKIPPATN
IP Patch Number
MKIPSWVR
IP Firmware Version
MKLTLOOP
ISS Loop Open/Closed
MKM1DIR
MTM1 Default Direction
MKM2DIR
MTM2 Default Direction
MKM1POS = int
Michelson Tuning Motor 1
Range: 0 - 179
MKM2POS = int
Michelson Tuning Motor 2
Range: 0 - 179
MKOVNMAX
Filter Oven Max Temp
MKPADIR
PAW Default Direction
MKPAPOS = int
Polarization Analyzer Wheel
Range: 0 - 143
MKPKFMID
Packet Format ID
MKPWAE
Application I/F Power
MKPWAM
Alignment  Power
MKPWC1
Cal 1 Power
MKPWC2
Cal 2 Power

MKPWCM
        CCD Camera Power
MKPWDM
        DMC Power
MKPWFD
        Door Power
MKPWHCM
        CCD Heater Power
MKPWHOP
        Prime Oven Heater Power
MKPWHPO
        Optics Package heater
MKPWHRO
        Red. Oven Heater Pwr
MKPWIP
        Image Processor Power
MKPWLT
        Limb Tracker Power
MKPWM1
        MTM 1 Power
MKPWM2
        MTM 2 Power
MKPWPA
        PAW Power
MKPWSH
        Shutter Power
MKPWTM
        IP TLM 1-2 Power
MKSHNEXP
        Shutter Nominal Exposure
MKSHPATH= char
        Shutter Default Beam Path
        'DARK    ', 'FULL DISK', 'HIREZ   ', and 'UNDEF   ' are valid values.
MKSQAPA1
        APU Address Frame 1
MKSQAPA2
        APU Address Frame 2
MKSQAPA3
        APU Address Frame 3
MKSQAPA4
        APU Address Frame 4
MKSQC1A1
        CAL1 Actual Position Frame 1
MKSQC1A2
        CAL1 Actual Position Frame 2
MKSQC1A3
        CAL1 Actual Position Frame 3
MKSQC1A4
        CAL1 Actual Position Frame 4
MKSQC2A1
        CAL2 Actual Position Frame 1
MKSQC2A2
        CAL2 Actual Position Frame 2
MKSQC2A3
        CAL2 Actual Position Frame 3
MKSQC2A4
        CAL2 Actual Position Frame 4
MKSQCA
        Sequence Current Address
MKSQDM11

DMA Control Word 1 Frame 1
MKSQDM12
DMA Control Word 1 Frame 2
MKSQDM13
DMA Control Word 9 Frame 3; should probably be Word 1
MKSQDM14
DMA Control Word 1 Frame 4
MKSQDM21
DMA Control Word 2 Frame 1
MKSQDM22
DMA Control Word 2 Frame 2
MKSQDM23
DMA Control Word 2 Frame 3
MKSQDM24
DMA Control Word 2 Frame 4
MKSQDM31
DMA Control Word 3 Frame 1
MKSQDM32
DMA Control Word 3 Frame 2
MKSQDM33
DMA Control Word 3 Frame 3
MKSQDM34
DMA Control Word 3 Frame 4
MKSQDM41
DMA Control Word 4 Frame 1
MKSQDM42
DMA Control Word 4 Frame 2
MKSQDM43
DMA Control Word 4 Frame 3
MKSQDM44
DMA Control Word 4 Frame 4
MKSQIP11
IP Parameter 1 Frame 1
MKSQIP12
IP Parameter 1 Frame 2
MKSQIP13
IP Parameter 1 Frame 3
MKSQIP14
IP Parameter 1 Frame 4
MKSQIP21
IP Parameter 2 Frame 1
MKSQIP22
IP Parameter 2 Frame 2
MKSQIP23
IP Parameter 2 Frame 3
MKSQIP24
IP Parameter 2 Frame 4
MKSQIP31
IP Parameter 3 Frame 1
MKSQIP32
IP Parameter 3 Frame 2
MKSQIP33
IP Parameter 3 Frame 3
MKSQIP34
IP Parameter 3 Frame 4
MKSQIP41
IP Parameter 4 Frame 1
MKSQIP42
IP Parameter 4 Frame 2
MKSQIP43

IP Parameter 4 Frame 3
MKSQIP44
IP Parameter 4 Frame 4
MKSQIP51
IP Parameter 5 Frame 1
MKSQIP52
IP Parameter 5 Frame 2
MKSQIP53
IP Parameter 5 Frame 3
MKSQIP54
IP Parameter 5 Frame 4
MKSQIP61
IP Parameter 6 Frame 1
MKSQIP62
IP Parameter 6 Frame 2
MKSQIP63
IP Parameter 6 Frame 3
MKSQIP64
IP Parameter 6 Frame 4
MKSQIP71
IP Parameter 7 Frame 1
MKSQIP72
IP Parameter 7 Frame 2
MKSQIP73
IP Parameter 7 Frame 3
MKSQIP74
IP Parameter 7 Frame 4
MKSQIP81
IP Parameter 8 Frame 1
MKSQIP82
IP Parameter 8 Frame 2
MKSQIP83
IP Parameter 8 Frame 3
MKSQIP84
IP Parameter 8 Frame 4
MKSQIPC1
IP Control Frame 1
MKSQIPC2
IP Control Frame 2
MKSQIPC3
IP Control Frame 3
MKSQIPC4
IP Control Frame 4
MKSQLBA
Sequence Last Branch Address
MKSQM1A1
MTM1 Actual Position Frame 1
MKSQM1A2
MTM1 Actual Position Frame 2
MKSQM1A3
MTM9 Actual Position Frame 3; should probably be MTM1
MKSQM1A4
MTM1 Actual Position Frame 4
MKSQM1P1
MTM1 Planned Position Frame 1
MKSQM1P2
MTM1 Planned Position Frame 2
MKSQM1P3
MTM9 Planned Position Frame 3; should probably be MTM1
MKSQM1P4

MTM1 Planned Position Frame 4
MKSQM2A1
MTM2 Actual Position Frame 1
MKSQM2A2
MTM2 Actual Position Frame 2
MKSQM2A3
MTM2 Actual Position Frame 3
MKSQM2A4
MTM2 Actual Position Frame 4
MKSQM2P1
MTM2 Planned Position Frame 1
MKSQM2P2
MTM2 Planned Position Frame 2
MKSQM2P3
MTM2 Planned Position Frame 3
MKSQM2P4
MTM2 Planned Position Frame 4
MKSQNFRM
Number of Frames
MKSQNK1
Compression Control Frame 1
MKSQNK2
Compression Control Frame 2
MKSQNK3
Compression Control Frame 3
MKSQNK4
Compression Control Frame 4
MKSQOCC1
Optics Config. Control Frame 1
MKSQOCC2
Optics Config. Control Frame 2
MKSQOCC3
Optics Config. Control Frame 3
MKSQOCC4
Optics Config. Control Frame 4
MKSQPAA1
PAW Actual Position Frame 1
MKSQPAA2
PAW Actual Position Frame 2
MKSQPAA3
PAW Actual Position Frame 3
MKSQPAA4
PA Actual Position Frame 4
MKSQPAP1
PAW Planned Position Frame 1
MKSQPAP2
PAW Planned Position Frame 2
MKSQPAP3
PAW Planned Position Frame 3
MKSQPAP4
PAW Planed Position Frame 4
MKSQREF0
LOBT Reference 0
MKSQREF1
LOBT Reference 1
MKSQREF2
LOBT Reference 2
MKSQREF3
LOBT Reference 3
MKSQREF4

LOBT Reference 4
MKSQREF5
LOBT Reference 5
MKSQRG0
Sequence Register 0
MKSQRG1
Sequence Register 1
MKSQRG2
Sequence Register 2
MKSQRG3
Sequence Register 3
MKSQRG4
Sequence Register 4
MKSQRG5
Sequence Register 5
MKSQRG6
Sequence Register 6
MKSQRG7
Sequence Register 7
MKSQRG8
Sequence Register 8
MKSQRG9
Sequence Register 9
MKSQRGA
Sequence Register A
MKSQRGB
Sequence Register B
MKSQRGC
Sequence Register C
MKSQRGD
Sequence Register D
MKSQRGE
Sequence Register E
MKSQRGF
Sequence Register F
MKSQRT1
Relative Time Frame 1
MKSQRT2
Relative Time Frame 2
MKSQRT3
Relative Time Frame 3
MKSQRT4
Relative Time Frame 4
MKSQSA
Sequence Start Address
MKSQSHC1
Shutter Close Time Frame 1
MKSQSHC2
Shutter Close Time Frame 2
MKSQSHC3
Shutter Close Time Frame 3
MKSQSHC4
Shutter Close Time Frame 4
MKSQSHO1
Shutter Open Time Frame 1
MKSQSHO2
Shutter Open Time Frame 2
MKSQSHO3
Shutter Open Time Frame 3
MKSQSHO4

Shutter Open Time Frame 4
MKSQSNUM
Sequence Serial Number
MKSQTMC1
Telemetry Control Frame 1
MKSQTMC2
Telemetry Control Frame 2
MKSQTMC3
Telemetry Control Frame 3
MKSQTMC4
Telemetry Control Frame 4
MKTCCNT
Command counter
MKTCLNVC
MLA of Latest NVCB
MKTCNVCB
Non-valid Command Blocks
MKTCNVER
Error code of latest NVCB
MKTCVCB
Valid Command Blocks
MTEPTS1
Electronics Temp 1  (PRT)
MTEPTS2
Electronics Temp 2  (Thermistor)
MTEPTS3
Electronics Temp 3  (Thermistor)
MTEPTS4
Electronics Temp 4  (PRT)
MTOPTS1
Optics Temp 1  (Thermistor)
MTOPTS2
Optics Temp 2  (Thermistor)
MTOPTS3
Optics Temp 3  (Thermistor)
MTOPTS4
Optics Temp 4  (Thermistor)
MTOPTS5
Optics Temp 5  (Thermistor)
MTOPTS6
Optics Temp 6  (PRT)
MTOPTS7
Optics Temp 7  (Thermistor)
MTOPTS8
Optics Temp 8  (PRT)
MTOPTS9
Optics Temp 9  (Thermistor)
MTOPTS10
Optics Temp 10  (Thermistor)
MTOPTS11
Optics Temp 11  (PRT)
MTOPTS12
Optics Temp 12  (Thermistor)
MTOPTS13
Optics Temp 13  (Thermistor)
MTOPTS14
Optics Temp 14  (Thermistor)
MTOPTS15
Optics Temp 15  (Thermistor)
MTOPTS16

Optics Temp 16  (Thermistor)
MUHOPCTL
Optics package Heater Controller
MUHPOBRG
Filter Oven Bridge Out
MUHPOLOP
Filter Oven Loop Out
MUHROCTL
Filter Oven Redundant Controller
MULTDIO1
Limb Tracker Photodiode 1
MULTDIO2
Limb Tracker Photodiode 2
MULTDIO3
Limb Tracker Photodiode 3
MULTDIO4
Limb Tracker Photodiode 4
MULTPZT1
Limb Tracker PZT1
MULTPZT2
Limb Tracker PZT2
MULTPZT3
Limb Tracker PZT3
MULTXAVE
Limb Tracker X Average Error
MULTYAVE
Limb Tracker Y AverageError
MULTXMAX
Limb Tracker X Max Error
MULTYMAX
Limb Tracker Y Max Error
MVM15AE
App. Elec  -15 volts; Cnv2; differs from proposed mnemonic M-15AE for
consistency with positive voltage keywords
MVM15CM
CCD Camera -15 volts; Cnv1; differs from proposed mnemonic M-15CM for
consistency with positive voltage keywords
MVP05AE
App. Elec  +5 volts; Cnv2; differs from proposed mnemonic M+5AE which
violates FITS standard
MVP05CM
CCD Camera +5 volts; Cnv1; differs from proposed mnemonic M+5CM which
violates FITS standard
MVP05IP
IP AND TLM +5 volts; Cnv1; differs from proposed mnemonic M+5IP which
violates FITS standard
MVP15AE
App. Elec  +15 volts; Cnv2; differs from proposed mnemonic M+15AE which
violates FITS standard
MVP15CM
CCD Camera +15 volts; Cnv1; differs from proposed mnemonic M+15CM which
violates FITS standard
MVP15HTR
Heater +15 volts; Cnv2; differs from proposed mnemonic M+15HTR which
violates FITS standard
MVP15MTR
Motor +15 volts; Cnv2; differs from proposed mnemonic M+15MTR which
violates FITS standard
MVP30CM
CCD Camera +30 volts; Cnv1; differs from proposed mnemonic M+30CM which

violates FITS standard
MVP75LT
Limb Tracker +75 volts; Cnv2; differs from proposed mnemonic M+75LT
which violates FITS standard
MXAM1POS
Leg Potentiometer Wiper 1
MXAM2POS
Leg Potentiometer Wiper 2
MXHPOREF
Temperature Sensor Reference
3.  Other SOI reserved keywords
A.  General
MISSING = real
B.  Ephemerides
HGDCLAT = real
Heliographic Latitude of observed Solar Disc Center
HGDCLON = real
Heliographic Longitude of observed Solar Disc Center
SUNRAD  = real
Apparent semidiameter of Solar Disc
SUNDIST =real
Distance from Observer to center of Sun
TIME-OBS= 'HH:MM:SS.XXX'
UT time of observation.  Values as described under TIME keyword
C.  Creation Parameters
TIME    = 'HH:MM:SS.XXX ZZZ'
Time at which the Header Data Unit was created.  HH and MM are
two-digit integers (with leading 0's) giving the hour and minute,
respectively.  SS.XXX, if present, is either a real or two digit
integer (with leading 0's) representing seconds.  ZZZ, if present,
is a character string representing time zone information.  If it is
not present, UT (Universal Time) is assumed.
4.  FITS Keywords for MDI PF Testing
The following is a list of proposed keywords for images taken by the MDI
breadboard camera (the parentheses indicate the corresponding f0 structure).
Not included are FITS keywords listed in Sections 1 and 2 above.
BLS_PIX = int
Number of BLS pixels (bls)
CENTER_X= int
x-Coordinate of image center in millipixels used in gain calculation
CENTER_Y= int
y-Coordinate of image center in millipixels used in gain calculation
RADIUS  = int
Radius of image in pixels used in gain calculation
IMAGENO = int
Image serial number (imnum)
EXT_COLS= int
Number of extended pixels (eper)
EXT_ROWS= int
Number of extended rows (eler)

FRAME   = char
        Frame position descriptor (frame)
FTYPE   = char
        Datatype descriptor (ftype)
SEQUENCE= char
        sequencename, incl. serial number (seq)
SITE    = char
        Observing site (site)
SUMMING = char
        Summing mode (sum)
TEMPn   = int           n = [0, 9]
        Temperature n from camera string (temp(n))
VIDEODSK= bool
        flag if image on videodisk (vdsk)

70E LISTING OF HOW MANY TIMES SOME CONTROL FILES WERE RUN (HIST COUNT OUTPUT)

E Listing of How Many Times Some Control Files were Run (HIST COUNT output)

HIST_COUNT  Ver 1.1    Program Run: 24-May-1994 18:55:08.00
Input file spans: 30-JUN-93 19:00:04 to 12-MAY-94 18:09:48
                 9306  9307  9308  9309  9310  9311  9312  9401  9402  9403  9404  9405   Total
mdipict             0     0     0     1     3     0    10     3     0     5    11     2      35
mdiltc              0     0     3     0     2    13    13     3     4    10    13     2      63
mdifocus            0     0     0     1     2     3    10     3     1     3     8     9      40
mdiiss              0     0     0     0     3     8     5     5     0     0     2     1      24
mdidust             0     0     0     0     3     0     0     0     0     0     0     0       3
mdilaser            0     0     0     0     0     0     1     2     0     5     5     1      14
mdidtune            0     0     0     0    29     6     4     0     0     0     0     0      39
big                 0    13    26     0    35    23     3     0     0     3     3     0     106
supersun            0     0     0     0     2     0     0     0     0     0     0     0       2
vfdlscan            0     3     7     0    12    15     0     0     4     4    22     0      67
mtmtune             0     0     0     0     0    70     0     0     0     0     2     0      72
maplaser            0     0     0     0     0     3     0     0     0     0     0     0       3
mapgrad             0     0     0     0     0     4     0     0     0     0     3     0       7
mdiprot             0     0     0     1     3     0     1     1     0     0     2     2      10
mdimrot             0     0     0     1     2     0     1     1     0     0     3     1       9
mdipdist            0     0     0     3     2     0     1     0     0     0     2     1       9
" mdion "           0    11    20     2    28    38    55    41    22    44    53    10     324
lfdclose            0     0    17     2    11    18    39    23    30    25    40    10     215
mbfdbc              0    62    69    30    35    33    45    24    28    25    35    11     397
lfdopen             0     0     8     2    27    22    29    16    30    39    25     5     203
mbfd1o              3    52    55    18    35    38    43    24    30    25    36    12     371

F Listing of When Control Files were Run (HIST SUMMARY output)

NOTE: It is possible to search the calibration log CAL LOG to list all datasets which were created using a certain control file. This list would most likely be a subset of the command summary list below because the are occassions when execution of control files is aborted before completion. The listing from the CAL LOG would include the CCD and oven temperatures. A sample use of the alias calgrep is:

% calgrep mdipict

HIST_SUMMARY  Ver 1.2    Program Run: 24-May-1994 18:00:23.00
Searching for: mdipict,mdiltc,mdifocus,mdiiss,mdidust,mdilaser,shutest,mdicom,lcmcom,ipcom,ipmemtst,iptbltst,ipdmatst,ipdynmac,ipseqtst,lsqcom,mdidtune,bigtune,supersun,vfdlscan,mtmtune,maplaser,mapgrad,mdiprot,mdimrot,mdipdist
Input file spans: 30-JUN-93 19:00:04 to 12-MAY-94 18:09:48

------------------------------ Searching for mdipict ------------------------------
930901_061019 <  >>     mdipict                                1-SEP-93 06:10:30
931001_194438    >> mdipict                                    1-OCT-93 19:44:05
931008_010431    >> mdipict                                    8-OCT-93 01:04:03
931018_210208 <  >>         mdipict                           18-OCT-93 21:01:14
931218_011846    >>     mdipict                               18-DEC-93 01:17:56
931218_214831    >> mdipict                                   18-DEC-93 21:47:07
931220_064308 <  >> mdipict                                   20-DEC-93 06:41:52
931221_194038    >> mdipict                                   21-DEC-93 20:32:53
931221_210002    >> mdipict                                   21-DEC-93 20:59:13
931222_134717    >>                             mdipict       22-DEC-93 16:36:45
931222_163837    >>                             mdipict       22-DEC-93 16:37:21
931223_012013 <  >> mdipict                                   23-DEC-93 01:19:35
931223_055631 <  >> mdipict                                   23-DEC-93 05:55:36
931223_163917    >>     mdipict                               23-DEC-93 16:38:13
940112_224104 <  >> mdipict                                   12-JAN-94 22:40:24
940120_115542    >>     mdipict                               20-JAN-94 11:54:31
940120_121041 <  >>     mdipict                               20-JAN-94 12:09:48
940304_214902 <  >>     mdipict                                4-MAR-94 21:48:18
940304_223746 <  >>         mdipict                            4-MAR-94 22:35:47
940315_210948 <  >>     mdipict                               14-MAR-94 21:07:23
940315_210948    >>     mdipict                               14-MAR-94 21:08:50
940327_150141 <  >>     mdipict                               27-MAR-94 15:00:55
940401_143902 <  >>     mdipict                                1-APR-94 14:37:51
940406_141503 <  >>     mdipict                                6-APR-94 14:13:43
940410_140323    >> mdipict                                   10-APR-94 14:02:27
940422_203742    >> mdipict                                   22-APR-94 20:36:35
940423_021755    >>             mdipict                       23-APR-94 02:16:40
940423_043634    >>         mdipict                           23-APR-94 04:35:20
940424_204659 <  >>     mdipict                               24-APR-94 20:46:05
XXXXXX_XXXXXX    >> mdipict                                   26-APR-94 03:45:14
940426_035009    >>     mdipict                               26-APR-94 03:48:58
940426_035538 <  >>     mdipict                               26-APR-94 03:54:49
940426_044621 <  >>             mdipict                       26-APR-94 04:45:27
940504_095350    >> mdipict                                    4-MAY-94 09:52:42

72 F LISTING OF WHEN CONTROL FILES WERE RUN (HIST SUMMARY OUTPUT)

940506_120808    >> mdipict                                    6-MAY-94 12:07:02
------------------------------ Searching for mdiltc ------------------------------
XXXXXX_XXXXXX    >>     mdiltc                                20-AUG-93 01:51:31
XXXXXX_XXXXXX    >>     mdiltc                                20-AUG-93 01:54:48
XXXXXX_XXXXXX    >>     mdiltc                                20-AUG-93 02:06:57
931001_200012 <  >> mdiltc                                     1-OCT-93 19:59:46
931008_011902 <  >> mdiltc                                     8-OCT-93 01:18:33
931103_013705    >> mdiltc                                     3-NOV-93 02:41:54
931103_013705    >> mdiltcs2                                   3-NOV-93 02:57:29
931103_013705    >> mdiltcs2                                   3-NOV-93 03:04:49
931103_013705    >> mdiltcs2                                   3-NOV-93 03:13:26
931103_013705    >> mdiltcs2                                   3-NOV-93 03:20:48
931103_013705    >> mdiltcs2                                   3-NOV-93 03:32:14
931122_235018    >> mdiltc                                    22-NOV-93 23:49:30
931123_085214 <  >>     mdiltc                                23-NOV-93 08:51:53
931123_232756    >> mdiltc                                    23-NOV-93 23:27:15
931124_082454 <  >> mdiltc                                    24-NOV-93 08:23:51
931124_083139    >> mdiltc                                    24-NOV-93 08:31:10
931124_083549 <  >> mdiltc                                    24-NOV-93 08:35:20
931124_083856 <  >> mdiltc                                    24-NOV-93 08:38:26
931218_010148 <  >> mdiltc                                    18-DEC-93 01:00:53
931218_220431 <  >> mdiltc                                    18-DEC-93 22:03:34
931220_065852 <  >> mdiltc                                    20-DEC-93 06:58:01
931223_013528 <  >> mdiltc                                    23-DEC-93 01:34:37
931223_061445 <  >> mdiltc                                    23-DEC-93 06:13:36
931223_165515 <  >>     mdiltc                                23-DEC-93 16:54:14
931228_163045    >> mdiltcs1                                  28-DEC-93 16:29:22
931228_172003 <  >> mdiltcs2                                  28-DEC-93 17:18:14
931228_183645    >> mdiltcs4                                  28-DEC-93 18:35:41
931228_195938    >> mdiltcs4                                  28-DEC-93 19:58:55
931228_221532    >> mdiltcs1                                  28-DEC-93 22:14:44
931228_235936 <  >> mdiltcs2                                  28-DEC-93 23:59:03
931229_134517 <  >> mdiltcs2                                  29-DEC-93 13:44:39
940112_230148 <  >> mdiltc                                    12-JAN-94 22:59:46
940119_163231    >> mdiltc                                    19-JAN-94 18:00:39
940119_180239    >>     mdiltc                                19-JAN-94 18:01:52
940220_071248 <  >>     mdiltc                                19-FEB-94 07:11:58
940224_162658 <  >>     mdiltc                                24-FEB-94 16:21:39
940227_145237 <  >>     mdiltc                                27-FEB-94 14:48:06
940227_145237    >>     mdiltc                                27-FEB-94 14:49:34
940304_212004 <  >>     mdiltc                                 4-MAR-94 21:18:58
940304_222132    >>         mdiltc                             4-MAR-94 22:20:29
940305_195329 <  >>             mdiltc                         5-MAR-94 19:50:18
940305_195329    >>             mdiltc                         5-MAR-94 19:51:47
940305_202112    >>             mdiltc                         5-MAR-94 20:19:42
940313_155528    >> mdiltc                                    12-MAR-94 15:54:16
940313_155528    >> mdiltc                                    12-MAR-94 15:56:31
940314_155356 <  >> mdiltc                                    13-MAR-94 15:51:40
940314_155356    >> mdiltc                                    13-MAR-94 15:52:48
940315_212548 <  >> mdiltc                                    14-MAR-94 21:24:55
940401_142048    >>     mdiltc                                 1-APR-94 14:19:44
940406_140020    >>     mdiltc                                 6-APR-94 13:59:27
940408_063328    >>     mdiltcs1                               8-APR-94 06:32:09
940408_072850    >>     mdiltcs1                               8-APR-94 07:27:37
940408_083541 <  >>     mdiltcs2                               8-APR-94 08:34:29
940408_094207 <  >> mdiltcs4                                   8-APR-94 09:40:52
940422_202556    >> mdiltc                                    22-APR-94 20:24:44
940423_023654 <  >>             mdiltc                        23-APR-94 02:36:10
940424_193854    >>                     mdiltc                24-APR-94 20:14:40

940424_202445    >>     mdiltc                                24-APR-94 20:23:41
940424_202445    >>     mdiltc                                24-APR-94 20:29:43
940424_203144 <  >>     mdiltc                                24-APR-94 20:30:53
940426_041407 <  >>     mdiltc                                26-APR-94 04:10:34
940504_101134    >> mdiltc                                     4-MAY-94 10:10:26
940506_122453 <  >> mdiltc                                     6-MAY-94 12:23:11
------------------------------ Searching for mdifocus ------------------------------
930901_063001 <  >>     mdifocus                               1-SEP-93 06:30:09
931001_202439    >> mdifocus                                   1-OCT-93 20:24:17
931008_013816    >>     mdifocus                               8-OCT-93 01:37:49
931121_200140    >> mdifocus                                  21-NOV-93 20:01:20
931121_200140    >> mdifocus                                  21-NOV-93 20:18:48
931123_001958    >> mdifocus                                  23-NOV-93 00:19:40
931218_013527 <  >>     mdifocus                              18-DEC-93 01:33:51
931218_221904 <  >> mdifocus                                  18-DEC-93 22:18:37
931220_071334 <  >> mdifocus                                  20-DEC-93 07:13:21
931220_223624    >> mdifocus                                  20-DEC-93 22:36:05
931222_202513    >>     mdifocus                              22-DEC-93 20:24:31
931223_010434 <  >> mdifocus                                  23-DEC-93 01:04:17
931223_013528    >>     mdifocus                              23-DEC-93 05:33:11
931223_053833    >> mdifocus                                  23-DEC-93 05:37:59
931223_170953 <  >>     mdifocus                              23-DEC-93 17:09:14
931223_174402    >>     mdifocus                              23-DEC-93 17:43:41
940106_232957    >>                     mdifocus               6-JAN-94 23:29:40
940112_233031    >> mdifocus                                  12-JAN-94 23:30:08
940120_131844    >> mdifocus                                  20-JAN-94 13:17:54
940219_025607    >>             mdifocus                      19-FEB-94 02:55:44
940305_081947    >>         mdifocus                           5-MAR-94 08:19:21
940322_034029    >>         mdifocus                          22-MAR-94 03:40:13
940326_010023    >> mdifocus                                  26-MAR-94 01:24:09
940402_045113    >> mdifocus                                   2-APR-94 04:50:59
940406_010943 <  >>     mdifocus                               6-APR-94 01:09:23
940406_012550    >> mdifocus                                   6-APR-94 01:25:29
940407_000702    >>     mdifocus                               7-APR-94 00:06:44
940407_151311    >> mdifocus                                   7-APR-94 15:12:53
940422_195952    >> mdifocus                                  22-APR-94 19:59:36
940422_200839    >> mdifocus                                  22-APR-94 20:08:22
940426_043043    >>             mdifocus                      26-APR-94 04:30:13
940503_162350    >> mdifocus                                   3-MAY-94 16:23:24
940504_080320    >>     mdifocus                               4-MAY-94 08:02:50
940504_092535    >>     mdifocus                               4-MAY-94 09:25:13
940505_115723    >> mdifocus                                   5-MAY-94 11:57:00
940505_124057    >> mdifocus                                   5-MAY-94 12:39:04
940505_141459    >> mdifocus                                   5-MAY-94 14:14:36
940505_171316    >> mdifocus                                   5-MAY-94 17:12:59
940506_102651    >>     mdifocus                               6-MAY-94 10:26:27
940506_110703    >> mdifocus                                   6-MAY-94 11:03:24
------------------------------ Searching for mdiiss ------------------------------
931001_204141 <  >> mdiiss                                     1-OCT-93 20:40:51
931001_205218 <  >> mdiiss                                     1-OCT-93 20:50:00
931008_015736 <  >>     mdiiss                                 8-OCT-93 01:56:40
931129_024915    >> mdiiss                                    29-NOV-93 02:47:53
931129_032516    >>     mdiiss                                29-NOV-93 03:21:30
931129_040346    >>     mdiiss                                29-NOV-93 04:03:08
931129_041502    >>     mdiiss                                29-NOV-93 04:14:23
931129_041502    >> mdiiss                                    29-NOV-93 04:45:07
931129_044855    >> mdiiss                                    29-NOV-93 04:48:13
931129_051101    >> mdiiss                                    29-NOV-93 05:09:44

931129_061106    >> mdiiss                                    29-NOV-93 06:10:25
931222_230333    >> mdiiss                                    23-DEC-93 00:39:52
931223_004152    >> mdiiss                                    23-DEC-93 00:40:53
931223_074514    >> mdiiss                                    23-DEC-93 07:43:51
931223_193352    >>     mdiiss                                23-DEC-93 21:46:41
931223_193352    >>     mdiiss                                23-DEC-93 21:47:22
940113_073052    >>     mdiiss                                13-JAN-94 07:29:55
940113_172735 <  >> mdiiss                                    13-JAN-94 17:24:21
940113_181243 <  >> mdiiss                                    13-JAN-94 18:10:17
940120_180210    >> mdiiss                                    20-JAN-94 18:01:08
940121_093056    >>     mdiiss                                21-JAN-94 09:29:13
940401_183733    >> mdiiss                                     1-APR-94 18:35:05
940425_213204    >> mdiiss                                    25-APR-94 21:29:47
940504_103717    >> mdiiss                                     4-MAY-94 10:36:14
------------------------------ Searching for mdidust ------------------------------
931001_234506    >> mdidust                                    1-OCT-93 23:44:52
931002_214453    >> mdidust                                    2-OCT-93 21:44:33
931008_201704 <  >>     mdidust                                8-OCT-93 20:16:19
------------------------------ Searching for mdilaser ------------------------------
931229_070623    >>                 mdilaser                  29-DEC-93 07:04:24
940113_011157    >> mdilaser                                  13-JAN-94 01:08:38
XXXXXX_XXXXXX    >>             mdilaser                      21-JAN-94 14:56:53
XXXXXX_XXXXXX    >> mdilaser                                  29-MAR-94 00:14:02
940329_003243    >> mdilaser                                  29-MAR-94 00:30:51
XXXXXX_XXXXXX    >>     mdilaser                              29-MAR-94 03:52:57
940329_035721    >>     mdilaser                              29-MAR-94 03:56:31
XXXXXX_XXXXXX    >>     mdilaser                              29-MAR-94 04:54:58
940402_033500    >> mdilaser                                   2-APR-94 03:33:55
940406_185418    >>                 mdilaser                   6-APR-94 18:52:56
940412_170906    >>         mdilaser                          12-APR-94 17:24:17
940412_183825 <  >> mdilaser                                  12-APR-94 18:36:11
940425_192438    >>         mdilaser                          25-APR-94 19:22:44
940504_143015    >> mdilaser                                   4-MAY-94 15:23:59
------------------------------ Searching for shutest ------------------------------
XXXXXX_XXXXXX    >>     shutest                               21-JUL-93 22:35:03
XXXXXX_XXXXXX    >>     shutest                               22-JUL-93 15:20:17
XXXXXX_XXXXXX    >>     shutest                               22-JUL-93 21:20:36
XXXXXX_XXXXXX    >>     shutest                               22-JUL-93 21:26:52
XXXXXX_XXXXXX    >>     shutest                               26-JUL-93 23:46:57
XXXXXX_XXXXXX    >>     shutest                               26-JUL-93 23:49:20
XXXXXX_XXXXXX    >>     shutest                               29-JUL-93 18:58:54
930729_200816    >>     shutest                               29-JUL-93 20:41:54
930729_200816    >>     shutest                               29-JUL-93 20:53:13
930729_200816    >>     shutest                               29-JUL-93 20:56:52
930729_200816    >>     shutest                               29-JUL-93 21:00:20
931008_152008    >>     shutest                                8-OCT-93 17:04:33
931008_152008    >>     shutest                                8-OCT-93 17:09:10
931223_193352    >>     shutest                               23-DEC-93 20:49:52
940505_124057    >> shutest                                    5-MAY-94 13:44:47
------------------------------ Searching for mdicom ------------------------------
XXXXXX_XXXXXX    >>     mdicom                                28-JUL-93 21:47:53
XXXXXX_XXXXXX    >>     mdicom                                28-JUL-93 22:26:47
XXXXXX_XXXXXX    >>     mdicom                                28-JUL-93 22:40:42
XXXXXX_XXXXXX    >>     mdicom                                 4-AUG-93 16:50:43
XXXXXX_XXXXXX    >>     mdicom                                 4-AUG-93 16:58:18
XXXXXX_XXXXXX    >>     mdicom                                 5-AUG-93 17:27:20

XXXXXX_XXXXXX    >>     mdicom                                19-AUG-93 05:23:05
XXXXXX_XXXXXX    >>     mdicom                                21-AUG-93 17:41:59
XXXXXX_XXXXXX    >>     mdicom                                12-SEP-93 20:19:04
XXXXXX_XXXXXX    >>     mdicom                                 1-OCT-93 15:16:29
XXXXXX_XXXXXX    >>     mdicom                                12-OCT-93 18:25:32
XXXXXX_XXXXXX    >>     mdicom                                12-OCT-93 18:45:40
XXXXXX_XXXXXX    >> mdicom                                    20-NOV-93 17:48:56
XXXXXX_XXXXXX    >>     mdicom                                20-NOV-93 20:11:31
XXXXXX_XXXXXX    >>     mdicom                                21-NOV-93 17:49:52
XXXXXX_XXXXXX    >>     mdicom                                22-NOV-93 21:43:40
XXXXXX_XXXXXX    >>     mdicom                                23-NOV-93 05:44:34
931123_103215    >>     mdicom                                23-NOV-93 21:37:18
XXXXXX_XXXXXX    >>     mdicom                                24-NOV-93 04:39:21
XXXXXX_XXXXXX    >>                 mdicom                     5-DEC-93 22:03:31
XXXXXX_XXXXXX    >>     mdicom                                14-DEC-93 00:40:03
XXXXXX_XXXXXX    >>     mdicom                                17-DEC-93 22:13:32
XXXXXX_XXXXXX    >>     mdicom                                18-DEC-93 20:23:14
931218_203257    >>         mdicom                            18-DEC-93 20:50:29
931218_203257    >>         mdicom                            18-DEC-93 20:51:07
931218_203257    >>         mdicom                            18-DEC-93 20:53:15
931218_203257    >>         mdicom                            18-DEC-93 21:01:23
XXXXXX_XXXXXX    >>     mdicom                                19-DEC-93 17:35:54
931220_071334    >> mdicom                                    20-DEC-93 07:31:46
931220_090106    >> mdicom                                    20-DEC-93 20:57:01
XXXXXX_XXXXXX    >>     mdicom                                21-DEC-93 07:46:30
XXXXXX_XXXXXX    >>     mdicom                                21-DEC-93 18:16:35
931222_030252    >>         mdicom                            22-DEC-93 03:11:18
931222_052755    >>         mdicom                            22-DEC-93 10:06:26
XXXXXX_XXXXXX <  >>     mdicom                                22-DEC-93 10:27:04
931222_163837    >>                                 mdicom    22-DEC-93 17:22:15
931222_173317    >>                                 mdicom    22-DEC-93 17:46:58
931223_111900    >>         mdicom                            23-DEC-93 13:36:49
XXXXXX_XXXXXX    >>     mdicom                                11-JAN-94 03:40:57
XXXXXX_XXXXXX    >>     mdicom                                12-JAN-94 14:40:18
XXXXXX_XXXXXX    >>     mdicom                                13-JAN-94 22:55:04
XXXXXX_XXXXXX    >>     mdicom                                19-JAN-94 15:29:07
XXXXXX_XXXXXX    >>     mdicom                                15-FEB-94 18:42:38
XXXXXX_XXXXXX    >>     mdicom                                 2-APR-94 18:55:13
940406_012550    >> mdicom                                     6-APR-94 02:22:47
XXXXXX_XXXXXX    >>     mdicom                                17-APR-94 01:59:57
XXXXXX_XXXXXX    >>     mdicom                                19-APR-94 21:13:44
XXXXXX_XXXXXX    >>     mdicom                                20-APR-94 18:44:45
940422_003739    >> mdicom                                    22-APR-94 02:27:59
XXXXXX_XXXXXX    >>     mdicom                                25-APR-94 16:30:31
XXXXXX_XXXXXX    >>     mdicom                                 3-MAY-94 10:53:50
940505_141459    >> mdicom                                     5-MAY-94 15:19:04
------------------------------ Searching for lcmcom ------------------------------
XXXXXX_XXXXXX               lcmcom                            19-AUG-93 05:26:26  ## Camera/Shutter commissioning
XXXXXX_XXXXXX           lcmcom                                19-AUG-93 05:26:26  ## Camera/Shutter commissioning
XXXXXX_XXXXXX               lcmcom                            21-AUG-93 17:53:49  ## Camera/Shutter commissioning
XXXXXX_XXXXXX           lcmcom                                21-AUG-93 17:53:49  ## Camera/Shutter commissioning
930908_220032    >>     lcmcom2                                9-SEP-93 00:29:35
930909_221948    >>     lcmcom2                                9-SEP-93 23:59:38
930910_004348    >>     lcmcom2                               10-SEP-93 00:48:18
930910_004348    >>     lcmcom2                               10-SEP-93 00:49:51
930912_203245 <             lcmcom                            12-SEP-93 20:34:31  ## Camera/Shutter commissioning
930912_203245           lcmcom                                12-SEP-93 20:34:31  ## Camera/Shutter commissioning
930912_203958 <         lcmcom                                12-SEP-93 20:40:21  ## Camera/Shutter commissioning
930912_203958               lcmcom                            12-SEP-93 20:40:21  ## Camera/Shutter commissioning

930912_205724 <         lcmcom                                12-SEP-93 20:57:59  ## Camera/Shutter commissioning
930912_205724               lcmcom                            12-SEP-93 20:57:59  ## Camera/Shutter commissioning
931001_152636               lcmcom                             1-OCT-93 15:26:37  ## Camera commissioning
XXXXXX_XXXXXX    >>     lcmcom                                 7-OCT-93 19:14:27
931007_223139               lcmcom                             7-OCT-93 22:31:40  ## Camera commissioning
931012_185700               lcmcom                            12-OCT-93 18:57:01  ## Camera commissioning
931120_175953           lcmcom                                20-NOV-93 17:59:54  ## Camera commissioning
931120_202146               lcmcom                            20-NOV-93 20:21:47  ## Camera commissioning
931121_180055               lcmcom                            21-NOV-93 18:00:56  ## Camera commissioning
931122_215549               lcmcom                            22-NOV-93 21:55:50  ## Camera commissioning
931123_055608               lcmcom                            23-NOV-93 05:56:09  ## Camera commissioning
931123_214817               lcmcom                            23-NOV-93 21:48:18  ## Camera commissioning
931124_044913               lcmcom                            24-NOV-93 04:49:14  ## Camera commissioning
931205_221408                           lcmcom                 5-DEC-93 22:14:09  ## Camera commissioning
931214_004957               lcmcom                            14-DEC-93 00:49:58  ## Camera commissioning
931217_222334               lcmcom                            17-DEC-93 22:23:35  ## Camera commissioning
931218_203257               lcmcom                            18-DEC-93 20:32:58  ## Camera commissioning
931219_174652               lcmcom                            19-DEC-93 17:46:53  ## Camera commissioning
931220_074608           lcmcom                                20-DEC-93 07:46:10  ## Camera commissioning
931220_210852           lcmcom                                20-DEC-93 21:08:53  ## Camera commissioning
931221_075531               lcmcom                            21-DEC-93 07:55:32  ## Camera commissioning
931221_182624               lcmcom                            21-DEC-93 18:26:25  ## Camera commissioning
931222_032258                   lcmcom                        22-DEC-93 03:22:59  ## Camera commissioning
931222_101439                   lcmcom                        22-DEC-93 10:14:40  ## Camera commissioning
931222_103549               lcmcom                            22-DEC-93 10:35:50  ## Camera commissioning
931222_173317                                         lcmcom  22-DEC-93 17:33:19  ## Camera commissioning
931223_134655                   lcmcom                        23-DEC-93 13:46:56  ## Camera commissioning
940111_035400               lcmcom                            11-JAN-94 03:54:01  ## Camera commissioning
940112_145145               lcmcom                            12-JAN-94 14:51:46  ## Camera commissioning
940119_154231               lcmcom                            19-JAN-94 15:42:32  ## Camera commissioning
940215_185257               lcmcom                            15-FEB-94 18:52:58  ## Camera commissioning
940402_191303               lcmcom                             2-APR-94 19:13:04  ## Camera commissioning
940406_023226           lcmcom                                 6-APR-94 02:32:28  ## Camera commissioning
940417_021220               lcmcom                            17-APR-94 02:12:22  ## Camera commissioning
940419_212453               lcmcom                            19-APR-94 21:24:54  ## Camera commissioning
940420_185642               lcmcom                            20-APR-94 18:56:43  ## Camera commissioning
940422_024117           lcmcom                                22-APR-94 02:41:18  ## Camera commissioning
940425_164209               lcmcom                            25-APR-94 16:42:11  ## Camera commissioning
940503_110612               lcmcom                             3-MAY-94 11:06:13  ## Camera commissioning
940505_152950           lcmcom                                 5-MAY-94 15:29:51  ## Camera commissioning
------------------------------ Searching for ipcom ------------------------------
XXXXXX_XXXXXX           lipcom                                19-AUG-93 05:24:40  ## IP commissioning
XXXXXX_XXXXXX               lipcom                            19-AUG-93 05:24:41  ## IP commissioning
XXXXXX_XXXXXX           lipcom                                21-AUG-93 17:48:03  ## IP commissioning
XXXXXX_XXXXXX               lipcom                            21-AUG-93 17:48:04  ## IP commissioning
XXXXXX_XXXXXX           lipcom                                21-AUG-93 17:52:22  ## IP commissioning
XXXXXX_XXXXXX               lipcom                            21-AUG-93 17:52:23  ## IP commissioning
930912_205724               lipcom                            12-SEP-93 22:32:54  ## IP commissioning
931008_214523 <  >>         lipcom                             8-OCT-93 21:44:53
931110_003654           lipcom                                10-NOV-93 00:36:24  ## IP commissioning
931124_084903    >> ipcom                                     24-NOV-93 21:30:32
931124_084903    >> ipcom                                     24-NOV-93 21:33:07
931124_084903    >> ipcom                                     24-NOV-93 22:05:20
931205_231740           ipcom                                  5-DEC-93 23:47:58  ## IP commissioning
931217_230607           ipcom                                 17-DEC-93 23:13:29  ## IP commissioning
931218_211911           ipcom                                 18-DEC-93 21:27:17  ## IP commissioning
931219_190432           ipcom                                 19-DEC-93 19:12:26  ## IP commissioning
931219_192526    >> tvipcom                                   20-DEC-93 00:57:00
931219_192526    >>     tvipcom                               20-DEC-93 01:06:48

931219_192526    >>     tvipcom                               20-DEC-93 01:08:47
931219_192526    >>     tvipcom                               20-DEC-93 01:10:49
931220_090106 <         ipcom                                 20-DEC-93 08:57:21  ## IP commissioning
931220_215334           ipcom                                 20-DEC-93 22:02:17  ## IP commissioning
931220_223624    >> ipcom                                     20-DEC-93 22:56:00
931221_082209           ipcom                                 21-DEC-93 08:30:21  ## IP commissioning
931221_191940           ipcom                                 21-DEC-93 19:28:11  ## IP commissioning
931222_105703               ipcom                             22-DEC-93 11:04:43  ## IP commissioning
931222_182058           ipcom                                 22-DEC-93 18:28:56  ## IP commissioning
931223_145437           ipcom                                 23-DEC-93 15:07:47  ## IP commissioning
XXXXXX_XXXXXX    >>         ipcom                             11-JAN-94 15:02:41
940112_154832           ipcom                                 12-JAN-94 15:58:26  ## IP commissioning
940120_141706    >> ipcom                                     20-JAN-94 15:21:21
940215_210515           ipcom                                 15-FEB-94 21:16:15  ## IP commissioning
940411_204121    >> ipcom                                     11-APR-94 22:25:12
940417_030620           ipcom                                 17-APR-94 03:13:57  ## IP commissioning
XXXXXX_XXXXXX    >>     ipcom                                 20-APR-94 01:21:36
940422_033011           ipcom                                 22-APR-94 03:39:44  ## IP commissioning
940503_123543    >>     ipcom                                  3-MAY-94 13:01:55
------------------------------ Searching for ipmemtst ------------------------------
931129_210217    >> ipmemtst                                  29-NOV-93 21:02:30
931129_210217    >> ipmemtst                                  29-NOV-93 21:05:25
931129_210217    >> ipmemtst                                  29-NOV-93 21:08:53
XXXXXX_XXXXXX    >>     ipmemtst                               4-DEC-93 16:45:42
931204_165529    >> ipmemtst                                   4-DEC-93 17:42:12
940111_151731    >> ipmemtst                                  11-JAN-94 16:29:24
XXXXXX_XXXXXX    >>         ipmemtst                           1-MAR-94 18:49:55
XXXXXX_XXXXXX    >>     ipmemtst                               1-MAR-94 19:54:55
XXXXXX_XXXXXX    >>     ipmemtst                               1-MAR-94 21:34:45
XXXXXX_XXXXXX    >>     ipmemtst                               1-MAR-94 22:00:55
940302_143456    >>     ipmemtst                               2-MAR-94 16:17:54
940303_003621    >> ipmemtst                                   3-MAR-94 02:05:20
940303_003621    >> ipmemtst                                   3-MAR-94 02:06:41
940303_003621    >> ipmemtst                                   3-MAR-94 04:09:24
XXXXXX_XXXXXX    >>     ipmemtst                               4-MAR-94 20:35:32
940309_022837    >>     ipmemtst                               9-MAR-94 14:29:54
940309_022837    >>     ipmemtst                               9-MAR-94 14:31:15
XXXXXX_XXXXXX               ipmemtst                          20-APR-94 01:28:03
940422_033011               ipmemtst                          22-APR-94 03:44:46
940503_123543               ipmemtst                           3-MAY-94 13:04:33
940506_110703    >> ipmemtst                                   6-MAY-94 11:38:13
------------------------------ Searching for iptbltst ------------------------------
931201_181306    >>     iptbltst                               1-DEC-93 18:13:13
931201_211929    >>                 iptbltst                   1-DEC-93 21:43:09
931204_165529    >>     iptbltst                               4-DEC-93 16:55:39
931222_030252    >>         iptbltst                          22-DEC-93 03:03:08
931222_221349    >> iptbltst                                  22-DEC-93 22:14:45
931229_134517    >> iptbltst                                  29-DEC-93 22:49:03
931229_134517    >> iptbltst                                  29-DEC-93 22:51:30
940111_151731    >>             iptbltst                      11-JAN-94 16:18:13
940111_151731    >>             iptbltst                      11-JAN-94 16:23:13
940111_151731    >>     iptbltst                              11-JAN-94 16:35:21
XXXXXX_XXXXXX    >>     iptbltst                              11-JAN-94 18:21:19
940111_185730    >>     iptbltst                              11-JAN-94 18:57:39
940228_220408    >> iptbltst                                  28-FEB-94 22:04:14
940228_220408    >> iptbltst                                  28-FEB-94 22:17:40
XXXXXX_XXXXXX    >>     iptbltst                               1-MAR-94 18:26:46
XXXXXX_XXXXXX    >>     iptbltst                               1-MAR-94 19:59:50

XXXXXX_XXXXXX    >>     iptbltst                               1-MAR-94 21:23:26
XXXXXX_XXXXXX    >>     iptbltst                               1-MAR-94 21:24:55
XXXXXX_XXXXXX    >>     iptbltst                               1-MAR-94 21:30:24
XXXXXX_XXXXXX    >>     iptbltst                               1-MAR-94 22:19:03
XXXXXX_XXXXXX    >>     iptbltst                               1-MAR-94 22:23:41
940302_143456    >>     iptbltst                               2-MAR-94 16:15:08
940302_143456    >>     iptbltst                               2-MAR-94 16:23:39
940303_003621    >> iptbltst                                   3-MAR-94 02:10:19
940303_003621    >> iptbltst                                   3-MAR-94 02:13:50
940303_003621    >> iptbltst                                   3-MAR-94 04:08:18
940303_003621    >> iptbltst                                   3-MAR-94 04:12:39
940303_041355    >> iptbltst                                   3-MAR-94 04:14:00
940303_041454 <  >> iptbltst                                   3-MAR-94 04:15:20
940303_145012    >>             iptbltst                       3-MAR-94 15:39:56
940303_145012    >>             iptbltst                       3-MAR-94 15:42:33
XXXXXX_XXXXXX    >>     iptbltst                               4-MAR-94 20:39:22
XXXXXX_XXXXXX    >>     iptbltst                               4-MAR-94 20:40:46
940304_233629    >>             iptbltst                       4-MAR-94 23:47:12
940305_174506    >>             iptbltst                       5-MAR-94 18:08:45
940311_200352    >>         iptbltst                          11-MAR-94 21:20:35
940312_161816    >>     iptbltst                              12-MAR-94 16:23:47
940316_164656    >>     iptbltst                              16-MAR-94 17:54:18
XXXXXX_XXXXXX    >>             iptbltst                      19-APR-94 15:11:28
XXXXXX_XXXXXX    >>             iptbltst                      19-APR-94 16:10:17
940420_013337               iptbltst                          20-APR-94 01:44:18
940422_035054               iptbltst                          22-APR-94 03:57:47
940423_223815    >> iptbltst                                  24-APR-94 00:02:22
940423_223815    >> iptbltst                                  24-APR-94 00:06:19
940503_131002               iptbltst                           3-MAY-94 13:20:35
------------------------------ Searching for ipdmatst ------------------------------
931204_165529    >>         ipdmatst                           4-DEC-93 16:57:50
XXXXXX_XXXXXX    >>     ipdmatst                              10-JAN-94 16:51:11
940420_013337               ipdmatst                          20-APR-94 01:44:58
940422_035054               ipdmatst                          22-APR-94 03:58:27
940503_131002                   ipdmatst                       3-MAY-94 13:21:16
------------------------------ Searching for ipdynmac ------------------------------
931105_210122    >> ipdynmac                                   5-NOV-93 21:30:25
931105_210122    >> ipdynmac                                   5-NOV-93 21:34:55
931124_084903    >> ipdynmac                                  24-NOV-93 21:57:56
931124_223018    >>     ipdynmac                              24-NOV-93 23:22:15
931219_192526    >>     ipdynmac                              19-DEC-93 20:31:32
931222_221349    >> ipdynmac                                  22-DEC-93 22:19:23
931222_221349    >> ipdynmac                                  22-DEC-93 22:21:13
931229_134517    >> ipdynmac                                  29-DEC-93 22:53:04
940111_151731    >>     ipdynmac                              11-JAN-94 16:36:15
940111_185730    >>     ipdynmac                              11-JAN-94 18:58:34
XXXXXX_XXXXXX    >>     ipdynmac                              11-JAN-94 20:22:52
940314_155356    >> ipdynmac                                  13-MAR-94 18:27:32
940314_155356    >> ipdynmac                                  13-MAR-94 18:34:53
940314_184353    >> ipdynmac                                  13-MAR-94 18:52:35
940314_194041    >> ipdynmac                                  13-MAR-94 19:47:38
940316_164656    >>     ipdynmac                              16-MAR-94 17:58:22
940411_223537    >> ipdynmac                                  11-APR-94 23:00:33
940420_013337               ipdynmac                          20-APR-94 02:01:26
940422_035054           ipdynmac                              22-APR-94 04:12:36
940503_131002           ipdynmac                               3-MAY-94 13:37:49
------------------------------ Searching for ipseqtst ------------------------------

931124_223018    >>     ipseqtst                              24-NOV-93 23:24:15
XXXXXX_XXXXXX    >>                 ipseqtst                  29-NOV-93 19:07:40
XXXXXX_XXXXXX    >>                 ipseqtst                  29-NOV-93 19:17:42
XXXXXX_XXXXXX    >> ipseqtst                                  29-NOV-93 19:30:51
931129_212531    >> ipseqtst                                  29-NOV-93 21:25:41
931129_213607    >> ipseqtst                                  29-NOV-93 21:36:17
931129_213607    >>     ipseqtst                              29-NOV-93 22:36:28
931129_213607    >>     ipseqtst                              29-NOV-93 22:41:58
931129_224230    >> ipseqtst                                  29-NOV-93 22:53:13
931129_225448    >> ipseqtst                                  29-NOV-93 22:58:05
931130_222455    >>     ipseqtst                              30-NOV-93 22:25:00
XXXXXX_XXXXXX    >>             ipseqtst                       1-DEC-93 21:17:31
XXXXXX_XXXXXX    >>         ipseqtst                           2-DEC-93 22:21:11
XXXXXX_XXXXXX    >>         ipseqtst                           2-DEC-93 22:22:44
XXXXXX_XXXXXX    >>         ipseqtst                           2-DEC-93 22:30:29
XXXXXX_XXXXXX    >>         ipseqtst                           3-DEC-93 00:37:28
931203_004511    >>         ipseqtst                           3-DEC-93 00:45:44
XXXXXX_XXXXXX    >>         ipseqtst                           3-DEC-93 19:20:13
931204_165529    >>         ipseqtst                           4-DEC-93 17:00:40
931204_165529    >>         ipseqtst                           4-DEC-93 17:16:13
XXXXXX_XXXXXX    >>     ipseqtst                               8-DEC-93 23:26:20
XXXXXX_XXXXXX    >>     ipseqtst                               8-DEC-93 23:27:43
XXXXXX_XXXXXX    >>     ipseqtst                               8-DEC-93 23:38:38
XXXXXX_XXXXXX    >>     ipseqtst                               9-DEC-93 00:48:56
XXXXXX_XXXXXX    >>     ipseqtst                               9-DEC-93 01:14:07
XXXXXX_XXXXXX    >> ipseqtst                                  10-DEC-93 22:18:06
XXXXXX_XXXXXX    >>     ipseqtst                              10-DEC-93 22:39:07
931214_020937    >> ipseqtst                                  14-DEC-93 02:28:26
XXXXXX_XXXXXX    >>     ipseqtst                              14-DEC-93 14:59:51
931215_163228    >> ipseqtst                                  15-DEC-93 16:32:37
931219_192526    >>     ipseqtst                              19-DEC-93 20:32:47
931219_192526    >>     ipseqtst                              19-DEC-93 21:35:13
931222_221349    >> ipseqtst                                  22-DEC-93 22:24:19
931222_224109    >> ipseqtst                                  22-DEC-93 22:41:21
931229_134517    >> ipseqtst                                  29-DEC-93 22:54:59
931229_225715    >> ipseqtst                                  29-DEC-93 23:21:57
940111_185730    >>     ipseqtst                              11-JAN-94 19:00:05
940314_184201    >> ipseqtst                                  13-MAR-94 18:43:01
940314_185641    >> ipseqtst                                  13-MAR-94 18:56:57
940411_223537    >> ipseqtst                                  11-APR-94 23:03:42
940420_013337               ipseqtst                          20-APR-94 02:02:31
940422_035054           ipseqtst                              22-APR-94 04:14:09
940503_131002           ipseqtst                               3-MAY-94 13:43:31
------------------------------ Searching for lsqcom ------------------------------
931001_164601           lsqcom                                 1-OCT-93 16:46:02  ## DEP sequence interpreter commissioning
931007_234103           lsqcom                                 7-OCT-93 23:41:04  ## DEP sequence interpreter commissioning
931012_210906           lsqcom                                12-OCT-93 21:09:08  ## DEP sequence interpreter commissioning
931120_184514           lsqcom                                20-NOV-93 18:45:16  ## DEP sequence interpreter commissioning
931120_210453           lsqcom                                20-NOV-93 21:04:55  ## DEP sequence interpreter commissioning
931121_192739           lsqcom                                21-NOV-93 19:27:40  ## DEP sequence interpreter commissioning
931122_230301           lsqcom                                22-NOV-93 23:03:03  ## DEP sequence interpreter commissioning
931123_070030           lsqcom                                23-NOV-93 07:00:31  ## DEP sequence interpreter commissioning
931123_225411           lsqcom                                23-NOV-93 22:54:12  ## DEP sequence interpreter commissioning
931124_055501           lsqcom                                24-NOV-93 05:55:02  ## DEP sequence interpreter commissioning
931124_055501    >> lsqcom                                    24-NOV-93 06:15:38
931124_055501    >> lsqcom                                    24-NOV-93 06:31:35
931205_231740           lsqcom                                 5-DEC-93 23:17:41  ## DEP sequence interpreter commissioning
931214_020937           lsqcom                                14-DEC-93 02:09:38  ## DEP sequence interpreter commissioning
931217_230607           lsqcom                                17-DEC-93 23:06:08  ## DEP sequence interpreter commissioning

931218_211911           lsqcom                                18-DEC-93 21:19:13  ## DEP sequence interpreter commissioning
931219_190432           lsqcom                                19-DEC-93 19:04:33  ## DEP sequence interpreter commissioning
931220_084656           lsqcom                                20-DEC-93 08:46:57  ## DEP sequence interpreter commissioning
931220_215334           lsqcom                                20-DEC-93 21:53:35  ## DEP sequence interpreter commissioning
931221_082209           lsqcom                                21-DEC-93 08:22:10  ## DEP sequence interpreter commissioning
931221_191940           lsqcom                                21-DEC-93 19:19:41  ## DEP sequence interpreter commissioning
931222_042245           lsqcom                                22-DEC-93 04:22:46  ## DEP sequence interpreter commissioning
931222_105703               lsqcom                            22-DEC-93 10:57:04  ## DEP sequence interpreter commissioning
931222_182058           lsqcom                                22-DEC-93 18:20:59  ## DEP sequence interpreter commissioning
931223_145437           lsqcom                                23-DEC-93 14:54:39  ## DEP sequence interpreter commissioning
940111_043557           lsqcom                                11-JAN-94 04:35:58  ## DEP sequence interpreter commissioning
940112_154832           lsqcom                                12-JAN-94 15:48:33  ## DEP sequence interpreter commissioning
940119_163231           lsqcom                                19-JAN-94 16:32:33  ## DEP sequence interpreter commissioning
940215_210515           lsqcom                                15-FEB-94 21:05:16  ## DEP sequence interpreter commissioning
940402_202654           lsqcom                                 2-APR-94 20:26:55  ## DEP sequence interpreter commissioning
940406_031203           lsqcom                                 6-APR-94 03:12:04  ## DEP sequence interpreter commissioning
940417_030620           lsqcom                                17-APR-94 03:06:21  ## DEP sequence interpreter commissioning
940419_221546           lsqcom                                19-APR-94 22:15:47  ## DEP sequence interpreter commissioning
940420_193742           lsqcom                                20-APR-94 19:37:43  ## DEP sequence interpreter commissioning
940422_033011           lsqcom                                22-APR-94 03:30:13  ## DEP sequence interpreter commissioning
940423_002948    >>         lsqcom                            23-APR-94 01:56:56
940425_173206           lsqcom                                25-APR-94 17:32:07  ## DEP sequence interpreter commissioning
940425_173206    >> lsqcom                                    25-APR-94 18:27:57
XXXXXX_XXXXXX    >>         lsqcom                            26-APR-94 02:48:30
XXXXXX_XXXXXX    >>         lsqcom3                           26-APR-94 02:57:40
XXXXXX_XXXXXX    >>         lsqcom                            26-APR-94 03:02:37
XXXXXX_XXXXXX    >>         lsqcom3                           26-APR-94 03:22:18
XXXXXX_XXXXXX    >> lsqcom                                    26-APR-94 03:27:21
940426_061816    >> lsqcom                                    26-APR-94 07:04:32
940503_123543 <     lsqcom                                     3-MAY-94 12:37:06
940503_123543    >>     lsqcom                                 3-MAY-94 12:37:07
940503_123543       lsqcom                                     3-MAY-94 12:41:41
940503_123543    >>     lsqcom                                 3-MAY-94 12:41:42
XXXXXX_XXXXXX               lsqcom                             4-MAY-94 12:30:02
940505_155753               lsqcom                             5-MAY-94 15:57:55  ## DEP sequence interpreter commissioning
------------------------------ Searching for mdidtune ------------------------------
931002_211056 <  >>                         mdidtune           2-OCT-93 21:10:32
931008_211021 <  >>     mdidtune                               8-OCT-93 21:09:44
931019_215619    >>     mdidtune                              19-OCT-93 21:56:02
931020_214216 <  >>             mdidtune                      20-OCT-93 21:41:54
931021_205245    >>                         mdidtune          21-OCT-93 23:07:52
931021_230906    >>                             mdidtune      21-OCT-93 23:08:49
931022_002305    >>                                 mdidtune  22-OCT-93 00:22:47
931022_015725    >> mdidtune                                  22-OCT-93 01:57:04
931023_000744    >>     mdidtune                              23-OCT-93 00:07:23
931023_013033    >>     mdidtune                              23-OCT-93 01:30:09
931027_233428    >>     mdidtune                              27-OCT-93 23:34:10
931028_002639    >> mdidtune                                  28-OCT-93 00:26:13
931028_155016    >>         mdidtune                          28-OCT-93 15:48:53
931028_183039    >>         mdidtune                          28-OCT-93 18:29:26
931028_190116    >>     mdidtune                              28-OCT-93 18:59:14
931028_195052    >>     mdidtune                              28-OCT-93 19:50:33
931028_212849    >> mdidtune                                  28-OCT-93 21:28:31
931028_223530    >>             mdidtune                      28-OCT-93 22:35:13
931028_234149    >>     mdidtune                              28-OCT-93 23:41:31
931029_014038    >>         mdidtune                          29-OCT-93 01:40:09
931029_021639    >>         mdidtune                          29-OCT-93 02:12:58
931029_024958    >> mdidtune                                  29-OCT-93 02:49:22
931029_034842    >> mdidtune                                  29-OCT-93 03:48:12

931029_041446    >> mdidtune                                  29-OCT-93 04:14:20
XXXXXX_XXXXXX    >>     mdidtune                              29-OCT-93 17:31:19
931029_173353    >> mdidtune                                  29-OCT-93 17:32:05
931029_191854    >>     mdidtune                              29-OCT-93 19:17:21
931029_224151    >> mdidtune                                  29-OCT-93 22:40:58
931029_225806    >> mdidtune                                  29-OCT-93 22:56:42
931104_234942    >>     mdidtune                               4-NOV-93 23:48:56
931105_002930    >>     mdidtune                               5-NOV-93 00:28:57
931105_182100    >>     mdidtune                               5-NOV-93 19:38:00
931105_194016    >>         mdidtune                           5-NOV-93 19:39:41
931105_201216    >> mdidtune                                   5-NOV-93 20:11:40
931126_200554 <  >>         mdidtune                          26-NOV-93 20:05:10
931201_202222 <  >>                 mdidtune                   1-DEC-93 20:22:09
931201_211929    >>                     mdidtune               1-DEC-93 22:10:21
931201_221344    >> mdidtune                                   1-DEC-93 22:15:26
931202_190458    >> mdidtune                                   2-DEC-93 19:04:24
------------------------------ Searching for bigtune ------------------------------
XXXXXX_XXXXXX    >>     vbigtune                              14-JUL-93 02:15:34
XXXXXX_XXXXXX    >>     vbigtune                              14-JUL-93 14:41:24
XXXXXX_XXXXXX    >>     vbigtune                              14-JUL-93 14:42:34
XXXXXX_XXXXXX    >>     vbigtune                              14-JUL-93 15:02:21
930714_161132    >>     vbigtune                              15-JUL-93 00:10:18
930714_161132    >>     vbigtune                              15-JUL-93 00:30:11
930714_161132    >>     vbigtune                              15-JUL-93 00:54:24
XXXXXX_XXXXXX    >>     vbigtune                              16-JUL-93 18:59:03
XXXXXX_XXXXXX    >>     vbigtune                              16-JUL-93 19:16:07
XXXXXX_XXXXXX    >>     vbigtune                              31-JUL-93 14:08:15
930731_144820    >>     vbigtune                              31-JUL-93 16:32:53
930731_164519    >>     vbigtune                              31-JUL-93 18:18:22
930731_183233    >>     vbigtune                              31-JUL-93 20:00:18
XXXXXX_XXXXXX    >>     vbigtune                               1-AUG-93 01:00:29
XXXXXX_XXXXXX    >>     vbigtune                               1-AUG-93 01:01:50
XXXXXX_XXXXXX    >>     vbigtune                               1-AUG-93 01:03:28
XXXXXX_XXXXXX    >>     vbigtune                               1-AUG-93 01:21:50
XXXXXX_XXXXXX    >>     vbigtune                               4-AUG-93 01:51:04
XXXXXX_XXXXXX    >>     vbigtune                               4-AUG-93 05:31:17
XXXXXX_XXXXXX    >>     vbigtune                               4-AUG-93 05:47:42
XXXXXX_XXXXXX    >>     vbigtune                               4-AUG-93 06:12:07
XXXXXX_XXXXXX    >>     vbigtune                               4-AUG-93 06:23:24
XXXXXX_XXXXXX    >>     vbigtune                               4-AUG-93 06:34:33
XXXXXX_XXXXXX    >>     vbigtune                               4-AUG-93 13:47:30
XXXXXX_XXXXXX    >>     bigtune                                9-AUG-93 22:59:49
XXXXXX_XXXXXX    >>     bigtune                                9-AUG-93 23:02:20
XXXXXX_XXXXXX    >>     bigtune                                9-AUG-93 23:45:09
XXXXXX_XXXXXX    >>     bigtune                               10-AUG-93 00:01:21
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 01:06:27
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 01:08:37
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 01:23:09
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 20:08:43
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 20:11:32
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 20:17:13
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 20:35:17
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 21:40:04
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 21:56:46
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 22:09:40
XXXXXX_XXXXXX    >>     bigtune                               21-AUG-93 22:28:06
931002_211056                                   bigtune        2-OCT-93 21:10:58
931008_211021               bigtune                            8-OCT-93 21:10:23
931019_212049               bigtune                           19-OCT-93 21:20:51

931019_212138                       bigtune                   19-OCT-93 21:21:40
931019_215619               bigtune                           19-OCT-93 21:56:21
931020_214216                       bigtune                   20-OCT-93 21:42:19
931021_230906                                       bigtune   21-OCT-93 23:09:27
931022_002305                                        bigtune  22-OCT-93 00:23:26
931022_015725           bigtune                               22-OCT-93 01:57:45
931023_000744               bigtune                           23-OCT-93 00:08:03
931023_013033               bigtune                           23-OCT-93 01:30:53
XXXXXX_XXXXXX               bigtune                           27-OCT-93 17:04:29
931027_203724               bigtune                           27-OCT-93 21:40:43
931027_233428               bigtune                           27-OCT-93 23:34:49
931028_002639           bigtune                               28-OCT-93 00:27:00
931028_155016                   bigtune                       28-OCT-93 15:50:37
931028_183039                   bigtune                       28-OCT-93 18:31:01
931028_190116               bigtune                           28-OCT-93 19:01:37
931028_195052               bigtune                           28-OCT-93 19:51:14
931028_212849           bigtune                               28-OCT-93 21:29:10
931028_223530                       bigtune                   28-OCT-93 22:35:51
931028_234149               bigtune                           28-OCT-93 23:42:10
931029_014038                   bigtune                       29-OCT-93 01:40:59
931029_021639                   bigtune                       29-OCT-93 02:17:01
931029_024958           bigtune                               29-OCT-93 02:50:19
931029_034842           bigtune                               29-OCT-93 03:49:05
931029_041446           bigtune                               29-OCT-93 04:15:07
XXXXXX_XXXXXX           bigtune                               29-OCT-93 15:57:09
931029_173353           bigtune                               29-OCT-93 17:34:14
931029_173353           bigtune                               29-OCT-93 18:03:12
931029_191854               bigtune                           29-OCT-93 19:19:15
931029_203324           bigtune                               29-OCT-93 20:55:17
931029_224151           bigtune                               29-OCT-93 22:42:13
931029_225806           bigtune                               29-OCT-93 22:58:26
931029_225806           bigtune                               29-OCT-93 23:19:33
931104_234942               bigtune                            4-NOV-93 23:50:03
931105_002930               bigtune                            5-NOV-93 00:29:51
931105_194016                   bigtune                        5-NOV-93 19:40:37
931105_201216           bigtune                                5-NOV-93 20:12:37
931106_073424                   bigtune                        6-NOV-93 07:35:38
931106_073424                   bigtune                        6-NOV-93 07:52:47
931106_073424                   bigtune                        6-NOV-93 08:09:50
931106_073424           bigtune                                6-NOV-93 08:27:18
931106_073424           bigtune                                6-NOV-93 08:44:02
931106_073424           bigtune                                6-NOV-93 09:01:10
931106_073424           bigtune                                6-NOV-93 09:18:17
931106_073424           bigtune                                6-NOV-93 09:36:19
931106_073424           bigtune                                6-NOV-93 09:53:02
931106_101607           bigtune                                6-NOV-93 10:17:30
931106_101607           bigtune                                6-NOV-93 10:33:45
931106_101607           bigtune                                6-NOV-93 10:49:55
931106_101607           bigtune                                6-NOV-93 11:05:50
931106_101607           bigtune                                6-NOV-93 11:21:45
931106_101607           bigtune                                6-NOV-93 11:42:04
931106_101607           bigtune                                6-NOV-93 11:58:15
931106_101607           bigtune                                6-NOV-93 12:15:58
931106_101607           bigtune                                6-NOV-93 12:31:53
931126_200554                   bigtune                       26-NOV-93 20:06:17
931201_202248                           bigtune                1-DEC-93 20:23:11
931201_221559           bigtune                                1-DEC-93 22:16:22
931202_190458           bigtune                                2-DEC-93 19:05:22
940409_221733 <  >> bigtune                                    9-APR-94 22:17:38

------------------------------ Searching for supersun ------------------------------
931025_220928    >>         supersun                          25-OCT-93 22:09:13
931025_231427    >>     supersun                              25-OCT-93 23:14:03
------------------------------ Searching for vfdlscan ------------------------------
XXXXXX_XXXXXX    >>     vfdlscan                              31-JUL-93 05:43:53
XXXXXX_XXXXXX    >>     vfdlscan                              31-JUL-93 05:45:45
XXXXXX_XXXXXX    >>     vfdlscan                              31-JUL-93 06:07:28
XXXXXX_XXXXXX    >>     vfdlscan                               4-AUG-93 01:40:19
XXXXXX_XXXXXX    >>     vfdlscan                               4-AUG-93 01:41:52
XXXXXX_XXXXXX    >>     vfdlscan                               4-AUG-93 01:43:41
XXXXXX_XXXXXX    >>     vfdlscan                               4-AUG-93 20:40:28
XXXXXX_XXXXXX    >>     vfdlscan                               4-AUG-93 20:43:18
XXXXXX_XXXXXX    >>     vfdlscan                               4-AUG-93 21:35:32
XXXXXX_XXXXXX    >>     vfdlscan                               4-AUG-93 21:36:59
931021_205245    >>                 vfdlscan                  21-OCT-93 22:23:08
931021_205245    >>                 vfdlscan                  21-OCT-93 22:25:11
931021_205245    >>                 vfdlscan                  21-OCT-93 22:28:00
931021_205245    >>                 vfdlscan                  21-OCT-93 22:31:33
931021_205245    >>                 vfdlscan                  21-OCT-93 22:33:44
931021_205245    >>                 vfdlscan                  21-OCT-93 22:36:54
931021_205245    >>                 vfdlscan                  21-OCT-93 22:39:58
931021_205245    >>                     vfdlscan              21-OCT-93 22:49:55
931021_205245    >>                     vfdlscan              21-OCT-93 22:57:05
931021_230906    >>                                 vfdlscan  22-OCT-93 00:06:56
931021_230906    >>                                 vfdlscan  22-OCT-93 00:20:02
931028_195052    >> vfdlscan                                  28-OCT-93 20:26:01
931104_195727    >> vfdlscan                                   4-NOV-93 22:55:05
931104_195727    >> vfdlscan                                   4-NOV-93 23:10:42
931104_195727    >> vfdlscan                                   4-NOV-93 23:13:05
931104_195727    >> vfdlscan                                   4-NOV-93 23:18:15
931104_195727    >> vfdlscan                                   4-NOV-93 23:20:20
931104_195727    >> vfdlscan                                   4-NOV-93 23:32:14
931104_195727    >> vfdlscan                                   4-NOV-93 23:33:55
931104_195727    >> vfdlscan                                   4-NOV-93 23:35:59
931104_195727    >>     vfdlscan                               4-NOV-93 23:41:28
931104_195727    >>     vfdlscan                               4-NOV-93 23:43:17
931105_195912    >> vfdlscan                                   5-NOV-93 19:59:22
931105_203049    >> vfdlscan                                   5-NOV-93 20:31:03
931106_043318    >>         vfdlscan                           6-NOV-93 07:23:49
931130_094918    >> vfdlscan                                  30-NOV-93 11:54:03
931130_094918    >> vfdlscan                                  30-NOV-93 12:00:15
940224_182543    >>     vfdlscan                              24-FEB-94 22:45:57
940224_182543    >>     vfdlscan                              24-FEB-94 22:51:26
940224_182543    >> vfdlscan                                  24-FEB-94 22:57:25
940224_182543    >> vfdlscan                                  24-FEB-94 23:20:03
XXXXXX_XXXXXX    >>         vfdlscan                          16-MAR-94 02:20:04
XXXXXX_XXXXXX    >>         vfdlscan                          16-MAR-94 02:24:34
XXXXXX_XXXXXX    >>         vfdlscan                          16-MAR-94 02:29:37
XXXXXX_XXXXXX    >>         vfdlscan                          16-MAR-94 02:32:35
940408_213442                   vfdlscan                       8-APR-94 21:44:17
940408_221234                   vfdlscan                       8-APR-94 22:22:06
940409_023735           vfdlscan                               9-APR-94 02:47:06
940409_023735               vfdlscan                           9-APR-94 02:48:51
940409_215413                   vfdlscan                       9-APR-94 22:12:19
940409_215413                       vfdlscan                   9-APR-94 22:14:11
940410_142922           vfdlscan                              10-APR-94 14:46:57
940410_142922           vfdlscan                              10-APR-94 14:48:46
940410_145207           vfdlscan                              10-APR-94 15:09:42
940410_145207           vfdlscan                              10-APR-94 15:11:31

940410_153430           vfdlscan                              10-APR-94 16:02:49
940410_153430           vfdlscan                              10-APR-94 16:04:41
940410_172906                       vfdlscan                  10-APR-94 17:57:34
940410_172906           vfdlscan                              10-APR-94 17:59:24
940410_194930           vfdlscan                              10-APR-94 20:23:26
940410_194930           vfdlscan                              10-APR-94 20:25:18
940410_204949           vfdlscan                              10-APR-94 21:21:10
940410_204949           vfdlscan                              10-APR-94 21:23:02
940410_231043           vfdlscan                              10-APR-94 23:41:12
940410_231043           vfdlscan                              10-APR-94 23:43:04
940411_010818           vfdlscan                              11-APR-94 01:40:51
940411_010818               vfdlscan                          11-APR-94 01:42:43
------------------------------ Searching for mtmtune ------------------------------
931105_232820 <  >> mtmtune                                    5-NOV-93 23:27:54
931105_234456 <  >> mtmtune                                    5-NOV-93 23:44:23
931105_235455 <  >> mtmtune                                    5-NOV-93 23:54:21
931106_000448 <  >> mtmtune                                    6-NOV-93 00:04:18
931106_000448    >> mtmtune                                    6-NOV-93 00:14:29
931106_004551    >>         mtmtune                            6-NOV-93 00:45:28
931106_005437 <  >>         mtmtune                            6-NOV-93 00:53:53
931106_013612 <  >>         mtmtune                            6-NOV-93 01:35:43
931106_014445 <  >>         mtmtune                            6-NOV-93 01:44:11
931106_015624 <  >>     mtmtune                                6-NOV-93 01:54:23
931106_020449 <  >>     mtmtune                                6-NOV-93 02:04:18
931106_021515 <  >>     mtmtune                                6-NOV-93 02:14:42
931106_022629 <  >>     mtmtune                                6-NOV-93 02:25:39
931106_023402 <  >>     mtmtune                                6-NOV-93 02:33:31
931106_024713 <  >>     mtmtune                                6-NOV-93 02:46:31
931106_030239 <  >>     mtmtune                                6-NOV-93 03:02:09
931106_033611 <  >>     mtmtune                                6-NOV-93 03:35:36
931106_034618 <  >>     mtmtune                                6-NOV-93 03:45:35
931106_035446 <  >>     mtmtune                                6-NOV-93 03:54:11
931106_041603 <  >> mtmtune                                    6-NOV-93 04:15:36
931106_042520 <  >> mtmtune                                    6-NOV-93 04:24:51
931106_043318 <  >> mtmtune                                    6-NOV-93 04:32:50
931126_210209 <  >>         mtmtune                           26-NOV-93 21:01:34
931126_230423 <  >> mtmtune                                   26-NOV-93 23:03:52
931126_231551 <  >> mtmtune                                   26-NOV-93 23:15:23
931126_232631 <  >> mtmtune                                   26-NOV-93 23:23:48
931126_233710 <  >> mtmtune                                   26-NOV-93 23:36:34
931126_234525 <  >> mtmtune                                   26-NOV-93 23:43:53
931126_235528 <  >> mtmtune                                   26-NOV-93 23:52:03
931127_000834 <  >> mtmtune                                   27-NOV-93 00:07:57
931127_001658 <  >> mtmtune                                   27-NOV-93 00:15:31
931127_003129    >> mtmtune                                   27-NOV-93 00:30:54
931127_004016 <  >> mtmtune                                   27-NOV-93 00:37:53
931127_004909 <  >> mtmtune                                   27-NOV-93 00:47:17
931127_005846 <  >> mtmtune                                   27-NOV-93 00:55:51
931127_010818 <  >> mtmtune                                   27-NOV-93 01:05:29
931127_011818 <  >> mtmtune                                   27-NOV-93 01:16:36
931127_012814 <  >> mtmtune                                   27-NOV-93 01:27:40
931127_013805    >> mtmtune                                   28-NOV-93 01:37:34
931127_031747    >>     mtmtune                               27-NOV-93 03:15:26
931127_032724 <  >>     mtmtune                               27-NOV-93 03:25:07
931127_035144 <  >>     mtmtune                               27-NOV-93 03:49:09
931127_040155 <  >>     mtmtune                               27-NOV-93 04:00:04
931127_040155    >>     mtmtune                               27-NOV-93 04:08:07
931127_042725 <  >> mtmtune                                   27-NOV-93 04:23:13
931130_173229 <  >> mtmtune                                   30-NOV-93 17:30:52

931130_173229    >> mtmtune                                   30-NOV-93 17:34:37
931130_173658    >> mtmtune                                   30-NOV-93 17:36:15
931130_174943 <  >> mtmtune                                   30-NOV-93 17:47:39
931130_180212 <  >> mtmtune                                   30-NOV-93 18:01:39
931130_181423    >> mtmtune                                   30-NOV-93 18:12:36
931130_181423    >> mtmtune                                   30-NOV-93 18:21:14
931130_182353    >> mtmtune                                   30-NOV-93 18:24:01
931130_183856 <  >> mtmtune                                   30-NOV-93 18:38:23
931130_185106 <  >> mtmtune                                   30-NOV-93 18:50:41
931130_190308 <  >> mtmtune                                   30-NOV-93 19:02:06
931130_190908 <  >> mtmtune                                   30-NOV-93 19:08:40
931130_192229 <  >> mtmtune                                   30-NOV-93 19:21:47
931130_193403 <  >> mtmtune                                   30-NOV-93 19:33:21
931130_194548 <  >> mtmtune                                   30-NOV-93 19:45:06
931130_195705 <  >> mtmtune                                   30-NOV-93 19:56:18
931130_201256 <  >> mtmtune                                   30-NOV-93 20:12:17
931130_202811 <  >> mtmtune                                   30-NOV-93 20:27:30
931130_203918 <  >> mtmtune                                   30-NOV-93 20:38:35
931130_205003 <  >> mtmtune                                   30-NOV-93 20:49:29
931130_210035 <  >> mtmtune                                   30-NOV-93 21:00:03
931130_211400 <  >> mtmtune                                   30-NOV-93 21:13:27
931130_212832 <  >> mtmtune                                   30-NOV-93 21:27:51
931130_214300 <  >> mtmtune                                   30-NOV-93 21:42:15
931130_220201 <  >> mtmtune                                   30-NOV-93 22:01:05
940408_202122 <  >>     mtmtune                                8-APR-94 20:20:42
940408_204844    >>     mtmtune                                8-APR-94 20:47:44
------------------------------ Searching for maplaser ------------------------------
931106_073424    >>         maplaser                           6-NOV-93 07:33:54
931106_101607 <  >> maplaser                                   6-NOV-93 10:14:35
931130_094918    >> maplaser                                  30-NOV-93 15:49:43
------------------------------ Searching for mapgrad ------------------------------
931130_070421    >>     mapgrad                               30-NOV-93 07:03:33
931130_070421    >>     mapgrad                               30-NOV-93 09:19:29
931130_094050    >>     mapgrad                               30-NOV-93 09:39:50
931130_094918    >>     mapgrad                               30-NOV-93 09:48:34
940408_012439    >>             mapgrad                        8-APR-94 01:23:47
940408_021453 <  >>             mapgrad                        8-APR-94 02:13:36
940408_021453    >> mapgrad                                    8-APR-94 05:41:34
------------------------------ Searching for mdiprot ------------------------------
930901_074636 <  >>     mdiprot                                1-SEP-93 07:46:43
931001_221920    >> mdiprot                                    1-OCT-93 22:18:43
931008_030542 <  >>     mdiprot                                8-OCT-93 03:05:06
931008_030542    >>     mdiprot                                8-OCT-93 03:06:02
931223_190425 <  >> mdiprot                                   23-DEC-93 19:02:14
940120_134308 <  >> mdiprot                                   20-JAN-94 13:41:31
940401_180345 <  >>     mdiprot                                1-APR-94 18:01:29
940426_060248 <  >> mdiprot                                   26-APR-94 06:01:36
940505_090234    >> mdiprot                                    5-MAY-94 10:04:11
940505_104648    >>     mdiprot                                5-MAY-94 10:45:45
------------------------------ Searching for mdimrot ------------------------------
930901_080639 <  >>     mdimrot                                1-SEP-93 08:06:43
931001_212119    >> mdimrot                                    1-OCT-93 21:20:56
931008_152008    >>     mdimrot                                8-OCT-93 15:19:39
931223_180435 <  >> mdimrot                                   23-DEC-93 18:03:28
940120_141706 <  >> mdimrot                                   20-JAN-94 14:16:00
940401_183733    >> mdimrot                                    1-APR-94 19:40:56

940401_194654    >> mdimrot                                    1-APR-94 19:45:39
940426_050146 <  >>             mdimrot                       26-APR-94 05:00:33
940505_090234    >>     mdimrot                                5-MAY-94 09:00:27
------------------------------ Searching for mdipdist ------------------------------
930901_070349    >>     mdipdist                               1-SEP-93 07:04:11
930901_071017 <  >>     mdipdist                               1-SEP-93 07:10:29
930901_072035 <  >>     mdipdist                               1-SEP-93 07:21:37
931001_224409    >> mdipdist                                   1-OCT-93 22:43:48
931008_022351 <  >>     mdipdist                               8-OCT-93 02:23:29
931223_193352 <  >> mdipdist                                  23-DEC-93 19:32:37
940401_171701    >>     mdipdist                               1-APR-94 17:15:51
940426_061816 <  >> mdipdist                                  26-APR-94 06:17:34
940504_110947    >> mdipdist                                   4-MAY-94 11:08:33

G Chronological Listing of When Control Files were Run (CAL LOG output)

You can search the file shown below for certain control files by using the alias calgrep. An example is:

% calgrep mdipict

To list all mdipict datasets. If you wanted to search for all mdilaser that were run in Apr-94, you would type:

% calgrep mdilaser | grep APR-94

CAL_LOG.PRO  Ver 2.2
Program Run:  1-Jun-1994 15:19:22.00 on sxt2
  Ser     Telemetry    #Img    Date      Time    Duration    Control File Info  On Disk?   Tape    CCD  Oven
                                                    (min)                        Tlm FITS  #  File
opt347                  206  1-JAN-90  00:00:00 ?   0.     vbigtune               No  No  40   1    ??    ??
opt348                   68  1-JAN-90  00:00:00 ?   0.      vdetune               No  No  40   1    ??    ??
opt013                    3  1-JAN-92  00:00:00 ?   0.       cracks               No  No  -1  -1    ??    ??
opt017                   36 10-NOV-92  19:09:00    46.      fdfocus               No  No   9   5    ??    ??
opt018                   11 10-NOV-92  20:01:00    26.      hrfocus               No  No   5   1    ??    ??
opt022                    4 13-NOV-92  21:36:00     3.       calpos               No  No   6   1    ??    ??
opt024                    4 14-NOV-92  01:17:00     3.       calpos               No  No   6   2    ??    ??
opt028                   45 17-NOV-92  02:06:00    36.  exposerence               No  No   9   2    ??    ??
opt029                   36 17-NOV-92  02:52:00    23.  exposercale               No  No   9   1    ??    ??
opt036                    1 17-NOV-92  16:15:00     0.      dpmtest               No  No   7   1    ??    ??
opt030                   42 17-NOV-92  22:41:00    26.     magdopfd               No  No   9   3    ??    ??
opt034                   16 18-NOV-92  19:57:00    18.    sunlimbhr               No  No  11   3    ??    ??
opt035                   14 18-NOV-92  21:05:00     7.   nilinetune               No  No  11   4    ??    ??
opt030a                  14 19-NOV-92  21:09:00     8.   nilinetune               No  No   9   3    ??    ??
opt031                   14 19-NOV-92  21:22:00     8.   nilinetune               No  No  11   1    ??    ??
opt037                    3 19-NOV-92  21:43:00     1.       irleak               No  No   7   2    ??    ??
opt040                   21 20-NOV-92  02:21:00    11.       nomove               No  No  11   5    ??    ??
opt042                   41 20-NOV-92  14:47:00    18.    sunmagdop               No  No   1   2    ??    ??
opt041                    3 20-NOV-92  22:26:00     1.        sunir               No  No   7   3    ??    ??
opt043                   44 20-NOV-92  23:29:00    18.    sunmagdop               No  No  13   1    ??    ??
opt044                    7 20-NOV-92  23:57:00     3.        sunir               No  No   7   4    ??    ??
opt045                   11 25-NOV-92  00:39:00     5.    covernair               No  No   7   5    ??    ??
opt046                   11 25-NOV-92  00:50:00     5.      covered               No  No   7   6    ??    ??
opt047                   10 25-NOV-92  01:13:00    10.     polcheck               No  No   5   2    ??    ??
opt048                 1695 25-NOV-92  12:45:00     0.    magdopser               No  No   2   1    ??    ??
opt049                  302 29-NOV-92  01:17:00   148.     m1rotser               No  No   9   4    ??    ??
opt050                  305 30-NOV-92  18:29:00   151.     m2rotser               No  No  10   1    ??    ??
opt051                    4  1-DEC-92  23:58:00    17.      windows               No  No   5   3    ??    ??
opt053                 1164  4-DEC-92  12:14:00   213.    sundopser               No  No   1   1    ??    ??
opt052                   19  4-DEC-92  19:38:00     9.    sundopser               No  No  11   6    ??    ??
opt054                    9  7-DEC-92  23:06:00   ****    lamponoff               No  No   5   4    ??    ??
opt059                   14  9-DEC-92  00:14:00    13.       lasdop               No  No   5   5    ??    ??
opt060                  206 10-DEC-92  00:52:00   100.      protser               No  No   1   3    ??    ??

88G CHRONOLOGICAL LISTING OF WHEN CONTROL FILES WERE RUN (CAL LOG OUTPUT)

opt062                  203 10-DEC-92  20:04:00   102.   tunerotser               No  No  11   2    ??    ??
opt063                  843 12-DEC-92  10:26:00   223.    sunobsser               No  No   3   1    ??    ??
opt065                   45 12-DEC-92  23:26:00    13.    sunmagtdt               No  No   5   6    ??    ??
opt066                  129 12-DEC-92  23:49:00    35.    suncaltdt               No  No   8   1    ??    ??
opt067                  867 13-DEC-92  18:51:00   230.    sunmixser               No  No  14   1    ??    ??
opt068                  106 14-DEC-92  19:53:00    32.      lamscan               No  No   1   4    ??    ??
opt069                   56 14-DEC-92  20:41:00    26.       sunmix               No  No   1   5    ??    ??
opt070                  399 14-DEC-92  21:31:00   105.    sunmixser               No  No  13   2    ??    ??
opt071                  724 15-DEC-92  17:05:00   192.    sunmixser               No  No  13   3    ??    ??
opt072                   74 15-DEC-92  20:38:00    23.      lamscan               No  No  12   1    ??    ??
opt073                   48 15-DEC-92  21:05:00    13.      sundop2               No  No  12   2    ??    ??
opt074                   60 15-DEC-92  21:19:00    17.      sundop3               No  No  12   3    ??    ??
opt075                   30 15-DEC-92  21:37:00     8.       sundop               No  No  12   4    ??    ??
opt076                   83 15-DEC-92  22:04:00    28.       sunmix               No  No  12   5    ??    ??
opt077                   45 15-DEC-92  22:37:00    21.  suncalblock               No  No  12   6    ??    ??
opt080                   41 16-DEC-92  17:39:00    62.       sundop               No  No  12   7    ??    ??
opt081                  202 16-DEC-92  18:43:00    83.       sundop               No  No  12   8    ??    ??
opt084                   72 16-DEC-92  23:31:00    51.        darks               No  No  12   9    ??    ??
opt085                   60 17-DEC-92  02:46:00    17.        darks               No  No  12  10    ??    ??
opt062a                   1 23-DEC-92  02:26:00     0.    tdtrotser               No  No  11   2    ??    ??
opt088                   41  1-JAN-93  00:00:00 ?   0.     firstiss               No  No  15   1    ??    ??
opt089                   34  1-JAN-93  00:00:00 ?   0.   protseriss               No  No  15   1    ??    ??
opt090                  177  1-JAN-93  00:00:00 ?   0.   protseriss               No  No  15   1    ??    ??
opt091                   25  1-JAN-93  00:00:00 ?   0.     apr15tst               No  No  15   1    ??    ??
opt092                   10  1-JAN-93  00:00:00 ?   0.     apr15tst               No  No  15   1    ??    ??
opt100                    8  1-JAN-93  00:00:00 ?   0.     lscanobs               No  No  15   1    ??    ??
opt101                   13  1-JAN-93  00:00:00 ?   0.     lscanobs               No  No  15   1    ??    ??
opt102                    7  1-JAN-93  00:00:00 ?   0.       sundop               No  No  15   1    ??    ??
opt103                    7  1-JAN-93  00:00:00 ?   0.    sundopcal               No  No  15   1    ??    ??
opt104                    3  1-JAN-93  00:00:00 ?   0. pfmwindowcal               No  No  15   1    ??    ??
opt105                   13  1-JAN-93  00:00:00 ?   0.     lscancal               No  No  16   1    ??    ??
opt106                    6  1-JAN-93  00:00:00 ?   0.       caldop               No  No  16   1    ??    ??
opt107                    6  1-JAN-93  00:00:00 ?   0.       sundop               No  No  16   1    ??    ??
opt108                    6  1-JAN-93  00:00:00 ?   0.       sundop               No  No  16   1    ??    ??
opt109                    6  1-JAN-93  00:00:00 ?   0.       caldop               No  No  16   1    ??    ??
opt110                   11  1-JAN-93  00:00:00 ?   0.       sunmag               No  No  16   1    ??    ??
opt111                   13  1-JAN-93  00:00:00 ?   0.       sunmag               No  No  16   1    ??    ??
opt112                   14  1-JAN-93  00:00:00 ?   0.     fdfocser               No  No  16   1    ??    ??
opt113                   12  1-JAN-93  00:00:00 ?   0.     hrfocser               No  No  16   1    ??    ??
opt114                    6  1-JAN-93  00:00:00 ?   0.      fdwldop               No  No  16   1    ??    ??
opt115                    6  1-JAN-93  00:00:00 ?   0.   fdwlcaldop               No  No  16   1    ??    ??
opt116                   12  1-JAN-93  00:00:00 ?   0.     fdfocser               No  No  15   1    ??    ??
opt117                   12  1-JAN-93  00:00:00 ?   0.     hrfocser               No  No  15   1    ??    ??
opt118                   12  1-JAN-93  00:00:00 ?   0.     hrfocser               No  No  15   1    ??    ??
opt119                   12  1-JAN-93  00:00:00 ?   0.     fdfocser               No  No  15   1    ??    ??
opt120                   10  1-JAN-93  00:00:00 ?   0.    calexpser               No  No  15   1    ??    ??
opt121                    7  1-JAN-93  00:00:00 ?   0.    sundopiss               No  No  16   1    ??    ??
opt122                    9  1-JAN-93  00:00:00 ?   0.    sunmagiss               No  No  16   1    ??    ??
opt123                   12  1-JAN-93  00:00:00 ?   0.    sunmagiss               No  No  16   1    ??    ??
opt124                   20  1-JAN-93  00:00:00 ?   0.   suncalgrad               No  No  16   1    ??    ??
opt125                    7  1-JAN-93  00:00:00 ?   0.    isspolser               No  No  16   1    ??    ??
opt126                   14  1-JAN-93  00:00:00 ?   0.    isspolser               No  No  16   1    ??    ??
opt130                   71  1-JAN-93  00:00:00 ?   0.     halfwave               No  No  15   1    ??    ??
opt131                   30  1-JAN-93  00:00:00 ?   0.       pretty               No  No  16   1    ??    ??
opt132                   29  1-JAN-93  00:00:00 ?   0.       pretty               No  No  16   1    ??    ??
opt133                   33  1-JAN-93  00:00:00 ?   0.       pretty               No  No  16   1    ??    ??
opt134                   25  1-JAN-93  00:00:00 ?   0.     halfwave               No  No  16   1    ??    ??
opt135                   84  1-JAN-93  00:00:00 ?   0.     fddetune               No  No  16   1    ??    ??
opt137                  114  1-JAN-93  00:00:00 ?   0.    sunmixser               No  No  15   1    ??    ??
opt138                  113  1-JAN-93  00:00:00 ?   0.    sunmixser               No  No  15   1    ??    ??

opt139                  104  1-JAN-93  00:00:00 ?   0.    sunmixser               No  No  16   1    ??    ??
opt140                  113  1-JAN-93  00:00:00 ?   0.    sunmixser               No  No  16   1    ??    ??
opt141                    6  1-JAN-93  00:00:00 ?   0.     newlscan               No  No  16   1    ??    ??
opt142                   33  1-JAN-93  00:00:00 ?   0.     newlscan               No  No  16   1    ??    ??
opt143                   91  1-JAN-93  00:00:00 ?   0.    sunmixsum               No  No  16   1    ??    ??
opt144                   51  1-JAN-93  00:00:00 ?   0.    sunmixsum               No  No  16   1    ??    ??
opt145                  127  1-JAN-93  00:00:00 ?   0.  sunmixsumdt               No  No  17   1    ??    ??
opt146                  129  1-JAN-93  00:00:00 ?   0.    newmixsum               No  No  17   1    ??    ??
opt147                  122  1-JAN-93  00:00:00 ?   0.     newlscan               No  No  17   1    ??    ??
opt148                  251  1-JAN-93  00:00:00 ?   0.        cycle               No  No  17   1    ??    ??
opt149                  122  1-JAN-93  00:00:00 ?   0.     newlscan               No  No  17   1    ??    ??
opt152                   50  1-JAN-93  00:00:00 ?   0.     newlscan               No  No  17   1    ??    ??
opt151                   54  1-JAN-93  00:00:00 ?   0.    newdetune               No  No  17   1    ??    ??
opt150                   81  1-JAN-93  00:00:00 ?   0.     newlscan               No  No  17   1    ??    ??
opt153                   62  1-JAN-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt154                   70  1-JAN-93  00:00:00 ?   0.      openbox               No  No  17   1    ??    ??
opt155                    6  1-JAN-93  00:00:00 ?   0.      openbox               No  No  17   1    ??    ??
opt156                   34  1-JAN-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt086                  234  3-FEB-93  23:23:00   116.     m2rotser               No  No   8   2    ??    ??
opt087                  189  4-FEB-93  18:50:00   154.     m1rotser               No  No   8   3    ??    ??
opt171                   76  1-MAY-93  00:00:00 ?   0.        hrmag               No  No  17   1    ??    ??
opt169                   95  1-MAY-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt168                   17  1-MAY-93  00:00:00 ?   0.        hrdop               No  No  17   1    ??    ??
opt170                  165  1-MAY-93  00:00:00 ?   0.       mixsum               No  No  18   1    ??    ??
opt161                   33  8-MAY-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt163                   34  8-MAY-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt162                    5  8-MAY-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt164                    8  8-MAY-93  00:00:00 ?   0.     newlscan               No  No  17   1    ??    ??
opt166                  120  8-MAY-93  00:00:00 ?   0.       mixsum               No  No  17   1    ??    ??
opt167                  236  8-MAY-93  00:00:00 ?   0.       mixsum               No  No  18   1    ??    ??
opt160                   21  8-MAY-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt165                   76  8-MAY-93  00:00:00 ?   0.     newlscan               No  No  17   1    ??    ??
opt159                  240  8-MAY-93  00:00:00 ?   0.       mixsum               No  No  17   1    ??    ??
opt158                   30  8-MAY-93  00:00:00 ?   0.       detune               No  No  17   1    ??    ??
opt157                   20  8-MAY-93  00:00:00 ?   0.      fdlscan               No  No  17   1    ??    ??
opt157a                  47  8-MAY-93  00:00:00 ?   0.     newlscan               No  No  17   1    ??    ??
opt172                    4 11-MAY-93  14:01:00     1.       detune               No  No  18   1    ??    ??
opt173                    4 11-MAY-93  14:03:00     2.       detune               No  No  18   1    ??    ??
opt174                   22 11-MAY-93  19:34:00    10.     sptfdfoc               No  No  18   1    ??    ??
opt175                   37 11-MAY-93  20:19:00    14.     sptfdfoc               No  No  18   1    ??    ??
opt176                   39 11-MAY-93  20:34:00    19.     spthrfoc               No  No  18   1    ??    ??
opt177                   24 11-MAY-93  21:06:00     4.     sptfdlsn               No  No  18   1    ??    ??
opt178                   26 11-MAY-93  21:11:00     0.     spthrlsn               No  No  18   1    ??    ??
opt179                   30 11-MAY-93  21:34:00    18.     sptccdlt               No  No  18   1    ??    ??
opt180                   26 11-MAY-93  23:13:00     7.       sptlsn               No  No  18   1    ??    ??
opt181                   20 11-MAY-93  23:21:00     4.     spthrlsn               No  No  18   1    ??    ??
opt182                   24 11-MAY-93  23:26:00     4.     spthrlsn               No  No  18   1    ??    ??
opt183                    9 11-MAY-93  23:55:00     4.        fdfoc               No  No  18   1    ??    ??
opt185                   11 12-MAY-93  00:00:00 ?   0.        fdfoc               No  No  18   1    ??    ??
opt186                    9 12-MAY-93  00:00:00 ?   0.        hrfoc               No  No  18   1    ??    ??
opt184                   11 12-MAY-93  00:00:00     7.        hrfoc               No  No  18   1    ??    ??
opt187                   36 12-MAY-93  17:43:00    13.     sptfdfoc               No  No  18   1    ??    ??
opt188                   36 12-MAY-93  17:57:00    18.     spthrfoc               No  No  18   1    ??    ??
opt189                   24 12-MAY-93  18:48:00     4.     sptfdlsn               No  No  18   1    ??    ??
opt190                   26 12-MAY-93  18:54:00     7.     spthrlsn               No  No  18   1    ??    ??
opt191                    9 12-MAY-93  20:13:00     3.        hrfoc               No  No  18   1    ??    ??
opt192                   12 12-MAY-93  20:17:00     8.        fdfoc               No  No  18   1    ??    ??
opt193                   13 12-MAY-93  20:50:00     5.      sptprot               No  No  18   1    ??    ??
opt194                   26 12-MAY-93  21:01:00    10.      sptprot               No  No  18   1    ??    ??
opt195                   41 12-MAY-93  21:33:00    16.    sptmtmrot               No  No  18   1    ??    ??

opt196                   11 12-MAY-93  22:17:00    12.    issoffset               No  No  18   1    ??    ??
opt197                   11 12-MAY-93  23:29:00     4.   laserlscan               No  No  20   1    ??    ??
opt197a                  28 13-MAY-93  00:07:00    13.    laserscan               No  No  20   1    ??    ??
opt198                   39 13-MAY-93  00:22:00    19.      ldetune               No  No  20   1    ??    ??
opt199                   12 13-MAY-93  22:22:00     2.    laserscan               No  No  20   1    ??    ??
opt200                   12 13-MAY-93  22:26:00     2.  laserhrscan               No  No  20   1    ??    ??
opt201                   42 13-MAY-93  22:39:00     9.      ldetune               No  No  20   1    ??    ??
opt202                   20 13-MAY-93  23:19:00     5.     fdsundop               No  No  20   1    ??    ??
opt203                   14 13-MAY-93  23:25:00     3.     sunlscan               No  No  20   1    ??    ??
opt204                   12 13-MAY-93  23:40:00     6.     fdsunmag               No  No  20   1    ??    ??
opt205                   16 14-MAY-93  15:56:00     3.        seq1_No  No  20   1    ??    ??
opt206                   67 14-MAY-93  20:50:00    13.    laserscan               No  No  20   1    ??    ??
opt207                   77 14-MAY-93  21:12:00    19.       detune               No  No  20   1    ??    ??
opt208                   26 14-MAY-93  21:32:00     6.     sunlscan               No  No  20   1    ??    ??
opt209                   26 14-MAY-93  22:23:00    17.       wlscan               No  No  20   1    ??    ??
opt210                   52 18-MAY-93  19:09:00    11.     sptfdlsn               No  No  18   1    ??    ??
opt211                   21 18-MAY-93  19:25:00     9.       sptfoc               No  No  18   1    ??    ??
opt212                   42 19-MAY-93  20:37:00     0.       sptfoc               No  No  18   1    ??    ??
opt213                   31 19-MAY-93  21:10:00    20.     sptccdlt               No  No  20   1    ??    ??
opt214                   13 20-MAY-93  01:06:00     3.     laslscan               No  No  20   1    ??    ??
opt215                  127 20-MAY-93  14:36:00    55.       detune               No  No  20   1    ??    ??
opt217                   43 20-MAY-93  17:54:00   133.   vacldetune               No  No  18   1    ??    ??
opt218                   40 20-MAY-93  20:53:00    13.   vacldetune               No  No  18   1    ??    ??
opt220                   33 20-MAY-93  21:00:00 ?   0. sunvacdetune               No  No  18   1    ??    ??
opt221                   14 20-MAY-93  21:00:00 ?   0.  sunvaclscan               No  No  18   1    ??    ??
opt219                   44 20-MAY-93  21:07:00    17.   vacldetune               No  No  18   1    ??    ??
opt222                   21 20-MAY-93  21:24:00    78.    vacsundop               No  No  18   1    ??    ??
opt216                   43 20-MAY-93  22:29:00     0.   vacldetune               No  No  20   1    ??    ??
opt223                   27 21-MAY-93  00:32:00   112.       sundop               No  No  23   1    ??    ??
opt224                   77 21-MAY-93  14:24:00    14.      vmixser               No  No  23   1    ??    ??
opt225                   88 21-MAY-93  14:39:00    20.      vmixser               No  No  23   1    ??    ??
opt226                   30 21-MAY-93  15:04:00     5.      vdetune               No  No  19   1    ??    ??
opt227                   45 21-MAY-93  15:11:00    18.     vfdlscan               No  No  19   1    ??    ??
opt228                   32 21-MAY-93  15:30:00     6.      vdetune               No  No  19   1    ??    ??
opt229                   42 21-MAY-93  16:48:00     7.     vldetune               No  No  19   1    ??    ??
opt230                   42 21-MAY-93  16:58:00     8.     vldetune               No  No  19   1    ??    ??
opt231                   42 21-MAY-93  17:13:00     8.     vldetune               No  No  19   1    ??    ??
opt232                   37 21-MAY-93  17:58:00    10.     vldetune               No  No  19   1    ??    ??
opt233                   67 21-MAY-93  20:19:00    84.     issonoff               No  No  19   1    ??    ??
opt234                   42 21-MAY-93  23:12:00    16.   protseriss               No  No  23   1    ??    ??
opt235                   37 21-MAY-93  23:55:00    10.      vdetune               No  No  19   1    ??    ??
opt236                   67 22-MAY-93  00:10:00    21.      vmixser               No  No  23   1    ??    ??
opt237                   40 22-MAY-93  14:19:00    14.     vldetune               No  No  20   1    ??    ??
opt238                   33 22-MAY-93  15:00:00     6.      vdetune               No  No  20   1    ??    ??
opt239                   86 22-MAY-93  15:17:00    23.      vmixser               No  No  20   1    ??    ??
opt240                  186 22-MAY-93  15:49:00    32.      vdopser               No  No  20   1    ??    ??
opt241                   45 22-MAY-93  17:01:00    20.     vldetune               No  No  20   1    ??    ??
opt242                   47 22-MAY-93  17:44:00     8.     vldetune               No  No  20   1    ??    ??
opt243                   43 22-MAY-93  18:06:00    13.     vldetune               No  No  20   1    ??    ??
opt244                   42 22-MAY-93  18:51:00     7.     vldetune               No  No 109   4    ??    ??
opt245                   42 22-MAY-93  20:03:00     8.     vldetune               No  No 109   4    ??    ??
opt246                  196 22-MAY-93  22:01:00    64.      vdetune               No  No  20   1    ??    ??
opt247                   44 25-MAY-93  04:43:00    11.     vldetune               No  No 109   4    ??    ??
opt248                   43 25-MAY-93  04:55:00     9.     vldetune               No  No 109   4    ??    ??
opt249                   43 25-MAY-93  05:33:00    14.     vldetune               No  No 109   4    ??    ??
opt250                   43 25-MAY-93  06:06:00     8.     vldetune               No  No 109   4    ??    ??
opt251                   43 25-MAY-93  07:10:00     9.     vldetune               No  No 109   4    ??    ??
opt252                   43 25-MAY-93  08:26:00     9.     vldetune               No  No 109   4    ??    ??
opt253                   43 25-MAY-93  09:10:00    10.     vldetune               No  No 109   4    ??    ??
opt254                   32 25-MAY-93  12:27:00     7.     vldetune               No  No 109   4    ??    ??

opt255                   41 25-MAY-93  12:38:00    21.     vldetune               No  No 109   4    ??    ??
opt256                   42 25-MAY-93  13:03:00    11.     vldetune               No  No 109   4    ??    ??
opt257                   44 25-MAY-93  13:25:00    10.     vldetune               No  No 109   4    ??    ??
opt258                   42 25-MAY-93  14:03:00     8.     vldetune               No  No 109   4    ??    ??
opt259                   42 25-MAY-93  14:13:00     8.     vldetune               No  No 109   4    ??    ??
opt260                   42 25-MAY-93  14:28:00     8.     vldetune               No  No 109   4    ??    ??
opt261                   42 25-MAY-93  14:41:00     8.     vldetune               No  No 109   4    ??    ??
opt262                   42 25-MAY-93  15:02:00     9.     vldetune               No  No 109   4    ??    ??
opt263                   42 25-MAY-93  15:26:00     9.     vldetune               No  No 109   4    ??    ??
opt264                   42 25-MAY-93  16:26:00     8.     vldetune               No  No 109   4    ??    ??
opt265                   42 25-MAY-93  16:37:00     8.     vldetune               No  No 109   4    ??    ??
opt266                   43 25-MAY-93  17:39:00     9.     vldetune               No  No 109   4    ??    ??
opt267                   42 25-MAY-93  18:04:00     8.     vldetune               No  No 109   4    ??    ??
opt268                   43 25-MAY-93  19:14:00    10.     vldetune               No  No 109   4    ??    ??
opt269                   42 25-MAY-93  21:24:00     8.     vldetune               No  No 109   4    ??    ??
opt270                   42 25-MAY-93  22:52:00     8.     vldetune               No  No 109   4    ??    ??
opt271                   44 26-MAY-93  00:17:00     8.     vldetune               No  No 109   4    ??    ??
opt272                   57 26-MAY-93  13:05:00     0.     vldetune               No  No 109   4    ??    ??
opt273                   46 26-MAY-93  14:37:00     9.     vldetune               No  No  20   1    ??    ??
opt274                   20 26-MAY-93  17:42:00    10.       sptfoc               No  No  20   1    ??    ??
opt275                   26 26-MAY-93  18:01:00    15.     sptlscan               No  No  20   1    ??    ??
opt276                   12 26-MAY-93  22:25:00    18.        align               No  No  20   1    ??    ??
opt277                   79 27-MAY-93  17:53:00    83.       djaiqt               No  No  20   1    ??    ??
opt278                   42 28-MAY-93  00:42:00    17.     m1rotser               No  No  20   1    ??    ??
opt279                   31 28-MAY-93  01:08:00    13.     m2rotser               No  No  20   1    ??    ??
opt280                   29 29-MAY-93  13:26:00     9.       cdpser               No  No  21   1    ??    ??
opt281                  208 29-MAY-93  13:38:00    36.      vdopser               No  No  21   1    ??    ??
opt282                   29 29-MAY-93  14:23:00    11.      cdopser               No  No  21   1    ??    ??
opt283                 1128 29-MAY-93  14:37:00   225.      vdopser               No  No  21   1    ??    ??
opt284                  752 29-MAY-93  18:31:00   145.      vdopser               No  No  21   1    ??    ??
opt285                   10  2-JUN-93  01:03:00     3.    grandhunt               No  No  21   1    ??    ??
opt286                   11  2-JUN-93  01:23:00     8.        roger               No  No  21   1    ??    ??
opt287                   30  2-JUN-93  15:04:00     6.      cdopser               No  No  21   1    ??    ??
opt288                   74  2-JUN-93  15:12:00    15.      vdopser               No  No  21   1    ??    ??
opt289                   13  2-JUN-93  17:28:00    24.      isstest               No  No  21   1    ??    ??
opt290                   82  2-JUN-93  22:14:00    32.       sptfoc               No  No  21   1    ??    ??
opt291                   50  2-JUN-93  22:48:00    20.      sptprot               No  No  21   1    ??    ??
opt292                   83  2-JUN-93  23:10:00    40.      sptmrot               No  No  21   1    ??    ??
opt293                   26  3-JUN-93  00:01:00     6.   sptfdlscan               No  No  21   1    ??    ??
opt294                   32  3-JUN-93  00:32:00    25.   sunisstest               No  No  21   1    ??    ??
opt295                   25  3-JUN-93  20:36:00     8.       wlscan               No  No  22   1    ??    ??
opt296                   26  3-JUN-93  20:54:00    11.     sunlscan               No  No  22   1    ??    ??
opt297                   32  3-JUN-93  21:48:00   119.     holemove               No  No  22   1    ??    ??
opt298                   33  4-JUN-93  20:04:00    39.      polhunt               No  No  22   1    ??    ??
opt299                   43  5-JUN-93  00:06:00     9.     vldetune               No  No  22   1    ??    ??
opt300                   10  5-JUN-93  00:48:00    11.     michtest               No  No  22   1    ??    ??
opt301                   26  7-JUN-93  19:40:00    15.      lfringe               No  No  22   1    ??    ??
opt302                   19  7-JUN-93  21:56:00    31.        qwave               No  No  22   1    ??    ??
opt303                   52  7-JUN-93  23:45:00   ****        qwave               No  No  22   1    ??    ??
opt304                   28  8-JUN-93  21:33:00    44.       finger               No  No  22   1    ??    ??
opt305                   10  8-JUN-93  23:54:00    40.        bdsbs               No  No  22   1    ??    ??
opt306                   51  9-JUN-93  02:50:00    17.     pawlscan               No  No  23   1    ??    ??
opt307                   12  9-JUN-93  17:51:00     9. beforecamera               No  No  23   1    ??    ??
opt309                  119  9-JUN-93  22:05:00   165.   diode_scan               No  No  22   1    ??    ??
opt308                   12  9-JUN-93  30:31:00     0.  aftercamera               No  No  23   1    ??    ??
opt310                   36 10-JUN-93  17:24:00    29. ccd_sat_test               No  No  23   1    ??    ??
opt311                   16 10-JUN-93  18:01:00    23.   window_off               No  No  22   1    ??    ??
opt312                    7 10-JUN-93  18:46:00     2.  window_stim               No  No  22   1    ??    ??
opt313                   31 11-JUN-93  02:46:00    77.      bs1_bbx               No  No  23   1    ??    ??
opt314                   27 18-JUN-93  00:00:00 ?   0.    oven_move               No  No  23   1    ??    ??

opt315                   11 21-JUN-93  23:39:00     0.      bs2_bbx               No  No  23   1    ??    ??
opt316                  100 22-JUN-93  17:34:00   359.      bs2_bbx               No  No  23   1    ??    ??
opt317                   38 25-JUN-93  01:48:00    14.      two_bit               No  No  23   1    ??    ??
opt318                   13 25-JUN-93  03:12:00     8.     sunblock               No  No  23   1    ??    ??
opt319                   58 26-JUN-93  00:55:00    30.      twobits               No  No  23   1    ??    ??
opt320                   44 26-JUN-93  02:41:00    12.     vldetune               No  No  23   1    ??    ??
cat000   930629_171246    5 29-JUN-93  17:17:00     1.                           Yes  No 115   2    ??    ??
cat001   930629_204340    3 29-JUN-93  20:44:00     0.                           Yes  No 109   2    ??    ??
cat004   930702_222529    1  2-JUL-93  22:25:29 ?   0.                  swldseq1 Yes  No  -1  -1    ??    ??
cat005   930702_230529    1  2-JUL-93  23:05:29 ?   0.                  swldseq1 Yes  No  -1  -1    ??    ??
cat011   930703_184101   15  3-JUL-93  18:42:00     7.                        tp Yes  No 115   2    ??    ??
cat012   930703_190236  189  3-JUL-93  19:04:00    49.                  swldseq2 Yes  No 115   2    ??    ??
cat013   930703_204142  820  3-JUL-93  20:49:00   163.                           Yes  No 115   2    ??    ??
opt321                   41  3-JUL-93  21:14:00    12.      vdetune               No  No  24   1    ??    ??
opt322                   27  3-JUL-93  21:28:00     5.      fdlscan               No  No  24   1    ??    ??
opt323                   30  4-JUL-93  15:02:00     5.      vdetune               No  No  25   1    ??    ??
opt324                   26  4-JUL-93  15:11:00     9.      fdlscan               No  No  25   1    ??    ??
cat014   930704_153134   90  4-JUL-93  15:32:00    18.                           Yes  No 115   3    ??    ??
cat015   930704_163051    3  4-JUL-93  16:32:00     0.                           Yes  No 115   3    ??    ??
cat016   930704_163310   39  4-JUL-93  16:34:00     7.                           Yes  No 115   3    ??  40.7
cat018   930704_184933   81  4-JUL-93  18:51:00    16.                           Yes  No 115   3    ??    ??
opt325                  105  4-JUL-93  20:14:00    20.          sun               No  No  25   1    ??    ??
opt326                  945  4-JUL-93  20:58:00   204.          sun               No  No  25   1    ??    ??
opt327                   68  5-JUL-93  01:13:00    19.          sun               No  No  25   1    ??    ??
opt328                  323  5-JUL-93  01:39:00    70.          sun               No  No  25   1    ??    ??
opt329                   57  5-JUL-93  03:00:00    11.          sun               No  No  25   1    ??    ??
cat019   930705_142958 3376  5-JUL-93  14:31:00   686.                           Yes  No 109   9    ??    ??
opt330                 3491  5-JUL-93  15:31:00   700.          sun               No  No  26   1    ??    ??
opt331                    8  6-JUL-93  03:16:00     2.          ccd               No  No  27   1    ??    ??
opt333                 2256  6-JUL-93  17:55:00   456.          sun               No  No  27   1    ??    ??
opt334                  330  7-JUL-93  01:34:00    64.          sun               No  No  27   1    ??    ??
opt336                  274  9-JUL-93  23:31:00    86.        focus               No  No  28   1    ??    ??
cat023   930710_155702 3103 10-JUL-93  15:58:00   607.                     pdown  No  No 109   3    ??    ??
opt337                 2979 10-JUL-93  15:59:00   606.       caldop               No  No  29   1    ??    ??
opt338                 6148 11-JUL-93  18:47:00   ****        darks               No  No  41   1    ??    ??
cat024   930711_184559 3355 11-JUL-93  18:47:00   671.                           Yes  No 115   1    ??    ??
opt339                   36 14-JUL-93  14:19:00    17.     fdfocser               No  No  39   1  56.0  32.9
opt340                  205 14-JUL-93  14:41:00    40.     vbigtune               No  No  39   1  56.0  32.9
opt341                  244 14-JUL-93  16:12:00    51.       depseq               No  No  39   1    ??  30.7
cat025   930714_161132  240 14-JUL-93  16:13:00    51.                  vbigtune  No  No 115   2    ??  30.7
opt342                  302 15-JUL-93  00:08:00    64.     vbigtune               No  No  39   1    ??  33.2
opt343                  661 15-JUL-93  14:03:00    50.      vmagser               No  No  39   1    ??  40.7
opt344                  144 15-JUL-93  15:05:00    28.      vmagser               No  No  39   1    ??  40.7
opt345                  372 16-JUL-93  01:01:00    26.      vdopmix               No  No  39   1    ??  35.1
opt346                  565 16-JUL-93  01:44:00    38.      vdopmix               No  No  39   1    ??  35.0
opt351                   22 19-JUL-93  22:32:00     6.     fdfocser               No  No  40   1    ??    ??
opt352                   20 19-JUL-93  23:00:00     4.     fdfocser               No  No  40   1    ??    ??
opt353                   24 19-JUL-93  23:56:00     7.     fdfocser               No  No  40   1    ??    ??
opt354                   24 20-JUL-93  00:23:00    10.     fdfocser               No  No  40   1    ??    ??
opt355                   24 20-JUL-93  22:48:00   102.       finger               No  No  43   1    ??    ??
opt356                   21 21-JUL-93  15:00:00    19.        darks               No  No  43   1    ??  40.7
opt357                   58 28-JUL-93  02:10:00    10.     fdfocrot               No  No  45   1    ??  25.6
opt358                   58 28-JUL-93  02:21:00     9.     hrfocrot               No  No  52   1    ??  25.6
opt359                   64 28-JUL-93  02:36:00    11.       pawrot               No  No  52   1    ??  25.6
opt360                   64 28-JUL-93  02:47:00    11.       pawrot               No  No  45   1    ??  25.6
opt361                  248 28-JUL-93  03:28:00    44.       mtmrot               No  No  -1  -1    ??  25.6
cat026   930729_200816  100 29-JUL-93  20:10:00    19.                   shutest Yes  No 115   1  26.7  25.6
opt407                  370 31-JUL-93  00:00:00 ?   0. qwavesnholes               No  No  81   1 -11.4  35.0
opt362                  104 31-JUL-93  14:03:00    10.    lasdetune               No  No  43   1 -48.6  35.2
cat027   930731_144820  345 31-JUL-93  14:50:00    86.                  vbigtune Yes  No 115   4 -48.6  35.2

opt363                  435 31-JUL-93  15:00:00    73.       suncal               No Yes  43   1 -48.6  35.2
opt364                  106 31-JUL-93  16:28:00    11.    lasdetune               No Yes  43   1 -48.6  35.2
opt365                  432 31-JUL-93  16:45:00    72.       suncal               No Yes 109   4 -48.6  35.2
cat028   930731_164519    1 31-JUL-93  16:45:19 ?   0.                  vbigtune Yes  No  -1  -1 -48.6  35.2
opt366                  106 31-JUL-93  18:12:00    11.    lasdetune               No  No  43   1 -48.6  35.1
opt367                  432 31-JUL-93  18:32:00    74.       suncal               No  No  43   1 -48.6  35.2
cat029   930731_183233  340 31-JUL-93  18:32:00    76.                  vbigtune Yes  No 115   4 -48.6  35.2
opt368                  106 31-JUL-93  19:55:00    10.    lasdetune               No  No  45   1 -48.6  35.2
opt369                  144 31-JUL-93  20:18:00     7.       suncal               No  No  45   1 -48.6  35.2
opt370                   20 31-JUL-93  20:29:00     1.       suncal               No  No  45   1 -48.6  35.1
opt371                   33 31-JUL-93  20:38:00     3.       suncal               No  No  45   1 -48.6  35.1
opt372                  430 31-JUL-93  20:44:00    73.       suncal               No  No  45   1 -45.7  35.1
cat032   930731_203902  335 31-JUL-93  20:54:00    66.                  calmodev Yes  No 115   3 -45.7  35.2
opt373                  140 31-JUL-93  22:12:00    31.    lasdetune               No  No  46   1 -45.7  35.2
opt374                  432 31-JUL-93  22:51:00    72.       suncal               No  No  43   1 -45.7  35.1
cat033   930731_225227  340 31-JUL-93  22:54:00    73.                           Yes  No 115   4 -48.6  35.0
opt408                  590  1-AUG-93  00:00:00 ?   0.    waveplate               No  No  81   1 -48.6  35.2
opt409                  103  1-AUG-93  00:00:00 ?   0.      bigtune               No  No  81   1 -48.6  35.2
opt375                  133  1-AUG-93  00:36:00    25.    lasdetune               No  No  46   1 -48.6  35.2
opt376                  106  1-AUG-93  01:18:00    10.    lasdetune               No  No  46   1 -48.6  35.2
opt377                  318  1-AUG-93  01:33:00    49.       suncal               No  No  45   1 -48.6  35.1
cat034   930801_013337  220  1-AUG-93  01:35:00    50.                  calmodev Yes  No 115   4 -48.6  35.2
opt378                  101  1-AUG-93  13:04:00    32.        darks               No  No  46   1 -48.6  35.2
opt379                    3  1-AUG-93  13:40:00     1.        darks               No  No  43   1 -48.6  35.2
opt380                   22  1-AUG-93  14:16:00     6.          mag               No  No  43   1 -45.7  35.2
opt381                   20  1-AUG-93  14:26:00     2.          mag               No  No  43   1 -48.6  35.2
opt382                    8  1-AUG-93  14:33:00     1.          dop               No  No  43   1 -48.6  35.2
opt383                 3193  1-AUG-93  14:45:00   637.          dop               No  No  45   1 -45.7  35.2
cat036   930801_144602 1790  1-AUG-93  14:48:00   357.                  obsmodev Yes  No 115   3 -48.6  35.2
opt384                  104  2-AUG-93  01:29:00    38.          mag               No  No  43   1 -45.7  35.1
opt385                  122  2-AUG-93  02:17:00    40.         door               No  No  45   1 -48.6  35.1
cat037   930802_021910  200  2-AUG-93  02:21:00    39.                     pdown Yes  No 115   1 -48.6  35.2
opt386                   14  2-AUG-93  14:40:00     4.     hrfocser               No  No  43   1 -51.4  35.3
opt387                   18  2-AUG-93  15:11:00     3.          ltc               No  No  46   1 -51.4  35.3
opt400                    9  4-AUG-93  00:00:00 ?   0. sorensondark               No  No  80   1    ??  40.7
opt388                  103  4-AUG-93  01:46:00     9.    lasdetune               No  No  46   1 -45.7  35.2
opt389                  103  4-AUG-93  05:27:00     8.    lasdetune               No  No  46   1 -48.6  35.2
opt390                  103  4-AUG-93  05:44:00     8.    lasdetune               No  No  46   1 -45.7  35.2
opt391                  103  4-AUG-93  06:09:00     7.    lasdetune               No  No  46   1 -48.6  35.2
opt392                  103  4-AUG-93  06:19:00     8.    lasdetune               No  No  46   1 -45.7  35.2
opt393                  103  4-AUG-93  06:30:00     9.    lasdetune               No  No  46   1 -51.4  35.2
opt394                  103  4-AUG-93  13:42:00    10.    lasdetune               No  No  46   1 -51.4  35.3
opt395                   29  4-AUG-93  20:01:00     9.     sunfdfoc               No  No  46   1 -45.7  33.4
opt396                   39  4-AUG-93  20:15:00    10.     sunhrfoc               No  No  46   1 -48.6  35.2
opt397                   19  4-AUG-93  20:30:00     4.     sunfdfoc               No  No  46   1 -45.7  35.5
opt398                   25  4-AUG-93  20:38:00     4.      fdlscan               No  No  46   1 -45.7  35.4
opt399                   25  4-AUG-93  21:32:00     3.      fdlscan               No  No  46   1 -48.6  35.4
opt401                   19  4-AUG-93  23:04:00     5.     fdfocser               No  No  46   1 -45.7  35.3
opt402                    6  6-AUG-93  21:31:00     3.      icehunt               No  No  47   1    ??    ??
opt403                    4  6-AUG-93  21:58:00     1.     issdarks               No  No  47   1  23.3  30.8
opt404                   19  6-AUG-93  22:36:00     5.     hrfocser               No  No  47   1  23.3  31.6
opt405                   11  6-AUG-93  23:05:00     2.     hrfocser               No  No  47   1  23.3  32.0
opt406                   26  7-AUG-93  00:49:00    27.      issperf               No  No  47   1  23.3  33.0
opt410                   65 18-AUG-93  14:45:00   314.       turnon               No  No  49   1    ??    ??
opt411                    4 18-AUG-93  20:10:00    98.          ice               No  No  50   1 -45.7  25.6
opt412                   78 18-AUG-93  23:27:00   137.        focus               No  No  50   1    ??  25.6
opt413                   26 19-AUG-93  02:19:00   261.          ice               No  No  49   1    ??    ??
opt414                    7 19-AUG-93  14:03:00     6.          ice               No  No  51   1 -51.4  25.9
opt415                    9 19-AUG-93  15:55:00     6.          ice               No  No  49   1 -48.6  31.7
opt416                  184 19-AUG-93  19:34:00   151.        focus               No  No  51   1 -48.6  26.7

opt417                   20 19-AUG-93  23:00:00     4.        focus               No  No  52   1 -48.6  25.6
opt418                   29 19-AUG-93  23:33:00    25.          ice               No  No  50   1 -48.6  25.6
opt419                    3 20-AUG-93  01:36:00     2.          ice               No  No  49   1 -45.7  25.6
opt420                   50 20-AUG-93  06:09:00    28.          foc               No  No  81   1 -14.3  25.6
opt421                   10 20-AUG-93  23:21:00     4.    southwall               No  No  49   1    ??    ??
opt422                   77 20-AUG-93  23:30:00    14.      bigtune               No  No  49   1    ??    ??
opt423                  207 21-AUG-93  01:04:00    29.      bigtune               No  No  49   1    ??    ??
opt424                  128 21-AUG-93  20:30:00    38.    waveplate               No  No  50   1    ??    ??
opt425                  409 21-AUG-93  21:24:00    83.    waveplate               No  No  50   1    ??    ??
opt426                   34 23-SEP-93  18:29:00    19.     sunhrfoc               No  No  52   1    ??    ??
opt427                    5 23-SEP-93  18:59:00     1.      bs1test               No  No  52   1    ??    ??
opt428                    5 23-SEP-93  19:55:00     3.      bs1test               No  No  52   1    ??    ??
opt429                    5 23-SEP-93  20:11:00     1.      bs1test               No  No  52   1    ??    ??
opt430                   10 23-SEP-93  20:21:00     5.      bs1test               No  No  52   1    ??    ??
opt431                    6 23-SEP-93  20:33:00     3.      bs1test               No  No  52   1    ??    ??
opt432                    5 23-SEP-93  20:48:00     2.      bs1test               No  No  52   1    ??    ??
opt433                    5 23-SEP-93  20:59:00     2.      bs1test               No  No  52   1    ??    ??
opt434                    5 23-SEP-93  21:15:00     2.      bs1test               No  No  52   1    ??    ??
opt435                   19 24-SEP-93  15:17:00     4.      stimfoc               No  No  80   1    ??    ??
opt436                   82 24-SEP-93  16:06:00    21.     sunhrfoc               No  No  52   1    ??    ??
opt437                   40 24-SEP-93  16:42:00    23.     sunholes               No  No  80   1    ??    ??
oct500   xxxxxx_xxxxxx   49  1-OCT-93  00:00:00 ?   0.       diodes               No  No  75   1    ??    ??
oct503   xxxxxx_xxxxxx   40  1-OCT-93  00:00:00 ?   0.      magneto               No Yes  74   1    ??    ??
oct502   xxxxxx_xxxxxx  121  1-OCT-93  00:00:00 ?   0.      magneto               No Yes  74   1    ??    ??
cat038   931001_152636    6  1-OCT-93  15:27:00     9.                    lcmcom  No Yes  82   1  23.3  25.6
cat039   931001_164601   34  1-OCT-93  16:47:00     6.                    lsqcom  No  No  83   1  23.3  27.6
cat040   931001_194438    7  1-OCT-93  19:45:00    12.                   mdipict  No  No  84   1  23.3  25.6
cat041   931001_200012   48  1-OCT-93  20:00:00    13.                    mdiltc  No  No  85   1  23.3  25.6
cat042   931001_202439   74  1-OCT-93  20:25:00    15.                  mdifocus  No  No  84   1  23.3  25.6
cat043   931001_204141    4  1-OCT-93  20:44:00     2.                    mdiiss  No  No  83   1  23.3  25.6
cat044   931001_205218   27  1-OCT-93  20:53:00    24.                    mdiiss  No  No  83   1  23.3  25.6
cat045   931001_212119  301  1-OCT-93  21:22:00    55.                   mdimrot  No  No  96   1  26.7  25.6
cat046   931001_221920  122  1-OCT-93  22:20:00    22.                   mdiprot  No  No  84   1  23.3  25.6
cat047   931001_224409   21  1-OCT-93  22:44:00    42.                  mdipdist  No  No  84   1  23.3  25.6
cat048   931001_234506  100  1-OCT-93  23:46:00    48.                   mdidust  No  No  85   1  26.7  25.6
opt438                   41  2-OCT-93  19:32:00    11.       sundop               No  No  -1  -1    ??    ??
cat049   931002_211056  101  2-OCT-93  21:11:00    21.                  mdidtune  No  No  86   1  23.3  25.6
cat050   931002_214453  100  2-OCT-93  21:45:00    32.                   mdidust  No  No  86   1  23.3  25.6
cat051   931007_223139    6  7-OCT-93  22:32:00     9.                    lcmcom  No  No  85   1  26.7  25.6
cat052   931007_234103   35  7-OCT-93  23:42:00     6.                    lsqcom  No  No  86   1  26.7  26.9
cat053   931008_010431    7  8-OCT-93  01:05:00    11.                   mdipict  No  No  86   1  26.7  25.6
cat054   931008_011902   48  8-OCT-93  01:19:00    13.                    mdiltc  No  No  85   1  23.3  25.6
cat055   931008_013816   73  8-OCT-93  01:39:00    12.                  mdifocus  No  No  82   1  23.3  25.6
cat056   931008_015736   27  8-OCT-93  01:58:00    21.                    mdiiss  No  No  84   1  23.3  25.6
cat057   931008_022351   21  8-OCT-93  02:24:00    38.                  mdipdist  No  No  87   1  23.3  25.6
cat058   931008_030637  122  8-OCT-93  03:07:00    26.                    mdioff  No  No  86   1  23.3  25.6
cat059   931008_152008  301  8-OCT-93  15:20:00    56.                   shutest  No  No  84   1  26.7  25.6
cat060   931008_201704  100  8-OCT-93  20:22:00    32.                   mdidust Yes  No  84   1  23.3  25.6
cat061   931008_211021  101  8-OCT-93  21:11:00    17.                  mdidtune  No  No  82   1  26.7  25.6
cat062   931012_185700    6 12-OCT-93  18:57:00     9.                    lcmcom  No  No  84   1  23.3  25.6
cat063   931012_210906   29 12-OCT-93  21:10:00     6.                    lsqcom  No  No  84   1  30.0  28.9
cat064   931013_174524    1 13-OCT-93  17:45:24 ?   0.                            No  No  -1  -1    ??  25.6
cat065   931013_181655    1 13-OCT-93  18:17:00     0.                            No  No  84   1    ??  25.6
cat066   931013_182509    1 13-OCT-93  18:25:00     0.                            No  No  84   1    ??  25.6
cat067   931013_183603    1 13-OCT-93  18:36:00     0.                   daveoff  No  No  84   1    ??  25.6
cat068   931013_185517    1 13-OCT-93  18:56:00     0.                            No  No  84   1    ??  25.6
cat069   931013_194318    1 13-OCT-93  19:44:00     0.                  mdiofred  No  No  84   1    ??  25.6
cat070   931014_214929    1 14-OCT-93  21:50:00     0.                            No  No  84   1    ??  25.6
cat071   931014_215820    1 14-OCT-93  21:59:00     0.                   daveoff  No  No  84   1    ??  25.6
cat072   931014_221648    1 14-OCT-93  22:17:00     0.                            No  No  84   1    ??  25.6

cat073   931014_222505    1 14-OCT-93  22:25:00     0.                  mdiofred  No  No  84   1    ??  25.6
cat074   931018_205132   51 18-OCT-93  20:52:00     9.                            No  No  84   1  40.0  38.6
cat075   931018_210208    8 18-OCT-93  21:02:00    14.                   mdipict  No  No  84   1  44.0  38.5
cat076   931019_212049    1 19-OCT-93  21:21:00     0.                   bigtune  No  No  84   1  30.0  35.6
oct501   931019_215619  101 19-OCT-93  21:56:00    17.          xxx     mdidtune  No  No  73   1  30.0  36.1
oct504   931020_214216  101 20-OCT-93  21:42:00    17.          xxx     mdidtune  No  No  74   1 -11.4  35.2
cat077   931021_205245  183 21-OCT-93  21:00:00    36.                  mdidtune  No  No  96   2 -17.1  35.9
oct505   931021_230906  104 21-OCT-93  23:09:00    16.          xxx     mdidtune  No  No  76   1 -14.3  35.9
oct506   931022_002305  104 22-OCT-93  00:23:00    16.          xxx     mdidtune  No  No  77   1  -8.6  34.7
oct507   931022_015725  104 22-OCT-93  01:57:00    16.          xxx     mdidtune  No  No  76   1  -8.6  36.0
cat078   931022_161640   90 22-OCT-93  16:19:00    18.                   obsmode  No  No  96   3 -14.3  33.9
oct508   931022_163919 2190 22-OCT-93  16:41:00   442.          sun               No  No  73   1 -14.3  35.8
oct509   931023_000744  105 23-OCT-93  00:08:00    16.          sun     mdidtune  No  No  74   1 -14.3  36.1
oct510   931023_013033  104 23-OCT-93  01:31:00    15.          xxx     mdidtune  No  No  73   1 -17.1  35.4
oct511   931023_141805  232 23-OCT-93  14:20:00    46.        darks      lfdopen  No  No  89   1 -14.3  35.0
oct512   931023_153723 1666 23-OCT-93  15:39:00   333.       satcal      obsmode  No  No  73   1 -14.3  33.6
oct513   931024_153756 2645 24-OCT-93  15:38:00   529.          sun     closelog  No Yes  76   1 -17.1  35.0
oct514   931025_205717   16 25-OCT-93  20:57:17 ?   0.     compress      openlog Yes Yes  -1  -1 -14.3  35.4
cat079   931025_212304    1 25-OCT-93  21:23:04 ?   0.                  closelog  No  No  -1  -1  56.0  32.9
oct515   931025_231427  388 25-OCT-93  23:15:00    61.          xxx     supersun  No  No  77   1 -14.3  34.8
oct516   xxxxxx_xxxxxx  203 26-OCT-93  18:05:00    22.        drift               No Yes  74   1 -14.3  36.2
oct518   931026_184901    1 26-OCT-93  18:51:00     0.          xxx     calmodev Yes  No  -1  -1 -14.3  35.0
oct517   xxxxxx_xxxxxx   58 26-OCT-93  19:12:00    61.      scatter               No  No  75   1 -14.3  35.0
oct519   xxxxxx_xxxxxx  328 26-OCT-93  20:40:00    26.    supertune               No Yes  90   1 -14.3  35.0
oct520   xxxxxx_xxxxxx  102 26-OCT-93  23:40:00     7.      fringes               No  No  78   1 -14.3  35.0
oct521   xxxxxx_xxxxxx  104 27-OCT-93  00:06:00    35.      fringes               No  No  90   1 -14.3  35.2
oct522   xxxxxx_xxxxxx  444 27-OCT-93  17:04:00    35.       newmix               No  No  79   1 -23.3  35.3
oct523   xxxxxx_xxxxxx  104 27-OCT-93  17:52:00    33.      fringes               No  No  75   1 -20.0  35.7
oct524   xxxxxx_xxxxxx  104 27-OCT-93  18:32:00    33.      fringes               No  No  75   1 -23.3  33.1
oct524a  931027_195942   20 27-OCT-93  20:00:00     5.          xxx               No  No  79   1 -20.0  35.3
oct524b  931027_201301   10 27-OCT-93  20:13:00     3.          xxx     lltsetup  No  No  79   1 -20.0  34.8
oct524c  931027_202241    9 27-OCT-93  20:23:00     2.          xxx     lltsetup  No  No  74   1 -23.3  34.9
oct524d  931027_202633   15 27-OCT-93  20:27:00     4.          xxx     lltsetup  No  No  79   1 -20.0  35.0
oct524e  931027_203119    5 27-OCT-93  20:32:00     0.          xxx               No  No  74   1 -20.0  35.0
oct524f  931027_203724   10 27-OCT-93  20:38:00     2.          xxx      bigtune  No  No  74   1 -23.3  34.9
oct525   xxxxxx_xxxxxx  454 27-OCT-93  21:39:00    43.       newmxr               No  No  79   1 -20.0  35.4
oct526   931027_233428  104 27-OCT-93  23:35:00    15.          xxx     mdidtune  No  No  75   1 -20.0  34.5
oct526a  931028_002639  104 28-OCT-93  00:27:00    16.          xxx     mdidtune  No  No  77   1 -20.0  35.7
oct527   931028_155016  104 28-OCT-93  15:50:00    16.          xxx     mdidtune  No  No  75   1 -20.0  35.2
oct528   931028_183039  102 28-OCT-93  18:31:00    16.          xxx     mdidtune  No  No  77   1 -23.3  34.8
oct528a  931028_190116   90 28-OCT-93  19:01:00    19.          xxx     mdidtune  No  No  75   1 -17.1  35.0
oct529   931028_195052  104 28-OCT-93  19:51:00    16.          xxx     mdidtune  No  No  77   1 -20.0  34.3
oct530   xxxxxx_xxxxxx    8 28-OCT-93  20:45:00     1.      lasdust               No  No  79   1 -20.0  35.2
oct531   931028_212849  104 28-OCT-93  21:29:00    16.          xxx     mdidtune  No  No  78   1 -20.0  35.3
oct532   xxxxxx_xxxxxx    8 28-OCT-93  21:58:00     2.      lasdust               No  No  79   1    ??    ??
oct533   931028_223530  104 28-OCT-93  22:36:00    15.          xxx     mdidtune  No  No  75   1 -20.0  35.9
oct534   xxxxxx_xxxxxx    8 28-OCT-93  23:27:00     0.      lasdust               No  No  76   1 -20.0  35.0
oct535   931028_234149  104 28-OCT-93  23:42:00    16.          xxx     mdidtune  No  No  77   1 -20.0  35.6
oct536   xxxxxx_xxxxxx    9 29-OCT-93  00:08:00    39.      lasdust               No  No  76   1 -20.0  34.0
oct537   931029_014038  104 29-OCT-93  01:41:00    16.          xxx     mdidtune  No  No  78   1    ??    ??
oct538   xxxxxx_xxxxxx    8 29-OCT-93  02:02:00     1.      lasdust               No  No  75   1 -20.0  35.4
oct539   931029_021639  104 29-OCT-93  02:17:00    17.          xxx     mdidtune  No  No  75   1 -23.3  34.7
oct540   xxxxxx_xxxxxx    8 29-OCT-93  02:38:00     1.      lasdust               No  No  76   1 -23.3  35.6
oct541   931029_024958  104 29-OCT-93  02:50:00    16.          xxx     mdidtune  No  No  77   1    ??    ??
oct542   xxxxxx_xxxxxx    8 29-OCT-93  03:19:00     1.      lasdust               No  No  75   1 -23.3  34.0
oct543   931029_034842  104 29-OCT-93  03:49:00    16.          xxx     mdidtune  No  No  78   1 -23.3  34.0
oct544   xxxxxx_xxxxxx    8 29-OCT-93  04:06:00     1.      lasdust               No  No  76   1 -20.0  35.6
oct545   931029_041446  104 29-OCT-93  04:15:00    19.          xxx     mdidtune  No  No  75   1 -23.3  35.2
oct546   xxxxxx_xxxxxx    8 29-OCT-93  04:36:00     0.      lasdust               No  No  75   1 -23.3  34.5
oct547   xxxxxx_xxxxxx  283 29-OCT-93  04:36:00    39.    supertune               No  No  75   1 -23.3  34.5

oct548   xxxxxx_xxxxxx  658 29-OCT-93  15:56:00    55.      newmixv               No  No  79   1 -23.3  35.8
oct549   931029_173353  104 29-OCT-93  17:34:00    16.     mdidtune     mdidtune  No  No  77   1 -26.7  35.7
oct550   xxxxxx_xxxxxx  625 29-OCT-93  18:02:00    50.      newmixv               No  No  75   1 -23.3  35.0
oct551   931029_191854  104 29-OCT-93  19:19:00    16.     mdidtune     mdidtune  No  No  78   1 -23.3  34.7
oct552   xxxxxx_xxxxxx  141 29-OCT-93  20:19:00    26.      vmagmix               No Yes  76   1 -23.3  33.2
cat080   931029_203324   32 29-OCT-93  20:34:00     3.                   bigtune  No  No  96   3 -20.0  35.5
oct553   xxxxxx_xxxxxx  650 29-OCT-93  20:54:00    52.      newmixv               No  No  78   1 -20.0  35.1
cat081   931029_224151    7 29-OCT-93  22:42:00     1.                  mdidtune  No  No  96   2 -23.3  35.3
oct554   931029_225806  104 29-OCT-93  22:58:00    16.          xxx     mdidtune  No  No  75   1 -20.0  35.2
oct555   xxxxxx_xxxxxx  247 29-OCT-93  23:18:00    52.      newmixv               No  No  77   1 -23.3  35.2
oct561   xxxxxx_xxxxxx   41 30-OCT-93  00:00:00 ?   0.      vmagmix               No Yes  95   2 -23.3  35.0
oct562   xxxxxx_xxxxxx  122 30-OCT-93  00:00:00 ?   0.      vmagmix               No Yes  95   1 -23.3  35.0
oct558   xxxxxx_xxxxxx   89 30-OCT-93  00:00:00 ?   0.    caldopser               No  No  95   1 -23.3  35.0
oct559   xxxxxx_xxxxxx  118 30-OCT-93  00:00:00 ?   0.      ldopser               No  No  95   1 -23.3  35.0
oct556   xxxxxx_xxxxxx  250 30-OCT-93  02:02:00    32.    supertune               No  No  76   1 -20.0  36.1
oct557   xxxxxx_xxxxxx  328 30-OCT-93  02:39:00    25.    supertune               No Yes  78   1 -20.0  34.7
cat082   931030_175712   90 30-OCT-93  17:58:00    18.                  obsmodev  No  No  96   2 -33.3  34.8
oct560   931030_190831  145 30-OCT-93  19:08:31 ?   0.    caldopser      vmagmix Yes Yes  -1  -1    ??  40.7
cat083   931030_221644    1 30-OCT-93  22:16:44 ?   0.                  rundepsq  No  No  96   4 -33.3  34.9
oct564   931030_222634  185 30-OCT-93  22:27:00    35.       dopser        sleep  No  No  95   4 -33.3  34.4
oct570   xxxxxx_xxxxxx  228  1-NOV-93  00:00:00 ?   0.      sundust               No  No  93   1 -33.3  34.8
oct569a  xxxxxx_xxxxxx  200  1-NOV-93  00:00:00 ?   0.      boiloff               No  No  -1  -1 -33.3  34.8
oct565   xxxxxx_xxxxxx  187  1-NOV-93  21:46:00   101.      sundust               No  No  78   1 -36.7  37.1
cat084   931102_144232    4  2-NOV-93  14:43:00    18.                            No  No  96   4    ??    ??
oct566   931102_163611 2300  2-NOV-93  16:37:00   461.          xxx      openlog  No  No  90   1    ??    ??
cat085   931103_005403    3  3-NOV-93  00:56:00     7.                            No  No  96   4    ??    ??
cat086   931103_011224    2  3-NOV-93  01:14:00     2.                            No  No  96   4    ??    ??
cat087   931103_012818    1  3-NOV-93  01:32:00     0.                            No  No  96   4    ??    ??
cat088   931103_013705    6  3-NOV-93  01:38:00    10.                    mdiltc  No  No  96   4    ??    ??
oct567   xxxxxx_xxxxxx   30  3-NOV-93  02:41:00    10.          ltc               No  No  93   1    ??    ??
oct568   xxxxxx_xxxxxx   21  3-NOV-93  03:20:00     2.          ltc               No  No  95   5    ??    ??
oct569   931103_155221 2460  3-NOV-93  15:51:00   491.      dopmixv      openlog  No  No  95   6    ??    ??
cat089   931104_155245    2  4-NOV-93  15:57:00     0.                   obsmode  No  No  96   4    ??    ??
cat090   931104_171445    2  4-NOV-93  17:17:00     0.                            No  No  96   4    ??    ??
cat091   931104_172850    2  4-NOV-93  17:31:00     2.                   errdump  No  No  96   4    ??    ??
cat092   931104_174502    4  4-NOV-93  17:45:00     8.                            No  No  96   4    ??    ??
cat093   931104_181657    2  4-NOV-93  18:20:00     0.                   calmode  No  No  -1  -1    ??    ??
cat094   931104_195727    8  4-NOV-93  19:57:00    14.                  vfdlscan  No  No 107   4    ??    ??
oct571   xxxxxx_xxxxxx   12  4-NOV-93  22:53:00     2.     sunlscan               No  No  93   1    ??    ??
oct572   931104_234942  104  4-NOV-93  23:50:00    16.     dyedtune     mdidtune  No  No  93   1    ??    ??
oct573   931105_002930  104  5-NOV-93  00:30:00    16.     dyedtune     mdidtune  No  No  93   1    ??    ??
oct574   xxxxxx_xxxxxx  100  5-NOV-93  01:09:00     7.    dyestable               No  No  93   1    ??    ??
oct575   931105_022023   24  5-NOV-93  02:20:00     4.          xxx               No  No  93   1    ??    ??
oct576   931105_022943   24  5-NOV-93  02:30:00     3.          xxx               No  No  93   1    ??    ??
oct577   931105_023819   24  5-NOV-93  02:38:00     4.          xxx               No  No  93   1    ??    ??
oct578   931105_024723   24  5-NOV-93  02:47:00     4.          xxx               No  No  93   1    ??    ??
oct579   931105_025443   24  5-NOV-93  02:55:00     3.          xxx               No  No  93   1    ??    ??
oct580   931105_031013   24  5-NOV-93  03:10:00     4.          xxx               No  No  93   1    ??    ??
oct581   931105_031848   24  5-NOV-93  03:19:00     4.          xxx               No  No  93   1    ??    ??
oct582   931105_033054   24  5-NOV-93  03:31:00     4.          xxx               No  No  93   1    ??    ??
oct583   931105_033938   24  5-NOV-93  03:40:00     3.          xxx               No  No  93   1    ??    ??
oct584   931105_040742   24  5-NOV-93  04:08:00     3.          xxx               No  No  93   1    ??    ??
oct585   931105_041540   24  5-NOV-93  04:16:00     3.          xxx               No  No  93   1    ??    ??
oct586   931105_042326   24  5-NOV-93  04:23:00     4.          xxx               No  No  93   1    ??    ??
oct587   931105_045923 1000  5-NOV-93  05:00:00   174.          xxx       wakeup  No  No  93   1    ??    ??
cat095   931105_164442    2  5-NOV-93  16:45:00     2.                   obsmode  No  No 107   4    ??    ??
cat096   931105_170855    2  5-NOV-93  17:09:00     2.                            No  No 107   4    ??    ??
cat097   931105_174036    2  5-NOV-93  17:41:00     0.                   errdump  No  No 115   4    ??    ??
cat098   931105_175852    2  5-NOV-93  17:59:00     0.                            No  No 115   4    ??    ??
cat099   931105_181831    2  5-NOV-93  18:19:00     0.                            No  No 115   4    ??    ??

cat100   931105_182100    2  5-NOV-93  18:21:00     0.                  mdidtune  No  No 115   4    ??    ??
oct588   931105_194016  104  5-NOV-93  19:40:00    16.          xxx     mdidtune  No  No  93   1    ??    ??
oct589   931105_195912   11  5-NOV-93  20:00:00     2.          xxx     vfdlscan  No  No  93   1    ??    ??
oct590   931105_201216  104  5-NOV-93  20:12:00    16.          xxx     mdidtune  No  No  93   1    ??    ??
oct591   931105_203049   11  5-NOV-93  20:32:00     2.          xxx     vfdlscan  No  No  93   1    ??    ??
cat101   931105_205546    2  5-NOV-93  20:56:00     2.                            No  No 107   4    ??    ??
cat102   931105_210122    4  5-NOV-93  21:01:00     6.                  ipdynmac  No  No 107   4    ??    ??
cat103   931105_214744    2  5-NOV-93  21:48:00     2.                   lmchson  No  No 107   4    ??    ??
oct592   931105_232820   29  5-NOV-93  23:28:00     5.          xxx      mtmtune  No  No  93   1    ??    ??
oct593   931105_234456   29  5-NOV-93  23:45:00     5.          xxx      mtmtune  No  No  93   1    ??    ??
oct594   931105_235455   29  5-NOV-93  23:55:00     5.          xxx      mtmtune  No  No  93   1    ??    ??
oct595   931106_000448   29  6-NOV-93  00:05:00     4.          xxx      mtmtune  No  No  93   1    ??    ??
oct596   931106_004551   29  6-NOV-93  00:46:00     4.          xxx      mtmtune  No  No  95   7    ??    ??
oct597   931106_005437   29  6-NOV-93  00:55:00     4.          xxx      mtmtune  No  No  93   1    ??    ??
oct598   931106_013612   29  6-NOV-93  01:36:00     5.          xxx      mtmtune  No  No  94   1    ??    ??
oct599   931106_014445   29  6-NOV-93  01:45:00     6.          xxx      mtmtune  No  No  94   1    ??    ??
oct600   931106_015624    1  6-NOV-93  01:56:00     0.          xxx      mtmtune  No  No  94   1    ??    ??
oct601   931106_020449   29  6-NOV-93  02:05:00     4.          xxx      mtmtune  No  No  94   1    ??    ??
oct602   931106_021515   29  6-NOV-93  02:15:00     6.          xxx      mtmtune  No  No  94   1    ??    ??
oct603   931106_022629   29  6-NOV-93  02:27:00     4.          xxx      mtmtune  No  No  94   1    ??    ??
oct604   931106_023402   29  6-NOV-93  02:34:00     5.          xxx      mtmtune  No  No  94   1    ??    ??
oct605   931106_024713   29  6-NOV-93  02:47:00     5.          xxx      mtmtune  No  No  94   1    ??    ??
oct606   931106_030239   29  6-NOV-93  03:03:00     4.          xxx      mtmtune  No  No  94   1    ??    ??
oct607   931106_033611   29  6-NOV-93  03:36:00     5.          xxx      mtmtune  No  No  94   1    ??    ??
oct608   931106_034618   29  6-NOV-93  03:46:00     5.          xxx      mtmtune  No  No  94   1    ??    ??
oct609   931106_035446   29  6-NOV-93  03:55:00     4.          xxx      mtmtune  No  No  94   1    ??    ??
oct610   931106_041603   29  6-NOV-93  04:16:00     5.          xxx      mtmtune  No  No  94   1    ??    ??
oct611   931106_042520   29  6-NOV-93  04:25:00     5.          xxx      mtmtune  No  No  94   1    ??    ??
oct612   931106_043318   29  6-NOV-93  04:33:00     5.          xxx     vfdlscan  No  No  94   1    ??    ??
oct613   931106_073424  964  6-NOV-93  07:35:00   152.          xxx      bigtune  No  No  94   1    ??    ??
oct614   931106_101607  963  6-NOV-93  10:16:00   150.          xxx      bigtune  No  No  94   1    ??    ??
oct615   xxxxxx_xxxxxx    9  7-NOV-93  00:41:00     1.       absimg               No  No  95   7 -17.1  35.9
oct616   xxxxxx_xxxxxx   81  7-NOV-93  01:03:00    11.        pdist               No  No  95   7 -17.1  34.3
oct617   xxxxxx_xxxxxx  210  7-NOV-93  01:53:00    76.     distgrid               No  No  93   1 -14.3  34.7
cat104   931120_175953    6 20-NOV-93  18:00:00     9.                    lcmcom  No  No 107   4  26.7  25.6
cat105   931120_184514   31 20-NOV-93  18:46:00     6.                    lsqcom  No  No 107   4  26.7  27.3
cat106   931120_202146    5 20-NOV-93  20:22:00     9.                    lcmcom  No  No 107   3  26.7  27.7
cat107   931120_210453   32 20-NOV-93  21:06:00     5.                    lsqcom  No  No 107   3  30.0  29.1
cat108   931121_180055    6 21-NOV-93  18:01:00     9.                    lcmcom  No  No 107   4  20.0  25.6
cat109   931121_192739   30 21-NOV-93  19:29:00     5.                    lsqcom  No  No 107   4  23.3  31.3
cat110   931121_200140   46 21-NOV-93  20:02:00    16.                  mdifocus  No  No 107   3  23.3  29.7
cat111   931121_201948   10 21-NOV-93  20:20:00     2.                    mdioff  No  No 107   3  26.7  29.3
cat112   931122_215549    6 22-NOV-93  21:56:00     9.                    lcmcom  No  No 107   4  48.0  40.2
cat113   931122_230301   34 22-NOV-93  23:04:00     6.                    lsqcom  No  No 107   3  48.0  40.3
cat114   931122_235018   48 22-NOV-93  23:50:00    10.                    mdiltc  No  No 107   4  44.0  40.1
cat115   931123_001958   73 23-NOV-93  00:21:00    13.                  mdifocus  No  No 107   4  44.0  40.1
cat116   931123_004238    7 23-NOV-93  00:43:00    11.                   brbpict  No  No 107   3  48.0  40.1
cat117   931123_055608    6 23-NOV-93  05:57:00     8.                    lcmcom  No  No 107   3  -8.6  37.9
cat118   931123_070030   33 23-NOV-93  07:01:00     6.                    lsqcom  No  No 107   4  -2.9  37.8
cat119   931123_083619    7 23-NOV-93  08:36:00    14.                   brbpict  No  No 107   4  -8.6  35.2
cat120   931123_085214   48 23-NOV-93  08:52:00    13.                    mdiltc  No  No 107   4  -8.6  34.8
cat121   931123_100920   73 23-NOV-93  10:10:00    12.                            No  No 107   3  -8.6  33.5
cat122   931123_102437    4 23-NOV-93  10:25:00     2.                  lltsetup  No  No 107   4  -5.7  33.4
cat123   931123_103215   27 23-NOV-93  10:33:00    22.                    mdicom  No  No 107   4  -8.6  33.2
cat124   931123_214817    6 23-NOV-93  21:49:00     8.                    lcmcom  No  No 107   3 -26.7  25.6
cat125   931123_225411   27 23-NOV-93  22:55:00     5.                    lsqcom  No  No 107   4 -23.3  25.6
cat126   931123_230951    7 23-NOV-93  23:10:00    12.                   brbpict  No  No 107   4 -20.0  25.6
cat127   931123_232756   32 23-NOV-93  23:28:00     6.                    mdiltc  No  No 107   4 -23.3  25.6
cat128   931123_233841   73 23-NOV-93  23:39:00    13.                  brbfocus  No  No 107   4 -23.3  25.6
cat129   931124_044913    6 24-NOV-93  04:50:00     8.                    lcmcom  No  No 107   3 -23.3  25.8

cat130   931124_072950   27 24-NOV-93  07:30:00    24.                            No  No 107   4 -20.0  35.6
cat131   931124_075730   43 24-NOV-93  07:58:00     7.                            No  No 107   3 -20.0  35.6
cat132   931124_081016   73 24-NOV-93  08:11:00    12.                  brbfocus  No  No 107   4 -20.0  35.7
cat133   931124_082454   14 24-NOV-93  08:25:00     2.                    mdiltc  No  No 107   3 -20.0  35.6
cat134   931124_083139   13 24-NOV-93  08:32:00     2.                    mdiltc  No  No 107   4 -20.0  35.6
cat135   931124_083549    1 24-NOV-93  08:36:00     0.                    mdiltc  No  No 107   3 -20.0  35.5
cat136   931124_083856   46 24-NOV-93  08:39:00     8.                    mdiltc  No  No 107   4 -20.0  35.6
cat137   931124_084903    7 24-NOV-93  08:49:00    12.                     ipcom  No  No 107   3 -20.0  35.6
cat138   931124_223018    1 24-NOV-93  22:30:18 ?   0.                  ipdynmac  No  No  -1  -1 -11.4  35.3
cat139   931125_230840    1 25-NOV-93  23:09:00     0.                  obsmodev  No  No 107   3  -5.7  33.2
nov618a  931125_231103    9 25-NOV-93  23:11:00    19.                  fdhrprot  No  No 107   9  -5.7  33.2
nov618b  931126_071012   10 26-NOV-93  07:10:00    19.                    fdprot  No  No 107   9   8.0  33.5
nov618c  931126_074541   10 26-NOV-93  07:46:00    19.                  fdhrprot  No  No 107  10  -8.6  33.2
nov618d  931126_081339   10 26-NOV-93  08:14:00    20.                  fdhrprot  No  No 107   9 -11.4  33.2
nov618e  931126_084410   10 26-NOV-93  08:44:00    18.                  fdhrprot  No  No 107  10 -11.4  33.2
nov618f  931126_091104   10 26-NOV-93  09:11:00    17.                  fdhrprot  No  No 107   9 -14.3  33.2
nov619   931126_095713  396 26-NOV-93  09:58:00   122.                  obsmodev  No  No 107  10 -14.3  33.2
nov620   931126_200554  104 26-NOV-93  20:06:00    18.                  mdidtune  No  No 107   9 -14.3  33.2
nov621   931126_223814   29 26-NOV-93  22:38:00     5.                            No  No 107   9 -11.4  33.2
nov622   931126_225405   29 26-NOV-93  22:54:00     5.                            No  No 107   9 -14.3  33.1
nov623   931126_230423   29 26-NOV-93  23:04:00     6.                   mtmtune  No  No 107   9 -14.3  33.2
nov624   931126_231551   30 26-NOV-93  23:16:00     6.                   mtmtune  No  No 107   9 -14.3  33.2
nov625   931126_232631   29 26-NOV-93  23:27:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov626   931126_233710   29 26-NOV-93  23:37:00     5.                   mtmtune  No  No 107   9 -14.3  33.2
nov627   931126_234525   29 26-NOV-93  23:46:00     5.                   mtmtune  No  No 107  10 -14.3  33.2
nov628   931126_235528   29 26-NOV-93  23:56:00     5.                   mtmtune  No  No 107   9 -14.3  33.2
nov629   931127_000834   29 27-NOV-93  00:09:00     5.                   mtmtune  No  No 107  10 -14.3  33.2
nov630   931127_001658   29 27-NOV-93  00:17:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov631   931127_003129   29 27-NOV-93  00:32:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov632   931127_004016   29 27-NOV-93  00:40:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov633   931127_004909   29 27-NOV-93  00:49:00     5.                   mtmtune  No  No 107   9 -14.3  33.2
nov634   931127_005846   29 27-NOV-93  00:59:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov635   931127_010818   29 27-NOV-93  01:08:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov636   931127_011818   29 27-NOV-93  01:18:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov637   931127_012814    8 27-NOV-93  01:28:00     2.                   mtmtune  No  No 116   5 -14.3  33.2
nov638   931127_031747   29 27-NOV-93  03:18:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov639   931127_032724   29 27-NOV-93  03:27:00     6.                   mtmtune  No  No 107   9 -14.3  33.1
nov640   931127_035144   29 27-NOV-93  03:52:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov641   931127_040155   29 27-NOV-93  04:02:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov642   931127_041431   29 27-NOV-93  04:15:00     5.                            No  No 107  10 -14.3  33.2
nov643   931127_042725   29 27-NOV-93  04:28:00     5.                   mtmtune  No  No 107   9 -14.3  33.1
nov644a  931128_014411   52 28-NOV-93  01:44:00    39.                  obsmodev  No  No 107  10 -14.3  40.1
nov644b  931128_045909   61 28-NOV-93  04:59:00    32.                  obsmodev  No  No 107   9 -17.1  40.1
nov645   931128_062210  100 28-NOV-93  06:22:00    87.                  legscntr  No  No 107  11 -14.3  40.1
nov646   931128_080748  104 28-NOV-93  08:08:00    53.                   leghyst  No  No 107  10 -14.3  40.1
nov647   931129_010728   30 29-NOV-93  01:08:00    14.                  obsmodev  No  No 107   9 -11.4  40.1
nov648   931129_012618   32 29-NOV-93  01:26:00    16.                   isshyst  No  No 107   9 -11.4  40.1
nov649   931129_015525   94 29-NOV-93  01:56:00    47.                   isshyst  No  No 107   9 -11.4  40.1
nov650   931129_024915   27 29-NOV-93  02:52:00    19.                    mdiiss  No Yes 107  11 -14.3  40.1
nov651   931129_032516   22 29-NOV-93  03:26:00    13.                    mdiiss  No Yes 107  10 -14.3  40.1
nov652   931129_041502   27 29-NOV-93  04:15:00    18.                    mdiiss  No Yes 107  11 -14.3  40.1
nov653   931129_044855   27 29-NOV-93  04:49:00    17.                    mdiiss  No Yes 107  10 -14.3  40.1
nov654   931129_051101   27 29-NOV-93  05:12:00    16.                    mdiiss  No Yes 107  11 -14.3  40.1
nov655   931129_061106   27 29-NOV-93  06:11:00    17.                    mdiiss  No  No 116   5 -14.3  40.0
nov656   931129_073924   29 29-NOV-93  07:39:00     6.                            No  No 107  10 -14.3  40.1
nov657   931129_080328   29 29-NOV-93  08:04:00     5.                            No  No 107   9 -14.3  40.1
nov658   931129_081446   29 29-NOV-93  08:15:00     5.                            No  No 107  11 -14.3  40.1
nov659   931129_082350   29 29-NOV-93  08:24:00     5.                            No  No 107  10 -14.3  40.1
nov660   931129_083342   29 29-NOV-93  08:34:00     5.                            No  No 107   9 -14.3  40.1
nov661   931129_084437   29 29-NOV-93  08:45:00     5.                            No  No 107  11 -14.3  40.0

nov662   931129_085417   29 29-NOV-93  08:54:00     5.                            No  No 107  10 -14.3  40.1
nov663   931129_090730   29 29-NOV-93  09:08:00     5.                            No  No 107   9 -14.3  40.1
nov664   931129_092044   29 29-NOV-93  09:21:00     5.                            No  No 107  11 -14.3  40.1
nov665   931129_093056   29 29-NOV-93  09:31:00     5.                            No  No 107  10 -14.3  40.1
nov666   931129_094248   29 29-NOV-93  09:43:00     5.                            No  No 107   9 -14.3  40.1
nov667   931129_095303   29 29-NOV-93  09:53:00     5.                            No  No 107  11 -14.3  40.1
nov668   931129_100542   29 29-NOV-93  10:06:00     5.                            No  No 107  10 -14.3  40.1
nov669   931129_101604   29 29-NOV-93  10:16:00     5.                            No  No 107   9 -14.3  40.1
nov670   931129_103018   29 29-NOV-93  10:30:00     5.                            No  No 107  11 -14.3  40.1
nov671   931129_103954   29 29-NOV-93  10:40:00     5.                            No  No 107  10 -14.3  40.1
nov672   931129_105309   29 29-NOV-93  10:53:00     5.                            No  No 107   9 -14.3  40.1
nov673   931129_110224   29 29-NOV-93  11:02:00     6.                            No  No 107  11 -14.3  40.1
nov674   931129_111340   29 29-NOV-93  11:14:00     5.                            No  No 107  10 -14.3  40.1
nov675   931129_112337   29 29-NOV-93  11:24:00     5.                            No  No 107   9 -14.3  40.1
nov676   931129_113649   29 29-NOV-93  11:37:00     5.                            No  No 107  11 -14.3  40.1
nov677   931129_121424   62 29-NOV-93  12:15:00    28.                            No  No 107  10 -14.3  40.1
nov678   931129_124511  243 29-NOV-93  12:46:00    60.                            No  No 107   9 -14.3  40.1
nov679   931129_140205   37 29-NOV-93  14:03:00     8.                            No  No 107  11 -14.3  40.1
nov680   931129_143330  243 29-NOV-93  14:34:00    53.                   openlog  No  No 107  10 -14.3  40.1
cat140   931129_202205    4 29-NOV-93  20:22:00     5.                            No  No 107   4 -14.3  35.1
cat141   931129_203723    2 29-NOV-93  20:38:00     2.                            No  No 107   4 -17.1  35.2
cat142   931129_212531    1 29-NOV-93  21:30:00     0.                  ipseqtst  No  No 107   3 -14.3  35.1
cat143   931129_213607   11 29-NOV-93  21:41:00    21.                  ipseqtst  No  No 107   3 -14.3  35.2
cat144   931129_225853    2 29-NOV-93  23:01:00     2.                  closelog  No  No 107   4 -14.3  35.2
nov681   931130_000601   57 30-NOV-93  00:06:00    28.                   pztgrid  No  No 107  11 -14.3  35.2
nov682   931130_003605  125 30-NOV-93  00:36:00    63.                   pzthyst  No  No 107  11 -14.3  35.2
nov683a  931130_021014   20 30-NOV-93  02:10:00    38.                            No  No 107  11 -14.3  35.2
nov683b  931130_025307   10 30-NOV-93  02:53:00    21.                  obsmodev  No  No 107   9 -14.3  34.4
nov684   931130_035815   33 30-NOV-93  03:58:00    16.                    pztmax  No  No 107  10 -14.3  35.2
nov685   931130_070421  242 30-NOV-93  07:05:00   130.                   mapgrad  No  No 107  10 -14.3  35.2
nov686   931130_094918  243 30-NOV-93  09:50:00    95.                  vfdlscan  No  No 107   9 -14.3  35.2
nov687   931130_155254  245 30-NOV-93  15:53:00    79.                            No  No 107  11  -8.6  35.2
cat145   931130_173229    6 30-NOV-93  17:33:00     1.                   mtmtune  No  No 107   4 -14.3  35.3
nov688   931130_173658   29 30-NOV-93  17:37:00     5.                   mtmtune  No  No 107  10 -14.3  35.2
nov689   931130_174943    9 30-NOV-93  17:50:00     1.                   mtmtune  No  No 107  10 -14.3  35.2
cat146   931130_181423    8 30-NOV-93  18:14:00     5.                   mtmtune  No  No 107   4 -14.3  35.2
cat147   931130_182423   29 30-NOV-93  18:24:00     6.                            No  No 107   4 -14.3  35.2
nov691   931130_183856   29 30-NOV-93  18:39:00     5.                   mtmtune  No  No 107  10 -14.3  35.2
nov692   931130_185106   29 30-NOV-93  18:51:00     5.                   mtmtune  No  No 107   9 -14.3  35.2
cat148   931130_190908   29 30-NOV-93  19:09:00     5.                   mtmtune  No  No 107   4 -14.3  35.0
nov693   931130_190908   29 30-NOV-93  19:09:00     5.                   mtmtune  No  No 107  10 -14.3  35.0
nov694   931130_192229   29 30-NOV-93  19:23:00     5.                   mtmtune  No  No 107   9 -14.3  35.2
nov695   931130_193403   29 30-NOV-93  19:34:00     5.                   mtmtune  No  No 107  10 -14.3  35.1
nov696   931130_194548   29 30-NOV-93  19:46:00     5.                   mtmtune  No  No 107   9 -14.3  35.2
nov697   931130_195705   29 30-NOV-93  19:57:00     5.                   mtmtune  No  No 107  10 -14.3  35.3
nov698   931130_201256   29 30-NOV-93  20:13:00     5.                   mtmtune  No  No 107   9 -14.3  35.2
nov699   931130_202811   29 30-NOV-93  20:28:00     5.                   mtmtune  No  No 107  10 -14.3  35.2
nov700   931130_203918   29 30-NOV-93  20:39:00     5.                   mtmtune  No  No 107   9 -14.3  35.2
nov701   931130_205003   29 30-NOV-93  20:50:00     5.                   mtmtune  No  No 107  10 -14.3  35.1
nov702   931130_210035   29 30-NOV-93  21:01:00     5.                   mtmtune  No  No 107  11 -14.3  35.3
nov703   931130_211400   29 30-NOV-93  21:14:00     5.                   mtmtune  No  No 107   9 -14.3  35.2
nov704   931130_212832   29 30-NOV-93  21:29:00     5.                   mtmtune  No  No 107  10 -14.3  35.2
nov705   931130_214300   29 30-NOV-93  21:43:00     5.                   mtmtune  No  No 107  11 -14.3  35.2
nov706   931130_220201   29 30-NOV-93  22:02:00     5.                   mtmtune  No  No 107   9 -14.3  35.2
nov707   931130_222455   11 30-NOV-93  22:28:00    19.                  ipseqtst  No  No 107  10 -14.3  35.2
cat149   931201_181306    3  1-DEC-93  18:14:00     0.                  iptbltst Yes  No 115   4  23.3  25.6
nov708   931201_211929   14  1-DEC-93  21:21:00    23.                  iptbltst  No  No 107  11  23.3  25.6
cat150   931201_221344    1  1-DEC-93  22:14:00     0.                  mdidtune  No  No 109   8  26.7  25.6
nov709   931201_221559  104  1-DEC-93  22:16:00    18.                   bigtune  No  No 107   9  23.3  25.6
nov710   931202_190458  103  2-DEC-93  19:05:00    19.                  mdidtune  No  No 107   9  23.3  25.6

cat151   931203_004511   12  3-DEC-93  00:48:00    26.                  ipseqtst  No  No 109   6  26.7  25.6
nov711   931203_192121    4  3-DEC-93  19:25:00     6.                   obsmode  No  No 116   1  26.7  25.6
nov712   931203_201023   42  3-DEC-93  20:12:00     8.                    fddair  No  No 107  13  26.7  25.6
nov713   931204_165529   16  4-DEC-93  16:56:00    43.                  ipmemtst  No  No 116   9  23.3  25.6
cat152   931205_221408    6  5-DEC-93  22:15:00     8.                    lcmcom  No Yes 109   9  23.3  25.6
cat153   931214_004957    6 14-DEC-93  00:50:00     9.                    lcmcom  No Yes 109   9  23.3  25.6
cat154   931215_024536    7 15-DEC-93  02:46:00    12.                   brbpict  No  No 109   8  26.7  25.6
cat155   931215_152905    1 15-DEC-93  15:29:05 ?   0.                  ipvelmac  No  No  -1  -1  26.7  25.6
cat156   931215_163228   11 15-DEC-93  16:38:00    21.                  ipseqtst  No  No 109   8  26.7  25.6
cat157   931217_022821    5 17-DEC-93  02:31:00    17.                    mdioff  No  No 109   9    ??  25.6
cat158   931217_222334    6 17-DEC-93  22:24:00     9.                    lcmcom  No  No 109   5  26.7  25.9
cat159   931217_230607   33 17-DEC-93  23:07:00     6.                     ipcom Yes  No 109   9  26.7  27.9
cat160   931217_233315    1 17-DEC-93  23:33:15 ?   0.                   lfdopen  No  No  -1  -1  26.7  27.0
cat161   931218_010148   48 18-DEC-93  01:02:00    13.                    mdiltc  No Yes 109   5  26.7  26.3
cat162   931218_011846    7 18-DEC-93  01:19:00    12.                   mdipict  No  No 109   5  26.7  26.2
cat163   931218_013527   73 18-DEC-93  01:36:00    13.                  mdifocus  No  No 115   5  30.0  26.2
cat164   931218_203257    6 18-DEC-93  20:33:00     9.                    mdicom  No  No 109   9  23.3  25.7
cat165   931218_211911   30 18-DEC-93  21:21:00     5.                     ipcom  No  No 107   6  26.7  31.8
cat166   931218_214831    7 18-DEC-93  21:49:00    11.                   mdipict  No  No 107   6  26.7  34.2
cat167   931218_220431   48 18-DEC-93  22:05:00    12.                    mdiltc  No Yes 107   7  26.7  35.2
cat168   931218_221904   73 18-DEC-93  22:20:00    15.                  mdifocus  No  No 115   5  26.7  35.2
cat169   931219_174652    6 19-DEC-93  17:47:00     9.                    lcmcom  No  No 107   7  -2.9  37.6
cat170   931219_190432   31 19-DEC-93  19:06:00     6.                     ipcom  No  No 109   9  -2.9  38.1
cat171   931219_192526    6 19-DEC-93  19:26:00    39.                   tvipcom  No  No 109   2  -2.9  38.1
cat172   931220_064308    7 20-DEC-93  06:43:00    12.                   mdipict  No  No 107   5 -17.1  27.3
cat173   931220_065852   48 20-DEC-93  06:59:00    13.                    mdiltc  No Yes 107   5 -17.1  28.5
cat174   931220_071334   73 20-DEC-93  07:14:00    14.                  mdifocus  No  No 107   5 -17.1  29.2
cat175   931220_074608    6 20-DEC-93  07:47:00     8.                    lcmcom  No  No 107   5 -20.0  30.9
cat176   931220_084656    2 20-DEC-93  08:50:00     2.                    lsqcom  No  No 107   5 -20.0  32.6
cat177   931220_090106    4 20-DEC-93  09:05:00    12.                    mdicom  No  No 107   5 -17.1  31.1
cat178   931220_210852    6 20-DEC-93  21:09:00     9.                    lcmcom  No  No 107   5 -11.4  35.2
cat179   931220_215334   36 20-DEC-93  21:55:00     6.                     ipcom  No  No 107   5  -8.6  35.8
cat180   931220_223624   73 20-DEC-93  22:37:00    16.                  mdifocus  No  No 107   5  -8.6  36.4
cat181   931220_231205   14 20-DEC-93  23:12:00    28.                   calmode  No  No 109   6  -8.6  36.7
cat182   931220_235934    5 21-DEC-93  00:00:00     7.                  closelog  No  No 107   5 -14.3  37.1
cat183   931221_075531    6 21-DEC-93  07:56:00     9.                    lcmcom  No  No 107   5 -42.9  35.8
cat184   931221_082209   27 21-DEC-93  08:24:00     5.                     ipcom  No  No 107   5 -45.7  35.7
cat185   931221_084934   14 21-DEC-93  08:50:00    27.                    mdioff  No  No 109   5 -48.6  35.7
cat186   931221_182624    6 21-DEC-93  18:27:00     9.                    lcmcom  No  No 107   5    ??    ??
cat187   931221_191940   30 21-DEC-93  19:21:00     6.                     ipcom  No  No 107   5 -33.3  36.1
cat188   931221_194038   14 21-DEC-93  19:41:00    28.                   mdipict  No  No 109   7 -30.0  36.2
cat189   931221_203349    1 21-DEC-93  20:34:00     0.                     llton  No  No 107   5 -30.0  36.4
cat190   931221_210002    7 21-DEC-93  21:00:00    12.                   mdipict  No  No 107   5 -30.0  36.4
cat191   931222_030252    1 22-DEC-93  03:02:52 ?   0.                    mdicom  No  No  -1  -1    ??  25.6
cat192   931222_032258    6 22-DEC-93  03:23:00     9.                    lcmcom  No  No 107   5 -48.6  35.8
cat193   931222_042245    1 22-DEC-93  04:22:45 ?   0.                    lsqcom  No  No  -1  -1 -51.4  35.7
cat194   931222_052755 1225 22-DEC-93  05:28:00   245.                    mdicom  No  No 107   8 -51.4  31.8
cat195   931222_103549    6 22-DEC-93  10:36:00     9.                    lcmcom  No  No 107   5 -33.3  34.9
cat196   931222_105703    5 22-DEC-93  10:59:00     5.                     ipcom  No  No 115   5 -33.3  35.0
cat197   931222_111407   14 22-DEC-93  11:14:00    25.                  closelog  No  No 109   5    ??    ??
cat198   931222_163837    7 22-DEC-93  16:39:00    11.                   mdipict  No  No 115   5 -54.3  35.5
cat199   931222_173317    6 22-DEC-93  17:34:00     8.                    mdicom  No  No 107   5 -60.0  35.7
cat200   931222_182058    3 22-DEC-93  18:23:00     3.                     ipcom  No  No 107   5 -60.0  35.7
cat201   931222_184012   14 22-DEC-93  18:41:00    26.                   lfdopen  No  No 109   2 -60.0  36.2
cat202   931222_202513   73 22-DEC-93  20:26:00    13.                  mdifocus  No  No 107   5 -65.7  36.4
cat203   931222_204349    2 22-DEC-93  20:45:00     3.                  lcalmode  No  No 115   5 -71.4  35.8
cat204   931222_221349    3 22-DEC-93  22:15:00     0.                  iptbltst Yes  No 109   5 -82.9  35.8
cat205   931222_224109    2 22-DEC-93  22:49:00     2.                  ipseqtst  No  No 109   9    ??  25.6
cat206   931223_004152   27 23-DEC-93  00:42:00    19.                    mdiiss  No  No 109   5 -88.6  35.6
cat207   931223_010434   73 23-DEC-93  01:05:00    14.                  mdifocus  No  No 107   5 -88.6  35.6

cat208   931223_012013    7 23-DEC-93  01:20:00    12.                   mdipict  No  No 109   9 -80.0  35.5
cat209   931223_013528   48 23-DEC-93  01:36:00    12.                    mdiltc  No Yes 107   5 -85.7  35.6
cat210   931223_053833   73 23-DEC-93  05:40:00    14.                  mdifocus  No  No 109   9 -74.3  35.3
cat211   931223_055631    7 23-DEC-93  05:57:00    11.                   mdipict  No  No 107   5 -74.3  35.2
cat212   931223_061445   48 23-DEC-93  06:15:00    13.                    mdiltc  No Yes 109   9 -74.3  35.2
cat213   931223_074514   27 23-DEC-93  07:45:00    18.                    mdiiss  No  No 107   5 -71.4  35.2
cat214   931223_092008    6 23-DEC-93  09:21:00    11.                  closelog  No  No 115   5 -74.3  35.1
cat215   931223_111900    6 23-DEC-93  11:19:00    12.                    mdicom  No  No 107   5 -80.0  35.2
cat216   931223_134655    6 23-DEC-93  13:47:00     9.                    lcmcom  No Yes 107   3 -85.7  35.4
cat217   931223_145437   37 23-DEC-93  14:58:00     6.                     ipcom  No  No 109   5 -85.7  34.4
cat218   931223_152145   14 23-DEC-93  15:22:00    30.                  lobsmode  No  No 109   5 -85.7  35.3
cat219   931223_163917    7 23-DEC-93  16:39:00    12.                   mdipict  No  No 109   5 -85.7  35.4
cat220   931223_165515   48 23-DEC-93  16:55:00    13.                    mdiltc  No Yes 107   3 -85.7  35.4
cat221   931223_170953   73 23-DEC-93  17:11:00    13.                  mdifocus  No  No 109   9 -85.7  35.4
cat222   931223_174402   73 23-DEC-93  17:45:00    14.                  mdifocus  No  No 107   5 -85.7  35.4
cat223   931223_180435  301 23-DEC-93  18:05:00    53.                   mdimrot  No  No 107   3 -85.7  35.4
cat224   931223_190425  122 23-DEC-93  19:05:00    25.                   mdiprot  No  No 109   9 -85.7  35.3
cat225   931223_193352   21 23-DEC-93  19:34:00    38.                    mdiiss  No  No 107   5 -85.7  35.5
cat226   931223_222528    1 23-DEC-93  22:25:28 ?   0.                  mdircsca  No  No  -1  -1 -85.7  35.3
cat227   931223_225040    3 23-DEC-93  22:51:00     0.                            No  No 109   6 -85.7  35.3
cat228   931223_235525    2 23-DEC-93  23:55:00     6.                   liptlm2  No  No 109   8 -85.7  33.1
nov714   931227_195520   10 27-DEC-93  19:56:00    18.                  lltsetup  No  No 107  10 -17.1  35.2
nov715   931227_213353   98 27-DEC-93  21:34:00   194.                  maphrfov  No  No 116   1 -17.1  35.3
nov716   931228_012649   10 28-DEC-93  01:27:00     2.                            No  No 107  13 -14.3  35.3
nov717   931228_015435   10 28-DEC-93  01:55:00     1.                  hrfocser  No  No 107  13 -17.1  35.2
nov718   931228_015834   10 28-DEC-93  01:59:00     1.                  hrfocser  No  No 107  12 -14.3  35.2
cat229   931228_020234   10 28-DEC-93  02:03:00     1.                  hrfocser  No  No 109   9 -14.3  35.3
nov719   931228_032318  202 28-DEC-93  03:23:00   412.                  closelog  No  No 107  13 -14.3  35.3
nov720   931228_163045   19 28-DEC-93  16:31:00    35.                  mdiltcs1  No  No 107  12 -14.3  35.2
nov721   931228_172003   15 28-DEC-93  17:20:00    29.                  mdiltcs2  No  No 107  10 -14.3  35.2
nov723   931228_195938   48 28-DEC-93  20:00:00    95.                  mdiltcs4  No  No 107  11 -14.3  35.2
nov724   931228_221532   48 28-DEC-93  22:16:00    95.                  mdiltcs1  No  No 107   9 -14.3  35.2
nov725   931228_235936   48 29-DEC-93  00:00:00    95.                  mdiltcs2  No  No 107  12 -14.3  35.2
nov726   931229_024235   52 29-DEC-93  02:43:00    99.                            No  No 107  10 -14.3  35.2
nov727a  931229_070623   64 29-DEC-93  07:06:00    39.                  mdilaser Yes  No 107  14 -14.3  35.2
nov727b  931229_075956   65 29-DEC-93  08:00:00    32.                  obsmodev  No  No 107  14 -14.3  35.3
nov728   931229_084005  328 29-DEC-93  08:40:00    56.                  suprtune  No  No 107  14 -14.3  35.2
nov729   931229_095517  206 29-DEC-93  09:55:00    84.                  closelog  No  No 107  11 -14.3  35.2
nov730   931229_112150   36 29-DEC-93  11:22:00    68.                     pztff  No  No 107   9 -14.3  35.2
nov731   931229_123316   36 29-DEC-93  12:33:00    70.                            No  No 107  12 -14.3  35.2
nov732   931229_134517   48 29-DEC-93  13:45:00    96.                  mdiltcs2  No  No 107  14 -14.3  35.2
cat230   931229_225715   11 29-DEC-93  22:58:00    21.                  ipseqtst  No  No 115   2    ??  25.6
cat231   931230_002209    1 30-DEC-93  00:22:09 ?   0.                   iptmtst  No  No  -1  -1  12.0  35.2
cat232   940104_190945    7  4-JAN-94  19:10:00    11.                  fdfocser  No  No 109   5  26.7  25.6
cat233   940104_192544    9  4-JAN-94  19:26:00    16.                  fdfocser  No  No 115   4  26.7  25.6
cat234   940104_194606    9  4-JAN-94  19:46:00    17.                  hrfocser  No  No 115   4  26.7  25.6
cat235   940104_211609    5  4-JAN-94  21:17:00     9.                  hrfocser  No  No 109   8  26.7  25.6
cat236   940104_212753   19  4-JAN-94  21:28:00    11.                  fdfocser  No  No 115   4  26.7  25.6
cat237   940104_215812    9  4-JAN-94  21:58:00     2.                  fdfocser  No  No 109   5  26.7  25.6
cat238   940104_220112   20  4-JAN-94  22:01:00     7.                  fdfocser  No  No 109   5  26.7  25.6
cat239   940104_222402   20  4-JAN-94  22:24:00     6.                  fdfocser  No  No 109   8  26.7  25.6
cat240   940104_234101    3  4-JAN-94  23:41:00     1.                  fdfocser  No  No 109   5  26.7  25.6
cat241   940104_234326   20  4-JAN-94  23:44:00     7.                  fdfocser  No  No 109   8  26.7  25.6
cat242   940104_235604   20  4-JAN-94  23:56:00     6.                  fdfocser  No  No 109   5  26.7  25.6
cat243   940105_000329   21  5-JAN-94  00:04:00     6.                  fdfocser  No  No 115   2  26.7  25.6
cat244   940105_001032   20  5-JAN-94  00:11:00     6.                  fdfocser  No  No 109   8  26.7  25.6
cat245   940105_001755   21  5-JAN-94  00:18:00     9.                  fdfocser  No  No 109   5  26.7  25.6
cat246   940105_015305   20  5-JAN-94  01:53:00     6.                  fdfocser  No  No 115   2  26.7  25.6
cat247   940106_232957   73  6-JAN-94  23:30:00    12.                  mdifocus  No  No 109   8  23.3  25.6
cat248   940110_192724    3 10-JAN-94  19:28:00     6.                   iptmtst  No  No 109   5  26.7  25.6

cat249   940110_195133    3 10-JAN-94  19:53:00     3.                   iptmtst  No  No 109   6  26.7  25.6
cat250   940111_012617  584 11-JAN-94  01:26:00    73.                    mdioff Yes  No  -1  -1  20.0  25.6
cat251   940111_035400    6 11-JAN-94  03:55:00     8.                    lcmcom  No  No 109   5  26.7  25.7
cat252   940111_043557   32 11-JAN-94  04:37:00     6.                    lsqcom  No  No 109   7  26.7  27.5
cat253   940111_151731   18 11-JAN-94  15:18:00    36.                  ipmemtst  No  No 109   5  23.3  25.6
cat254   940111_185730   12 11-JAN-94  18:58:00    23.                  iptbltst  No  No 109   2  26.7  25.6
cat255   940111_202554    5 11-JAN-94  20:29:00     8.                    mdioff Yes  No 115   2  26.7  25.6
cat256   940112_145145    6 12-JAN-94  14:52:00     9.                    lcmcom  No  No 109   5  23.3  25.6
cat257   940112_154832   35 12-JAN-94  15:50:00     6.                     ipcom Yes  No 109   2  26.7  27.7
cat258   940112_161848   21 12-JAN-94  16:19:00    59.                  closelog Yes  No 109   7  26.7  26.7
cat259   940112_224104    7 12-JAN-94  22:41:00    12.                   mdipict  No  No 109   5  26.7  25.6
cat260   940112_230148   48 12-JAN-94  23:02:00    13.                    mdiltc Yes  No 109   9  26.7  25.6
cat261   940112_233031   73 12-JAN-94  23:31:00    14.                  mdifocus  No  No 109   8  26.7  25.6
cat262   940113_011157  245 13-JAN-94  01:12:00    63.                  mdilaser Yes  No 115   2  26.7  25.6
cat263   940113_073052   27 13-JAN-94  07:31:00    23.                    mdiiss  No  No 109   5  23.3  25.6
cat264   940113_172735   27 13-JAN-94  17:28:00    31.                    mdiiss  No  No 109   9  23.3  25.6
cat265   940113_181243   27 13-JAN-94  18:14:00    23.                    mdiiss Yes  No 109   5  26.7  25.6
cat266   940120_115542    7 20-JAN-94  11:56:00    12.                   mdipict  No  No 109   8  16.0  25.6
cat267   940120_121041    7 20-JAN-94  12:11:00    12.                   mdipict Yes  No 109   9  20.0  25.6
cat268   940120_131844   73 20-JAN-94  13:20:00    19.                  mdifocus  No  No 109   6  20.0  25.6
cat269   940120_134308   82 20-JAN-94  13:43:00    28.                   mdiprot Yes  No 109   5  20.0  25.6
cat270   940120_141706  201 20-JAN-94  14:17:00    57.                     ipcom  No  No 109   5  20.0  25.6
cat271   940120_153426   21 20-JAN-94  15:35:00    52.                    mdioff  No  No 109   5  20.0  25.6
cat272   940120_180210   12 20-JAN-94  18:03:00    13.                    mdiiss Yes  No 109   5  20.0  25.6
cat273   940121_093056   27 21-JAN-94  09:32:00    28.                    mdiiss Yes  No 109   5  20.0  25.6
cat274   940121_150107  245 21-JAN-94  15:01:00    65.                           Yes  No 109   6  20.0  25.6
cat275   940121_161142  245 21-JAN-94  16:12:00    55.                  closelog Yes  No 109   5  23.3  25.6
cat276   940215_185257    6 15-FEB-94  18:53:00     9.                    lcmcom Yes  No 109   2  23.3  25.6
cat277   940215_210515    1 15-FEB-94  21:05:15 ?   0.                     ipcom Yes  No  -1  -1  26.7  27.0
cat278   940215_212633   21 15-FEB-94  21:27:00    55.                   ldmccom  No  No 109   9  26.7  26.2
cat279   940218_020247    3 18-FEB-94  02:03:00     4.                       ice Yes  No 109   5 -65.7  25.6
cat280   940218_023104    3 18-FEB-94  02:31:00     4.                   obsmode Yes  No 109   2 -65.7  25.6
cat281   940218_160854    3 18-FEB-94  16:09:00     4.                       ice Yes  No 109   5 -71.4  25.6
cat282   940218_161705    3 18-FEB-94  16:17:00     4.                       ice  No  No 109   2 -68.6  25.6
cat283   940218_180531    3 18-FEB-94  18:06:00     3.                   calmode Yes  No 115   4 -68.6  25.6
cat284   940218_221804    3 18-FEB-94  22:18:00     4.                   obsmode  No  No 109   5 -65.7  25.6
cat285   940219_025607   73 19-FEB-94  02:56:00    13.                  mdifocus  No  No 109   5 -65.7  25.6
cat286   940219_032649    3 19-FEB-94  03:27:00     4.                       ice  No  No 109   2 -65.7  25.6
cat287   940219_155830    3 19-FEB-94  15:59:00     4.                   calmode  No  No 109   5 -71.4  25.6
cat288   940219_160956    3 19-FEB-94  16:10:00     4.                       ice  No  No 115   4 -74.3  25.6
cat289   940219_215106    3 19-FEB-94  21:51:00     4.                   calmode  No  No 109   5 -71.4  25.6
cat290   940219_215846    3 19-FEB-94  21:59:00     4.                       ice  No  No 109   2 -71.4  25.6
cat291   940220_070232    3 20-FEB-94  07:03:00     4.                       ice  No  No 109   5    ??  40.7
cat292   940220_071248   48 20-FEB-94  07:13:00    13.                    mdiltc  No  No 115   4 -71.4  25.6
cat293   940220_153003    3 20-FEB-94  15:30:00     4.                   calmode  No  No 109   5    ??  40.7
cat294   940220_154207    3 20-FEB-94  15:42:00     4.                       ice  No  No 109   5    ??  40.7
cat295   940220_234953    3 20-FEB-94  23:50:00     4.                   calmode  No  No 109   2    ??  40.7
cat296   940220_235817    3 20-FEB-94  23:59:00     3.                       ice Yes  No 109   5    ??  40.7
cat297   940221_070028    3 21-FEB-94  07:01:00     3.                       ice  No  No 109   9    ??  40.7
cat298   940221_160347    3 21-FEB-94  16:04:00     4.                       ice Yes  No 109   6    ??  40.7
cat299   940221_234759    3 21-FEB-94  23:48:00     4.                       ice  No  No 109   5 -65.7  25.6
cat300   940222_150142    3 22-FEB-94  15:02:00     4.                       ice Yes  No 109   5    ??  40.7
cat301   940222_190140    2 22-FEB-94  19:03:00     0.                    lltcal  No  No 115   4    ??  40.7
nov735a  940222_190446   58 22-FEB-94  19:06:00    57.                    lltcal Yes  No 116   2    ??  40.7
cat303   940222_201609   12 22-FEB-94  20:19:00     9.                    lltcal Yes  No 109   2 -65.7  25.6
nov735b  940222_203749   22 22-FEB-94  20:38:00    17.                    lltcal Yes  No 116   9 -65.7  25.6
cat305   940223_012551    3 23-FEB-94  01:26:00     4.                  closelog Yes  No 109   9    ??    ??
cat306   940223_015615    3 23-FEB-94  01:57:00     3.                       ice Yes  No 109   2    ??    ??
cat307   940223_021025    3 23-FEB-94  02:11:00     3.                       ice Yes  No 109   6    ??  40.7
cat308   940223_151923    3 23-FEB-94  15:20:00     3.                       ice Yes  No 109   5    ??  40.7

cat309   940223_201113    3 23-FEB-94  20:11:00     4.                       ice Yes  No 109   5 -62.9  28.5
nov735c  940224_005004   59 24-FEB-94  00:51:00    45.                    lltcal Yes  No 116   7 -62.9  36.7
nov735d  940224_014138   52 24-FEB-94  01:43:00    38.                  closelog Yes  No 116  10 -62.9  35.2
cat312   940224_022501    3 24-FEB-94  02:25:00     4.                    lltoff Yes  No 109   9 -62.9  35.3
cat313   940224_160638    3 24-FEB-94  16:07:00     4.                       ice Yes  No 109   2 -68.6  35.2
cat314   940224_161413    3 24-FEB-94  16:14:00     4.                       ice Yes  No 109   6 -65.7  35.3
cat315   940224_162658   48 24-FEB-94  16:27:00    13.                    mdiltc Yes  No 109   5 -65.7  35.2
nov735e  940224_182543   60 24-FEB-94  18:26:00    46.                  vfdlscan Yes  No 116   8 -65.7  35.3
cat317   940225_022745    3 25-FEB-94  02:28:00     4.                  obsmodev  No  No 109   5 -60.0  35.2
cat318   940225_161235    3 25-FEB-94  16:13:00     3.                       ice  No  No 109   9 -65.7  35.5
cat319   940226_004937    3 26-FEB-94  00:50:00     3.                       ice  No  No 109   2 -60.0  35.2
cat320   940226_150554    3 26-FEB-94  15:06:00     4.                       ice Yes  No 115   4 -65.7  35.3
cat321   940226_151309    3 26-FEB-94  15:13:00     4.                       ice Yes  No 109   6 -65.7  35.2
cat322   940226_151943    3 26-FEB-94  15:20:00     4.                       ice Yes  No 109   5 -65.7  35.2
cat323   940227_034259    3 27-FEB-94  03:43:00     4.                       ice Yes  No 109   5 -65.7  35.2
cat324   940227_143117    3 27-FEB-94  14:31:00     4.                       ice Yes  No 109   5 -65.7  35.2
cat325   940227_143909    3 27-FEB-94  14:39:00     4.                       ice Yes  No 109   9 -65.7  35.1
cat326   940227_145237   48 27-FEB-94  14:53:00    12.                    mdiltc Yes  No 109   2 -65.7  35.2
cat327   940228_055509    3 28-FEB-94  05:55:00     4.                       ice  No  No 115   4 -65.7  35.3
cat328   940228_060250    3 28-FEB-94  06:03:00     4.                       ice  No  No 109   6 -65.7  35.3
cat329   940228_151100    3 28-FEB-94  15:11:00     4.                       ice  No  No 109   5 -62.9  35.0
cat330   940228_212319    2 28-FEB-94  21:24:00     1.                            No  No 109   5   8.0  35.0
cat331   940228_213019    2 28-FEB-94  21:31:00     1.                           Yes  No 109   5   8.0  35.0
cat332   940228_214304    2 28-FEB-94  21:43:00     2.                            No  No 109   9   8.0  35.0
cat333   940228_220408    3 28-FEB-94  22:05:00     0.                  iptbltst Yes  No 115   4  12.0  35.0
cat334   940301_035022    3  1-MAR-94  03:51:00     3.                           Yes  No 109   2 -60.0  35.0
cat335   940301_142036    3  1-MAR-94  14:21:00     4.                            No  No 109   5 -62.9  35.4
cat336   940302_004342    3  2-MAR-94  00:44:00     4.                       ice Yes  No 115   4 -60.0  35.2
cat337   940302_143456    3  2-MAR-94  14:35:00     4.                  ipmemtst Yes  No 109   6 -65.7  35.2
cat338   940302_164144    2  2-MAR-94  16:42:00     2.                            No  No 109   7 -60.0  35.2
cat339   940302_171309    4  2-MAR-94  17:14:00     0.                  nwtbltst Yes  No 109   5 -60.0  35.1
cat340   940303_003621    3  3-MAR-94  00:37:00     3.                  ipmemtst  No  No 109   5 -62.9  34.1
cat341   940303_041552    3  3-MAR-94  04:16:00     1.                     sleep Yes  No 109   7 -60.0  35.1
cat342   940303_145012    3  3-MAR-94  14:50:00     4.                  iptbltst Yes  No 109   7 -62.9  35.3
cat343   940303_160110    9  3-MAR-94  16:05:00     3.                  nwtbltst Yes  No 109   7   0.0  34.3
cat344   940304_210400    4  4-MAR-94  21:05:00     0.                  nwtbltst  No  No 109   2  23.3  33.5
cat345   940304_222132   48  4-MAR-94  22:22:00    13.                    mdiltc  No  No 109   2  23.3  35.1
cat346   940304_223746    7  4-MAR-94  22:38:00    12.                   mdipict Yes  No 109   5  23.3  35.2
cat347   940304_225422    3  4-MAR-94  22:55:00     3.                       ice  No  No 109   6  23.3  35.2
cat348   940304_233629   16  4-MAR-94  23:37:00    11.                  iptbltst  No  No 109   7  23.3  35.0
cat349   940305_004602    4  5-MAR-94  00:47:00     0.                  nwtbltst Yes  No 109   5  26.7  35.0
cat350   940305_005046    5  5-MAR-94  00:51:00     3.                           Yes  No 109   5  26.7  35.0
cat351   940305_005426    5  5-MAR-94  00:54:00     9.                  lobsmode Yes  No 109   5  26.7  35.0
cat352   940305_010802    3  5-MAR-94  01:08:00     4.                       ice  No  No 109   7  26.7  35.2
cat353   940305_021427    5  5-MAR-94  02:15:00     7.                       ice Yes  No 109   5  -5.7  35.2
cat354   940305_023056    5  5-MAR-94  02:31:00     8.                       ice  No  No 109   2 -17.1  34.9
cat355   940305_025054    5  5-MAR-94  02:51:00     8.                       ice Yes  No 109   6 -30.0  35.2
cat356   940305_030844    5  5-MAR-94  03:09:00     7.                       ice  No  No 109   7 -40.0  35.0
cat357   940305_032836    5  5-MAR-94  03:29:00     7.                       ice Yes  No 109   5 -48.6  35.6
cat358   940305_035540    5  5-MAR-94  03:56:00     7.                       ice  No  No 109   5 -57.1  35.2
cat359   940305_044052    5  5-MAR-94  04:41:00     8.                       ice Yes  No 109   5 -62.9  35.0
cat360   940305_080636    5  5-MAR-94  08:07:00     7.                       ice Yes  No 109   2 -62.9  35.3
cat361   940305_081947   73  5-MAR-94  08:21:00    13.                  mdifocus  No  No 109   6 -62.9  35.2
cat362   940305_174506    5  5-MAR-94  17:45:00     8.                  iptbltst Yes  No 115   4 -65.7  35.2
cat363   940305_181445    6  5-MAR-94  18:16:00     4.                  nwtbltst Yes  No 109   7 -65.7  35.2
cat364   940305_182619    3  5-MAR-94  18:27:00     2.                           Yes  No 109   5 -62.9  35.3
cat365   940305_183501    7  5-MAR-94  18:35:00    15.                  lobsmode  No  No 109   9 -65.7  35.3
cat366   940305_191026    3  5-MAR-94  19:11:00     2.                  closelog Yes  No 109   7 -62.9  35.1
cat367   940305_200817    5  5-MAR-94  20:08:00     8.                  closelog Yes  No 109   5 -62.9  35.2
cat368   940305_202112   48  5-MAR-94  20:21:00    14.                    mdiltc Yes  No 109   5 -62.9  35.2

cat369   940306_033557    5  6-MAR-94  03:36:00     8.                       ice Yes  No 109   5 -65.7  35.5
cat370   940306_040224    3  6-MAR-94  04:03:00     3.                            No  No 109   2 -65.7  35.2
cat371   940306_041358    3  6-MAR-94  04:14:00     3.                     crop1 Yes  No 109   7 -65.7  35.2
cat372   940306_042432    3  6-MAR-94  04:25:00     3.                   errdump Yes  No 109   6 -65.7  35.2
cat373   940306_163420    5  6-MAR-94  16:35:00     7.                       ice  No  No 115   4 -65.7  35.2
cat374   940307_040448    5  7-MAR-94  04:05:00     7.                       ice Yes  No 109   2 -65.7  35.1
cat375   940307_162906    1  7-MAR-94  16:29:00     0.                  ldcrpdmp Yes  No 109   7 -62.9  35.3
cat376   940307_164649    3  7-MAR-94  16:47:00     4.                  closelog  No  No 109   7 -62.9  35.3
cat377   940307_170019    5  7-MAR-94  17:00:00     8.                       ice Yes  No 109   6 -62.9  35.2
cat378   940307_205914    1  7-MAR-94  20:59:00     0.                  ldcrpdmp Yes  No 109   5 -62.9  35.2
cat379   940307_211717    2  7-MAR-94  21:18:00     2.                     crop1  No  No 109   5 -62.9  35.1
cat380   940307_214852    2  7-MAR-94  21:49:00     2.                           Yes  No 109   5 -62.9  35.2
cat381   940307_220050    2  7-MAR-94  22:02:00     2.                           Yes  No 109   9 -60.0  35.2
cat382   940307_220949    2  7-MAR-94  22:11:00     2.                   dpstart Yes  No 109   7 -60.0  35.2
cat383   940308_020544   12  8-MAR-94  02:06:00     7.                   lltcal1  No  No 109   5 -60.0  35.0
cat384   940308_021438   26  8-MAR-94  02:15:00    14.                    lltoff Yes  No 109   5 -60.0  35.0
cat385   940308_024939    5  8-MAR-94  02:50:00     7.                       ice Yes  No 109   5 -60.0  35.0
nov736a  940308_153620    5  8-MAR-94  15:37:00     7.                       ice  No  No 116  12 -65.7  35.1
nov736b  940308_174522  104  8-MAR-94  17:46:00    62.                  lltsetup Yes  No 116   6 -62.9  35.1
cat386   940309_020737    3  9-MAR-94  02:08:00    14.                  ldcrpdmp  No  No 109   5 -60.0  35.2
cat387   940309_022837    3  9-MAR-94  02:29:00     2.                  ipmemtst  No  No 115   4 -60.0  35.1
cat388   940309_144135    1  9-MAR-94  14:42:00     0.                            No  No 109   9 -65.7  35.2
cat389   940309_144714    2  9-MAR-94  14:48:00     1.                     crop1  No  No 109   9 -62.9  35.2
cat390   940309_153344    3  9-MAR-94  15:34:00     2.                   lfdopen  No  No 109   7 -65.7  35.3
cat391   940309_211426    1  9-MAR-94  21:15:00     0.                  ldcrpdmp  No  No 109   5 -62.9  35.0
cat392   940309_212153    1  9-MAR-94  21:22:00     0.                            No  No 109   5 -60.0  35.2
cat393   940309_212819    3  9-MAR-94  21:29:00     2.                            No  No 109   5 -62.9  35.2
cat394   940309_214441    2  9-MAR-94  21:45:00     2.                            No  No 109   2 -60.0  35.1
cat395   940309_215120    2  9-MAR-94  21:52:00     1.                     llton  No  No 109   6 -60.0  35.0
cat396   940309_235527    5  9-MAR-94  23:56:00     7.                       ice  No  No 109   5 -60.0  35.0
cat397   940311_002501    5 11-MAR-94  00:25:00     8.                       ice  No  No 115   3 -65.7  34.6
cat398   940311_003808   12 11-MAR-94  00:38:00     3.                   rbihunt  No  No 109   5 -65.7  35.2
cat399   940311_004114   12 11-MAR-94  00:42:00     2.                            No  No 109   9 -65.7  35.2
cat400   940311_004445   12 11-MAR-94  00:45:00     3.                            No  No 109   5 -62.9  35.2
cat401   940311_004824   12 11-MAR-94  00:49:00     3.                            No  No 109   9 -62.9  35.2
cat402   940311_005218   12 11-MAR-94  00:53:00     4.                            No  No 109   5 -62.9  35.3
cat403   940311_005645   12 11-MAR-94  00:57:00     5.                            No  No 109   9 -62.9  35.2
cat404   940311_010216   12 11-MAR-94  01:03:00     7.                            No  No 115   5 -62.9  35.4
cat405   940311_010956   12 11-MAR-94  01:10:00    12.                     sleep  No  No 109   8 -62.9  35.2
nov733a  940311_160912    1 11-MAR-94  16:09:12 ?   0.                   rbihunt Yes  No  -1  -1 -65.7  35.2
nov733b  940311_161057   12 11-MAR-94  16:11:00     2.                            No  No 116   2 -65.7  35.3
nov733c  940311_161309   12 11-MAR-94  16:14:00     1.                            No  No 116   9 -68.6  35.3
nov733d  940311_161529   12 11-MAR-94  16:16:00     2.                            No  No 116  10 -65.7  35.2
nov733e  940311_161804   12 11-MAR-94  16:18:00     3.                           Yes  No 116   2 -65.7  35.2
nov733f  940311_162112   12 11-MAR-94  16:22:00     3.                            No  No 116   9 -65.7  35.3
nov733g  940311_162525   12 11-MAR-94  16:26:00     6.                            No  No 116  10 -62.9  35.1
nov733h  940311_163143   12 11-MAR-94  16:32:00    10.                           Yes  No 116   2 -65.7  35.3
cat406   940311_164524    5 11-MAR-94  16:46:00     7.                       ice Yes  No 115   4 -65.7  35.3
nov734a  940311_194156    1 11-MAR-94  19:41:56 ?   0.                  closelog Yes  No 116   9 -62.9  35.2
nov734b  940311_194336   12 11-MAR-94  19:44:00     2.                            No  No 116   9 -62.9  35.2
nov734c  940311_194543   12 11-MAR-94  19:46:00     2.                            No  No 116   7 -62.9  35.2
nov734d  940311_194757   12 11-MAR-94  19:48:00     2.                            No  No 116  10 -62.9  35.2
nov734e  940311_195027   12 11-MAR-94  19:51:00     2.                            No  No 116   2 -62.9  35.1
nov734f  940311_195330   12 11-MAR-94  19:54:00     3.                            No  No 116   9 -62.9  35.2
nov734g  940311_195737   12 11-MAR-94  19:58:00     6.                           Yes  No 116   7 -62.9  35.2
nov734h  940311_200352   12 11-MAR-94  20:04:00    10.                  iptbltst Yes  No 116  10 -62.9  35.1
cat407   940311_213455    1 11-MAR-94  21:36:00     0.                  ldlmbdmp  No  No 115   2 -62.9  35.0
cat408   940311_220645   11 11-MAR-94  22:07:00    17.                  lobsmode  No  No 115   2 -62.9  34.9
cat409   940311_230237    2 11-MAR-94  23:03:00     0.                            No  No 109   5 -60.0  35.0
cat410   940311_231528    2 11-MAR-94  23:16:00     0.                            No  No 109   5 -62.9  35.2

cat411   940311_232118    5 11-MAR-94  23:21:00     8.                            No  No 109   9 -60.0  34.9
cat412   940311_233916    4 11-MAR-94  23:39:00     2.                  lasdtun2  No  No 115   5 -62.9  35.2
nov737   940312_025326  212 12-MAR-94  02:54:00    48.                  calmlasr Yes  No 116   9 -57.1  35.0
nov738a  940312_035601   13 12-MAR-94  03:56:00     2.                  calmlasr Yes  No 116   2 -57.1  34.9
nov738   940312_041421   13 12-MAR-94  03:56:00     2.                  lyotscan  No  No 116   2 -57.1  34.9
nov739   940312_052339  212 12-MAR-94  05:24:00    58.                  obsmlasr Yes  No 116  10 -57.1  35.0
nov740a  940312_145619    1 12-MAR-94  14:56:19 ?   0.                   rbihunt Yes  No 116  10 -62.9  35.3
nov740b  940312_145804   12 12-MAR-94  14:58:00     2.                           Yes  No 116  10 -62.9  35.3
nov740c  940312_150016   12 12-MAR-94  15:01:00     1.                           Yes  No 116   6 -62.9  35.2
nov740d  940312_150237   12 12-MAR-94  15:03:00     2.                           Yes  No 116   3 -62.9  35.2
nov740e  940312_150512   12 12-MAR-94  15:06:00     2.                            No  No 116   9 -62.9  35.3
nov740f  940312_150819   12 12-MAR-94  15:09:00     3.                           Yes  No 116   8 -62.9  35.2
nov740g  940312_151231   12 12-MAR-94  15:13:00     6.                            No  No 116   4 -65.7  35.2
nov740h  940312_151853   12 12-MAR-94  15:19:00    10.                           Yes  No 116   4 -62.9  35.2
cat413   940312_153119    5 12-MAR-94  15:32:00     7.                       ice  No  No 109   5 -62.9  35.1
cat414   940312_161816    3 12-MAR-94  16:19:00     3.                  iptbltst Yes  No 109   8 -62.9  35.1
cat415   940312_162558    3 12-MAR-94  16:26:00     2.                     limb2 Yes  No 115   5 -62.9  35.2
cat416   940312_165104    2 12-MAR-94  16:51:00     2.                     limb2  No  No 109   6 -62.9  35.2
cat417   940312_165822    7 12-MAR-94  16:59:00     5.                     sleep Yes  No 115   3 -62.9  35.3
nov741a  940313_150536    1 13-MAR-94  15:05:36 ?   0.                   rbihunt Yes  No 116  10 -51.4  35.3
nov741b  940313_150719   12 13-MAR-94  15:08:00     1.                            No  No 116  10 -51.4  35.3
nov741c  940313_150932   12 13-MAR-94  15:10:00     2.                           Yes  No 116   6 -51.4  35.3
nov741d  940313_151152   12 13-MAR-94  15:12:00     2.                           Yes  No 116   3 -54.3  35.2
nov741e  940313_151429   12 13-MAR-94  15:15:00     2.                           Yes  No 116   9 -54.3  35.3
nov741f  940313_151738   12 13-MAR-94  15:18:00     4.                            No  No 116   8 -51.4  35.2
nov741g  940313_152148   12 13-MAR-94  15:22:00     6.                           Yes  No 116   4 -51.4  35.2
nov741h  940313_152808   12 13-MAR-94  15:28:00    11.                            No  No 116   4 -51.4  35.2
cat418   940313_154101    5 13-MAR-94  15:42:00     7.                       ice Yes  No 109   8 -51.4  35.3
cat419   940313_155528    1 13-MAR-94  15:56:00     0.                    mdiltc Yes  No 115   5 -51.4  35.2
nov743a  940314_042336    0 13-MAR-94  15:58:00     0.                  closelog Yes  No   0   0 -51.4  35.1
nov742   940313_155733   48 13-MAR-94  15:58:00    12.                  closelog Yes  No 116   1 -51.4  35.1
nov743b  940314_042520   12 14-MAR-94  04:26:00     1.                           Yes  No 116   2 -36.7  35.1
nov743c  940314_042732   12 14-MAR-94  04:28:00     2.                           Yes  No 116  10 -36.7  35.2
nov743d  940314_042951   12 14-MAR-94  04:30:00     2.                           Yes  No 116   6 -40.0  35.2
nov743e  940314_043226   12 14-MAR-94  04:33:00     2.                           Yes  No 116   1 -36.7  35.2
nov743f  940314_043533   12 14-MAR-94  04:36:00     4.                           Yes  No 116   3 -36.7  35.0
nov743g  940314_043945   12 14-MAR-94  04:40:00     6.                           Yes  No 116   9 -36.7  35.1
nov743h  940314_044606   12 14-MAR-94  04:46:00    11.                           Yes  No 116   8 -40.0  35.2
nov744   940314_050027    5 14-MAR-94  05:01:00     7.                       ice Yes  No 116   3 -40.0  35.3
nov745a  940314_144255    1 14-MAR-94  14:42:55 ?   0.                           Yes  No 116   4 -20.0  35.2
nov745b  940314_144439   12 14-MAR-94  14:45:00     2.                           Yes  No 116   4 -20.0  35.2
nov745c  940314_144651   12 14-MAR-94  14:47:00     2.                            No  No 116   4 -20.0  35.2
nov745d  940314_144910   12 14-MAR-94  14:50:00     2.                           Yes  No 116   2 -20.0  35.2
nov745e  940314_145144   12 14-MAR-94  14:52:00     3.                           Yes  No 116  10 -20.0  35.2
nov745f  940314_145451   12 14-MAR-94  14:55:00     4.                           Yes  No 116   6 -20.0  35.2
nov745g  940314_145903   12 14-MAR-94  14:59:00     6.                            No  No 116   1 -20.0  35.3
nov745h  940314_150524   12 14-MAR-94  15:06:00    10.                            No  No 116   3 -20.0  35.1
cat420   940314_153806    1 14-MAR-94  15:38:00     8.                       ice Yes  No 115   5 -20.0  35.0
nov746   940314_155356   48 14-MAR-94  15:54:00    14.                    mdiltc Yes  No 116   9 -17.1  35.1
cat421   940314_195747    5 14-MAR-94  19:58:00     2.                  lmtm2com Yes  No 109   6 -17.1  35.0
cat422   940314_215403    1 14-MAR-94  21:54:00     0.                     sleep Yes  No 109   7  -5.7  35.0
nov747a  940315_140136    1 15-MAR-94  14:01:36 ?   0.                   rbihunt Yes  No 116   9  -5.7  35.1
nov747b  940315_140320   12 15-MAR-94  14:04:00     1.                            No  No 116   9  -5.7  35.0
nov747c  940315_140531   12 15-MAR-94  14:06:00     2.                            No  No 116   8  -5.7  35.2
nov747d  940315_140751   12 15-MAR-94  14:08:00     2.                            No  No 116   4  -5.7  35.0
nov747e  940315_141027   12 15-MAR-94  14:11:00     2.                            No  No 116   4  -5.7  35.0
nov747f  940315_141334   12 15-MAR-94  14:14:00     4.                            No  No 116   2  -5.7  35.0
nov747g  940315_141747   12 15-MAR-94  14:18:00     6.                            No  No 116  10  -5.7  35.0
nov747h  940315_142407   12 15-MAR-94  14:24:00    11.                     sleep  No  No 116   6  -5.7  35.1
nov748   940315_210948    7 15-MAR-94  21:10:00    12.                   mdipict  No  No 116   8  -2.9  35.0

nov749   940315_212548   48 15-MAR-94  21:26:00    13.                    mdiltc  No  No 116   4  -2.9  35.0
cat423   940315_224633    6 15-MAR-94  22:47:00     6.                            No  No 109   7  -5.7  35.0
cat424   940315_225859    2 15-MAR-94  22:59:00     1.                  lcalmode  No  No 115   4  -2.9  35.0
cat425   940316_024130   20 16-MAR-94  02:42:00     3.                  lasdtune  No  No 115   4    ??    ??
cat426   940316_025102   20 16-MAR-94  02:51:00     3.                  calmlasr  No  No 109   6    ??    ??
nov750a  940316_041320   21 16-MAR-94  04:13:00     4.                  obsmlasr  No  No 116   9 -20.0  35.0
nov750b  940316_041906   21 16-MAR-94  04:19:00     3.                            No  No 116   5 -17.1  35.0
nov750c  940316_042529   21 16-MAR-94  04:26:00     3.                            No  No 116   6 -20.0  34.7
nov750d  940316_043154   21 16-MAR-94  04:32:00     3.                            No  No 116   3 -20.0  35.0
nov750e  940316_043822   21 16-MAR-94  04:38:00     4.                  calmlasr  No  No 116   8 -20.0  35.0
nov751a  940316_045128   22 16-MAR-94  04:52:00     3.                  lyotscan  No  No 116   1 -20.0  35.0
nov751b  940316_045732   22 16-MAR-94  04:58:00     3.                            No  No 116  10 -20.0  34.8
nov751c  940316_050426   22 16-MAR-94  05:05:00     3.                            No  No 116   9 -20.0  35.0
nov751d  940316_051119   22 16-MAR-94  05:12:00     3.                            No  No 116   4 -20.0  35.0
nov751e  940316_051812   22 16-MAR-94  05:18:00     4.                            No  No 116   4 -20.0  35.0
nov755   940317_215816    6 17-MAR-94  21:58:00     2.                      grad Yes  No 116  10    ??  25.6
nov756   940317_220244    6 17-MAR-94  22:03:00     1.                   obsmode Yes  No 116  10    ??  25.6
nov757   940318_214229    6 18-MAR-94  21:43:00     1.                   obsmode  No  No 116  10    ??  25.6
nov758   940318_215143    6 18-MAR-94  21:52:00     1.                      grad Yes  No 116   2    ??  25.6
nov759   940318_215500    6 18-MAR-94  21:55:00     1.                           Yes  No 116  10    ??  25.6
nov764   940323_004101    6 23-MAR-94  00:41:00     1.                   obsmode Yes  No 116   2    ??  25.6
nov765   940323_005048    6 23-MAR-94  00:51:00     1.                   openlog Yes  No 116   9    ??  25.6
nov766   940329_050729  245 29-MAR-94  05:08:00    63.                  closelog Yes  No 116   1    ??  25.6
cat436   940330_181138    5 30-MAR-94  18:12:00     2.                           Yes Yes  -1  -1  26.7  25.6
cat440   940331_011035    1 31-MAR-94  01:10:35 ?   0.                           Yes  No  -1  -1  23.3  25.6
apr800   940401_025941   65  1-APR-94  03:00:00   137.                           Yes Yes  -1  -1  23.3  25.6
apr801   940401_061358   56  1-APR-94  06:14:00   125.                    mdioff Yes Yes  -1  -1  26.7  25.6
apr802   940401_142048   48  1-APR-94  14:21:00    13.                    mdiltc Yes Yes  -1  -1  23.3  25.6
apr803   940401_143902    7  1-APR-94  14:39:00    12.                   mdipict Yes Yes  -1  -1  23.3  25.6
apr804   940401_171701   21  1-APR-94  17:17:00    39.                  mdipdist Yes Yes  -1  -1  26.7  25.6
apr805   940401_180345   61  1-APR-94  18:04:00    12.                   mdiprot Yes Yes  -1  -1  26.7  25.6
apr806   940401_183733   61  1-APR-94  18:38:00    25.                    mdiiss Yes Yes  -1  -1  26.7  25.6
apr807   940401_194654   61  1-APR-94  19:47:00    53.                   mdimrot Yes Yes  -1  -1  26.7  25.6
apr808   940401_230257    2  1-APR-94  23:03:00     3.                   lfdopen Yes Yes  -1  -1  23.3  25.6
apr809   940402_033500  245  2-APR-94  03:35:00    63.                  mdilaser Yes Yes  -1  -1  23.3  25.6
apr810   940402_045113   73  2-APR-94  04:52:00    16.                  mdifocus Yes Yes  -1  -1  23.3  25.6
apr811   940402_191303    6  2-APR-94  19:14:00     8.                    lcmcom Yes Yes  -1  -1  20.0  25.9
apr812   940402_202654   31  2-APR-94  20:28:00     6.                    lsqcom Yes Yes  -1  -1  23.3  28.8
apr814   940402_213421   14  2-APR-94  21:34:00     3.                  mapfdfov Yes Yes  -1  -1  30.0  28.7
apr813   940402_213420    1  2-APR-94  21:34:20 ?   0.                           Yes  No  -1  -1  30.0  28.7
apr815   940402_215656   56  2-APR-94  21:57:00   120.                    mdioff Yes Yes  -1  -1  23.3  29.8
apr816   940404_202008    2  4-APR-94  20:20:00     2.                           Yes Yes  -1  -1  23.3  25.6
apr817   940404_203401    1  4-APR-94  20:34:01 ?   0.                     dpc84 Yes  No  -1  -1  23.3  25.6
apr818   940404_205345    1  4-APR-94  20:53:45 ?   0.                           Yes  No  -1  -1  23.3  25.6
apr819   940404_205622    1  4-APR-94  20:56:22 ?   0.                  lfdclose Yes  No  -1  -1  23.3  25.6
apr820   940404_205857    1  4-APR-94  20:58:57 ?   0.                     dpc82 Yes  No  -1  -1  23.3  25.6
apr821   940404_210847    1  4-APR-94  21:08:47 ?   0.                    mdioff Yes  No  -1  -1  23.3  25.6
apr822   940405_225452    0  5-APR-94  22:54:52     0.                           Yes  No   0   0  23.3  25.6
apr823   940405_225802    0  5-APR-94  22:58:02     0.                    mdioff Yes  No   0   0  20.0  25.6
apr824   940406_003403    0  6-APR-94  00:34:03     0.                   lfdopen Yes  No   0   0  23.3  25.6
apr825   940406_004420    0  6-APR-94  00:44:20     0.                    sw1x1s Yes  No   0   0  23.3  25.6
apr826   940406_010943    0  6-APR-94  01:09:43     0.                  mdifocus Yes  No   0   0  23.3  25.6
apr827   940406_012550    0  6-APR-94  01:25:50     0.                  mdifocus Yes  No   0   0  23.3  25.6
apr828   940406_023226    0  6-APR-94  02:32:26     0.                    lcmcom Yes  No   0   0  23.3  25.6
apr829   940406_031203    0  6-APR-94  03:12:03     0.                    lsqcom Yes  No   0   0  26.7  27.0
apr830   940406_140020    0  6-APR-94  14:00:20     0.                    mdiltc Yes  No   0   0  23.3  25.6
apr831   940406_141503    0  6-APR-94  14:15:03     0.                   mdipict Yes  No   0   0  23.3  25.6
apr832   940406_154133    0  6-APR-94  15:41:33     0.                  imgburst Yes  No   0   0  23.3  25.6
apr833   940406_160422    0  6-APR-94  16:04:22     0.                    set_nk Yes  No   0   0  23.3  25.6
apr834   940406_181137    0  6-APR-94  18:11:37     0.                           Yes  No   0   0  23.3  25.6

apr835   940406_181524    0  6-APR-94  18:15:24     0.                           Yes  No   0   0  23.3  25.6
apr836   940406_184839    0  6-APR-94  18:48:39     0.                           Yes  No   0   0  23.3  25.6
apr837   940406_185418    0  6-APR-94  18:54:18     0.                  mdilaser Yes  No   0   0  23.3  25.6
apr838   940407_000702    0  7-APR-94  00:07:02     0.                  mdifocus Yes  No   0   0  23.3  30.0
apr839   940407_013442    0  7-APR-94  01:34:42     0.                   ldsosld Yes  No   0   0    ??  25.6
apr840   940407_014019    0  7-APR-94  01:40:19     0.                           Yes  No   0   0    ??  26.8
apr841   940407_061445    0  7-APR-94  06:14:45     0.                  tsthrfov Yes  No   0   0  26.7  35.2
apr850   940407_151311   73  7-APR-94  15:14:00    13.                  mdifocus Yes Yes  -1  -1 -11.4  35.0
apr852   940407_154635   28  7-APR-94  15:47:00    49.                           Yes  No  -1  -1 -11.4  35.2
apr853   940407_170218    0  7-APR-94  17:02:18     0.     bigpdist     bigpdist Yes  No   0   0 -14.3  35.0
apr854   940407_184651    0  7-APR-94  18:46:51     0. bigpdist: di     bigpdist Yes  No   0   0 -14.3  35.1
apr855   940408_012439    0  8-APR-94  01:24:39     0. mapgrad: Ray      mapgrad Yes  No   0   0 -57.1  35.2
apr856   940408_021453    0  8-APR-94  02:14:53     0. mapgrad: rad      mapgrad Yes  No   0   0 -60.0  35.1
apr857   940408_063328    0  8-APR-94  06:33:28     0.     mdiltcs1     mdiltcs1 Yes  No   0   0 -57.1  35.2
apr858   940408_072850    0  8-APR-94  07:28:50     0.     mdiltcs1     mdiltcs1 Yes  No   0   0 -60.0  35.2
apr859   940408_083541    0  8-APR-94  08:35:41     0.     mdiltcs2     mdiltcs2 Yes  No   0   0 -60.0  35.2
apr860   940408_094207    0  8-APR-94  09:42:07     0.     mdiltcs4     mdiltcs4 Yes  No   0   0 -60.0  35.2
apr861   940408_105052    0  8-APR-94  10:50:52     0. legff (marke        legff Yes  No   0   0 -62.9  35.2
apr862   940408_112834    0  8-APR-94  11:28:34     0.        legff        legff Yes  No   0   0 -60.0  35.1
apr863   940408_125346    0  8-APR-94  12:53:46     0. pztff, cal m        pztff Yes  No   0   0 -60.0  35.3
apr864   940408_134413    0  8-APR-94  13:44:13     0.   pztff, cal               No  No   0   0 -60.0  35.3
apr865   940408_162705    0  8-APR-94  16:27:05     0.     maphrfov     maphrfov Yes  No   0   0 -60.0  35.2
apr866   940408_183154    0  8-APR-94  18:31:54     0.     mapfdfov     mapfdfov Yes  No   0   0 -60.0  35.2
apr867   940408_204844    0  8-APR-94  20:48:44     0. mtmtune: tes      mtmtune Yes  No   0   0 -60.0  35.0
apr868   940408_221234    0  8-APR-94  22:12:34     0. sunclasr: so     vfdlscan Yes  No   0   0 -60.0  35.2
apr869   940408_222903    0  8-APR-94  22:29:03     0. lyotscan, li     lyotscan Yes  No   0   0 -60.0  35.2
apr870   940408_223506    0  8-APR-94  22:35:06     0. lyotscan, li              Yes  No   0   0 -60.0  35.2
apr871   940408_224150    0  8-APR-94  22:41:50     0. lyotscan, li              Yes  No   0   0 -60.0  35.2
apr872   940408_224831    0  8-APR-94  22:48:31     0. lyotscan, li              Yes  No   0   0 -60.0  35.2
apr873   940408_225515    0  8-APR-94  22:55:15     0. lyotscan, li              Yes  No   0   0 -60.0  35.3
apr874   940409_000359    0  9-APR-94  00:03:59     0. lyotscan, bl     lyotscan Yes  No   0   0 -45.7  35.3
apr875   940409_001002    0  9-APR-94  00:10:02     0. lyotscan, bl              Yes  No   0   0 -45.7  35.1
apr876   940409_001645    0  9-APR-94  00:16:45     0. lyotscan, bl              Yes  No   0   0 -42.9  35.3
apr877   940409_002323    0  9-APR-94  00:23:23     0. lyotscan, bl              Yes  No   0   0 -42.9  35.2
apr878   940409_003005    0  9-APR-94  00:30:05     0. lyotscan, bl     calmlasr Yes  No   0   0 -42.9  35.3
apr879   940409_023735    0  9-APR-94  02:37:35     0. sunclasr, di     vfdlscan Yes  No   0   0 -57.1  35.2
apr880   940409_025529    0  9-APR-94  02:55:29     0. lyotscan, li     lyotscan Yes  No   0   0 -57.1  35.2
apr881   940409_030130    0  9-APR-94  03:01:30     0. lyotscn, pos              Yes  No   0   0 -57.1  35.0
apr882   940409_030808    0  9-APR-94  03:08:08     0. lyotscan, po              Yes  No   0   0 -60.0  35.2
apr883   940409_031446    0  9-APR-94  03:14:46     0. lyotscan, po              Yes  No   0   0 -60.0  35.1
apr884   940409_032125    0  9-APR-94  03:21:25     0. lyotscan, po       fdprot Yes  No   0   0 -60.0  35.1
apr885   940409_040550    0  9-APR-94  04:05:50     0. dark, fdprot     calmlasr Yes  No   0   0 -60.0  35.2
apr886   940409_041006    0  9-APR-94  04:10:06     0. lyotscan, la     lyotscan Yes  No   0   0 -60.0  34.9
apr887   940409_041609    0  9-APR-94  04:16:09     0. lyotscan, re              Yes  No   0   0 -60.0  35.3
apr888   940409_042257    0  9-APR-94  04:22:57     0. lyotscan, re              Yes  No   0   0 -60.0  35.3
apr889   940409_042936    0  9-APR-94  04:29:36     0. lyotscan, re              Yes  No   0   0 -60.0  35.2
apr890   940409_043619    0  9-APR-94  04:36:19     0. lyotscan, th     calmlasr Yes  No   0   0 -60.0  35.2
apr896   940409_050245    0  9-APR-94  05:02:45     0. dark, fdprot     calmlasr Yes  No   0   0 -60.0  35.3
apr891   940409_050707    0  9-APR-94  05:07:07     0. lyotscan, re     lyotscan Yes  No   0   0 -62.9  35.2
apr892   940409_051308    0  9-APR-94  05:13:08     0. lyotscan, re              Yes  No   0   0 -60.0  35.3
apr893   940409_051951    0  9-APR-94  05:19:51     0. lyotscan, re              Yes  No   0   0 -60.0  35.2
apr894   940409_052634    0  9-APR-94  05:26:34     0. lyotscan, re              Yes  No   0   0 -60.0  35.0
apr895   940409_053314    0  9-APR-94  05:33:14     0. lyotscan, re     calmlasr Yes  No   0   0 -60.0  35.2
apr897   940409_055703    0  9-APR-94  05:57:03     0. dark, fdprot     calmlasr Yes  No   0   0 -60.0  35.2
apr898   940409_060135    0  9-APR-94  06:01:35     0. lyotscan, re     lyotscan Yes  No   0   0 -60.0  35.3
apr899   940409_060736    0  9-APR-94  06:07:36     0. lyotscan, re              Yes  No   0   0 -60.0  35.1
apr900   940409_061421    0  9-APR-94  06:14:21     0. lyotscan, re              Yes  No   0   0 -60.0  35.3
apr901   940409_062109    0  9-APR-94  06:21:09     0. lyotscan, re              Yes  No   0   0 -60.0  35.2
apr902   940409_062746    0  9-APR-94  06:27:46     0. lyotscan, re        sleep Yes  No   0   0 -60.0  35.2
apr903   940409_162955    0  9-APR-94  16:29:55     0. lsltune, lyo      lsltune Yes  No   0   0 -65.7  35.3

apr904   940409_164227    0  9-APR-94  16:42:27     0. lsltune, red      lsltune Yes  No   0   0 -62.9  35.4
apr905   940409_165719    0  9-APR-94  16:57:19     0. lsltune, red      lsltune Yes  No   0   0 -62.9  35.3
apr906   940409_171247    0  9-APR-94  17:12:47     0. lsltune, red      lsltune Yes  No   0   0 -62.9  35.3
apr907   940409_180731    0  9-APR-94  18:07:31     0. lsltune, blu      lsltune Yes  No   0   0 -62.9  35.2
apr908   940409_182056    0  9-APR-94  18:20:56     0. lsltune, blu      lsltune Yes  No   0   0 -62.9  35.2
apr909   940409_183439    0  9-APR-94  18:34:39     0. lsltune, blu      lsltune Yes  No   0   0 -62.9  35.3
apr910   940409_215413    0  9-APR-94  21:54:13     0. suncals: Dop     vfdlscan Yes  No   0   0 -60.0  35.3
apr911   940409_221733    0  9-APR-94  22:17:33     0. bigtune: old      bigtune Yes  No   0   0 -60.0  35.1
apr912   940409_223834    0  9-APR-94  22:38:34     0. sundrift: dr     sundrift Yes  No   0   0 -57.1  35.2
apr913   940409_233130    0  9-APR-94  23:31:30     0. sunmagc: 102     obsmsunv Yes  No   0   0 -60.0  35.2
apr914   940410_005300    0 10-APR-94  00:53:00     0. sundust: dus     obsmsunv Yes  No   0   0 -60.0  35.1
apr914a  940410_025745    0 10-APR-94  02:57:45     0.      unknown              Yes  No   0   0 -60.0  35.0
apr915   940410_140323    0 10-APR-94  14:03:23     0.  mdipict_ice      mdipict Yes  No   0   0 -65.7  35.2
apr916   940410_142922    0 10-APR-94  14:29:22     0.      suncals     vfdlscan Yes  No   0   0 -68.6  35.1
apr917   940410_145207    0 10-APR-94  14:52:07     0.      suncals     vfdlscan Yes  No   0   0 -65.7  35.2
apr918   940410_153430    0 10-APR-94  15:34:30     0. suncals: Dop     vfdlscan Yes  No   0   0 -65.7  35.2
apr919   940410_161313    0 10-APR-94  16:13:13     0. sunff: conti        dpc84 Yes  No   0   0 -65.7  35.3
apr920   940410_164347    0 10-APR-94  16:43:47     0. sunff: conti        sunff Yes  No   0   0 -62.9  35.2
apr921   940410_172906    0 10-APR-94  17:29:06     0. suncals:mult     vfdlscan Yes  No   0   0 -62.9  35.1
apr922   940410_180747    0 10-APR-94  18:07:47     0.  sunmagc: FD              Yes  No   0   0 -62.9  35.5
apr923   940410_182213    0 10-APR-94  18:22:13     0.  sunmagc: HR              Yes  No   0   0 -62.9  35.3
apr924   940410_183254    0 10-APR-94  18:32:54     0. sunmagc: cal              Yes  No   0   0 -62.9  35.3
apr925   940410_184448    0 10-APR-94  18:44:48     0. sunmagc: big              Yes  No   0   0 -62.9  35.2
apr926   940410_185600    0 10-APR-94  18:56:00     0. sunmagc: big              Yes  No   0   0 -62.9  35.2
apr927   940410_190643    0 10-APR-94  19:06:43     0. sunmagc: vfd              Yes  No   0   0 -62.9  35.2
apr928   940410_194930    0 10-APR-94  19:49:30     0.      suncals     vfdlscan Yes  No   0   0 -62.9  35.1
apr929   940410_203946    0 10-APR-94  20:39:46     0.      sunmisc      sunmisc Yes  No   0   0 -62.9  35.0
apr930   940410_204949    0 10-APR-94  20:49:49     0.      suncals     vfdlscan Yes  No   0   0 -62.9  35.1
apr931   940410_212750    0 10-APR-94  21:27:50     0.     bigfdmag              Yes  No   0   0 -62.9  35.0
apr932   940410_213926    0 10-APR-94  21:39:26     0.     bighrmag              Yes  No   0   0 -62.9  34.9
apr933   940410_215014    0 10-APR-94  21:50:14     0. vfdmag+vhrma              Yes  No   0   0 -62.9  35.0
apr934   940410_215806    0 10-APR-94  21:58:06     0. sunff, 1024^sunff Yes  No   0   0 -62.9  35.0
apr935   940410_231043    0 10-APR-94  23:10:43     0.      suncals     vfdlscan Yes  No   0   0 -62.9  35.3
apr936   940410_234821    0 10-APR-94  23:48:21     0.      fringes      fringes Yes  No   0   0 -62.9  35.2
apr937   940411_010818    0 11-APR-94  01:08:18     0.      suncals     vfdlscan Yes  No   0   0 -62.9  35.3
apr938   940411_014643    0 11-APR-94  01:46:43     0. suncals, abo      suncals Yes  No   0   0 -62.9  35.2
apr939   940411_021952    0 11-APR-94  02:19:52     0. ccdheat: noi     closelog Yes  No   0   0 -62.9  35.2
apr939a  940411_142101    0 11-APR-94  14:21:01     0.      unknown     rundepsq Yes  No   0   0 -65.7  35.3
apr939b  940411_143218    0 11-APR-94  14:32:18     0.      unknown     rundepsq Yes  No   0   0 -65.7  35.3
apr940   940411_150425    0 11-APR-94  15:04:25     0.         ldop              Yes  No   0   0 -65.7  35.2
apr941   940411_162620    0 11-APR-94  16:26:20     0.      ldop941        llton Yes  No   0   0 -62.9  35.3
apr941a  940411_172420    0 11-APR-94  17:24:20     0.      unknown              Yes  No   0   0 -65.7  35.2
apr942   940411_175446    0 11-APR-94  17:54:46     0. sunmagc cont              Yes  No   0   0 -62.9  35.3
apr942a  940411_180531    0 11-APR-94  18:05:31     0.      unknown              Yes  No   0   0 -62.9  35.4
apr943   940411_182654    0 11-APR-94  18:26:54     0. sunmagc bett        dpc84 Yes  No   0   0 -62.9  35.3
apr944   940411_184856    0 11-APR-94  18:48:56     0. sunmagc FD w              Yes  No   0   0 -62.9  35.2
apr945   940411_190129    0 11-APR-94  19:01:29     0.   sunmagc HR              Yes  No   0   0 -62.9  35.2
apr945a  940411_191213    0 11-APR-94  19:12:13     0. sunmagc 4x4       sunmagc Yes  No   0   0 -62.9  35.0
apr945b  940411_193313    0 11-APR-94  19:33:13     0.      unknown        dpc84 Yes  No   0   0 -62.9  35.1
apr946   940411_201246    0 11-APR-94  20:12:46     0. sunmagc FD h              Yes  No   0   0 -62.9  35.3
apr947   940411_203008    0 11-APR-94  20:30:08     0. sunmagc HR h              Yes  No   0   0 -62.9  35.6
apr948   940411_204121    0 11-APR-94  20:41:21     0. sunmagc 4x4         ipcom Yes  No   0   0 -62.9  35.2
apr949   940411_223537    0 11-APR-94  22:35:37     0. ip commissio     ipdynmac Yes  No   0   0 -62.9  35.1
apr950   940411_234155    0 11-APR-94  23:41:55     0. sundust: ape      sundust Yes  No   0   0 -60.0  35.0
apr951   940412_012151    0 12-APR-94  01:21:51     0. rundepsq, in     obsmsunv Yes  No   0   0 -60.0  35.2
apr951a  940412_132049    0 12-APR-94  13:20:49     0.      unknown      0010_05 Yes  No   0   0 -65.7  35.3
apr951b  940412_144232    0 12-APR-94  14:42:32     0.      unknown     rundepsq Yes  No   0   0 -65.7  35.1
apr952   940412_152042    0 12-APR-94  15:20:42     0. solar dopple     calmodev Yes  No   0   0 -65.7  35.2
apr953   940412_155609    0 12-APR-94  15:56:09     0. calmode dopp               No  No   0   0 -65.7  35.1
apr953a  940412_155823    0 12-APR-94  15:58:23     0.      unknown     calmodev Yes  No   0   0 -62.9  35.3

apr953b  940412_163913    0 12-APR-94  16:39:13     0.      unknown     calmodev Yes  No   0   0 -62.9  35.4
apr954   940412_170906    0 12-APR-94  17:09:06     0.     lasdtune     mdilaser Yes  No   0   0 -62.9  35.2
apr955   940412_172802    0 12-APR-94  17:28:02     0. mdilaser, ra              Yes  No   0   0 -62.9  35.2
apr956   940412_183825    0 12-APR-94  18:38:25     0. mdilaser, ra     mdilaser Yes  No   0   0 -62.9  35.3
apr957   940412_193641    0 12-APR-94  19:36:41     0. lyotscan, be     lyotscan  No  No   0   0 -65.7  35.1
apr958   940412_194244    0 12-APR-94  19:42:44     0.     lyotscan               No  No   0   0 -65.7  35.1
apr959   940412_194924    0 12-APR-94  19:49:24     0.     lyotscan               No  No   0   0 -65.7  35.2
apr960   940412_195604    0 12-APR-94  19:56:04     0.     lyotscan               No  No   0   0 -65.7  35.2
apr961   940412_200245    0 12-APR-94  20:02:45     0.     lyotscan               No  No   0   0 -65.7  35.3
apr962   940412_201045    0 12-APR-94  20:10:45     0.     lyotscan               No  No   0   0 -65.7  35.3
apr963   940412_201647    0 12-APR-94  20:16:47     0.     lyotscan               No  No   0   0 -65.7  35.2
apr964   940412_202327    0 12-APR-94  20:23:27     0.     lyotscan               No  No   0   0 -68.6  35.2
apr965   940412_203007    0 12-APR-94  20:30:07     0.     lyotscan               No  No   0   0 -68.6  35.0
apr966   940412_203646    0 12-APR-94  20:36:46     0.     lyotscan               No  No   0   0 -65.7  34.6
apr967   940412_204344    0 12-APR-94  20:43:44     0.     lyotscan               No  No   0   0 -65.7  34.1
apr968   940412_204945    0 12-APR-94  20:49:45     0.     lyotscan              Yes  No   0   0 -65.7  33.9
apr969   940412_205625    0 12-APR-94  20:56:25     0.     lyotscan               No  No   0   0 -65.7  33.5
apr970   940412_210307    0 12-APR-94  21:03:07     0.     lyotscan              Yes  No   0   0 -68.6  33.2
apr971   940412_210949    0 12-APR-94  21:09:49     0.     lyotscan              Yes  No   0   0 -65.7  32.9
apr972   940412_211658    0 12-APR-94  21:16:58     0.     lyotscan               No  No   0   0 -65.7  32.5
apr973   940412_212300    0 12-APR-94  21:23:00     0.     lyotscan              Yes  No   0   0 -65.7  32.4
apr974   940412_212940    0 12-APR-94  21:29:40     0.     lyotscan              Yes  No   0   0 -65.7  32.2
apr975   940412_213620    0 12-APR-94  21:36:20     0.     lyotscan               No  No   0   0 -65.7  32.0
apr976   940412_214301    0 12-APR-94  21:43:01     0.     lyotscan              Yes  No   0   0 -65.7  31.7
apr977   940412_215011    0 12-APR-94  21:50:11     0.     lyotscan              Yes  No   0   0 -65.7  31.5
apr978   940412_215613    0 12-APR-94  21:56:13     0.     lyotscan               No  No   0   0 -65.7  31.3
apr979   940412_220254    0 12-APR-94  22:02:54     0.     lyotscan              Yes  No   0   0 -65.7  31.3
apr980   940412_220935    0 12-APR-94  22:09:35     0.     lyotscan              Yes  No   0   0 -65.7  31.0
apr981   940412_221614    0 12-APR-94  22:16:14     0.     lyotscan               No  No   0   0 -65.7  30.8
apr982   940412_222325    0 12-APR-94  22:23:25     0.     lyotscan              Yes  No   0   0 -65.7  30.6
apr983   940412_222926    0 12-APR-94  22:29:26     0.     lyotscan              Yes  No   0   0 -65.7  30.5
apr984   940412_223605    0 12-APR-94  22:36:05     0.     lyotscan               No  No   0   0 -65.7  30.3
apr985   940412_224244    0 12-APR-94  22:42:44     0.     lyotscan              Yes  No   0   0 -65.7  30.1
apr986   940412_224923    0 12-APR-94  22:49:23     0.     lyotscan              Yes  No   0   0 -65.7  30.1
apr987   940412_225623    0 12-APR-94  22:56:23     0.     lyotscan               No  No   0   0 -65.7  29.7
apr988   940412_230225    0 12-APR-94  23:02:25     0.     lyotscan              Yes  No   0   0 -65.7  29.5
apr989   940412_230905    0 12-APR-94  23:09:05     0.     lyotscan              Yes  No   0   0 -65.7  29.5
apr990   940412_231548    0 12-APR-94  23:15:48     0.     lyotscan               No  No   0   0 -65.7  29.3
apr991   940412_232229    0 12-APR-94  23:22:29     0.     lyotscan              Yes  No   0   0 -65.7  29.1
apr992   940412_232937    0 12-APR-94  23:29:37     0.     lyotscan              Yes  No   0   0 -62.9  29.1
apr993   940412_233932    0 12-APR-94  23:39:32     0.     lyotscan               No  No   0   0 -65.7  28.8
apr994   940412_234613    0 12-APR-94  23:46:13     0.     lyotscan              Yes  No   0   0 -62.9  28.7
apr995   940412_235252    0 12-APR-94  23:52:52     0.     lyotscan               No  No   0   0 -62.9  28.6
apr996   940412_235932    0 12-APR-94  23:59:32     0.     lyotscan              Yes  No   0   0 -62.9  28.4
apr997   940413_000635    0 13-APR-94  00:06:35     0.     lyotscan              Yes  No   0   0 -62.9  28.3
apr998   940413_001237    0 13-APR-94  00:12:37     0.     lyotscan               No  No   0   0 -62.9  28.1
apr999   940413_001917    0 13-APR-94  00:19:17     0.     lyotscan              Yes  No   0   0 -65.7  28.1
apr1000  940413_002556    0 13-APR-94  00:25:56     0.     lyotscan              Yes  No   0   0 -65.7  27.9
apr1001  940413_003237    0 13-APR-94  00:32:37     0.     lyotscan               No  No   0   0 -65.7  27.8
apr1002  940413_003948    0 13-APR-94  00:39:48     0.     lyotscan              Yes  No   0   0 -65.7  27.7
apr1003  940413_004552    0 13-APR-94  00:45:52     0.     lyotscan              Yes  No   0   0 -62.9  27.4
apr1004  940413_005234    0 13-APR-94  00:52:34     0.     lyotscan               No  No   0   0 -62.9  27.5
apr1005  940413_005914    0 13-APR-94  00:59:14     0.     lyotscan              Yes  No   0   0 -65.7  27.4
apr1006  940413_010556    0 13-APR-94  01:05:56     0.     lyotscan              Yes  No   0   0 -62.9  27.2
apr1007  940413_011801    0 13-APR-94  01:18:01     0.     lyotscan               No  No   0   0 -65.7  27.0
apr1008  940413_012403    0 13-APR-94  01:24:03     0.     lyotscan              Yes  No   0   0 -62.9  26.9
apr1009  940413_013044    0 13-APR-94  01:30:44     0.     lyotscan               No  No   0   0 -62.9  26.8
apr1010  940413_013726    0 13-APR-94  01:37:26     0.     lyotscan              Yes  No   0   0 -62.9  26.8
apr1011  940413_014409    0 13-APR-94  01:44:09     0.     lyotscan               No  No   0   0 -65.7  26.6
apr1012  940413_015115    0 13-APR-94  01:51:15     0.     lyotscan              Yes  No   0   0 -65.7  26.6

apr1013  940413_015720    0 13-APR-94  01:57:20     0.     lyotscan              Yes  No   0   0 -65.7  26.4
apr1014  940413_020403    0 13-APR-94  02:04:03     0.     lyotscan               No  No   0   0 -62.9  26.2
apr1015  940413_021045    0 13-APR-94  02:10:45     0.     lyotscan              Yes  No   0   0 -62.9  26.3
apr1016  940413_021726    0 13-APR-94  02:17:26     0.     lyotscan              Yes  No   0   0 -62.9  26.2
apr1017  940413_022428    0 13-APR-94  02:24:28     0.     lyotscan               No  No   0   0 -62.9  26.0
apr1018  940413_023125    0 13-APR-94  02:31:25     0.     lyotscan              Yes  No   0   0 -62.9  25.8
apr1019  940413_023804    0 13-APR-94  02:38:04     0.     lyotscan              Yes  No   0   0 -62.9  25.8
apr1020  940413_024741    0 13-APR-94  02:47:41     0.     lyotscan              Yes  No   0   0 -65.7  25.7
apr1021  940413_025421    0 13-APR-94  02:54:21     0.     lyotscan               No  No   0   0 -62.9  25.6
apr1022  940413_030126    0 13-APR-94  03:01:26     0.     lyotscan              Yes  No   0   0 -62.9  25.6
apr1023  940413_030728    0 13-APR-94  03:07:28     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1024  940413_031409    0 13-APR-94  03:14:09     0.     lyotscan              Yes  No   0   0 -62.9  25.6
apr1025  940413_032049    0 13-APR-94  03:20:49     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1026  940413_032731    0 13-APR-94  03:27:31     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1027  940413_033440    0 13-APR-94  03:34:40     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1028  940413_034042    0 13-APR-94  03:40:42     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1029  940413_034722    0 13-APR-94  03:47:22     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1030  940413_035403    0 13-APR-94  03:54:03     0.     lyotscan              Yes  No   0   0 -62.9  25.6
apr1031  940413_040042    0 13-APR-94  04:00:42     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1032  940413_040754    0 13-APR-94  04:07:54     0.     lyotscan               No  No   0   0 -62.9  25.6
apr1033  940413_041356    0 13-APR-94  04:13:56     0.     lyotscan              Yes  No   0   0 -62.9  25.6
apr1034  940413_042037    0 13-APR-94  04:20:37     0.     lyotscan              Yes  No   0   0 -62.9  25.6
apr1035  940413_042715    0 13-APR-94  04:27:15     0.     lyotscan               No  No   0   0 -62.9  25.6
apr1036  940413_043354    0 13-APR-94  04:33:54     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1037  940413_044052    0 13-APR-94  04:40:52     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1038  940413_044655    0 13-APR-94  04:46:55     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1039  940413_045336    0 13-APR-94  04:53:36     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1040  940413_050015    0 13-APR-94  05:00:15     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1041  940413_050656    0 13-APR-94  05:06:56     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1042  940413_051406    0 13-APR-94  05:14:06     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1043  940413_052009    0 13-APR-94  05:20:09     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1044  940413_052649    0 13-APR-94  05:26:49     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1045  940413_053327    0 13-APR-94  05:33:27     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1046  940413_054008    0 13-APR-94  05:40:08     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1047  940413_054719    0 13-APR-94  05:47:19     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1048  940413_055320    0 13-APR-94  05:53:20     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1049  940413_060000    0 13-APR-94  06:00:00     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1050  940413_060639    0 13-APR-94  06:06:39     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1051  940413_061319    0 13-APR-94  06:13:19     0.     lyotscan              Yes  No   0   0 -62.9  25.6
apr1052  940413_062032    0 13-APR-94  06:20:32     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1053  940413_062634    0 13-APR-94  06:26:34     0.     lyotscan               No  No   0   0 -65.7  25.6
apr1054  940413_063316    0 13-APR-94  06:33:16     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1055  940413_063954    0 13-APR-94  06:39:54     0.     lyotscan              Yes  No   0   0 -65.7  25.6
apr1056  940413_064634    0 13-APR-94  06:46:34     0.     lyotscan     obsmodev  No  No   0   0 -65.7  25.6
apr1057  940414_030507    0 14-APR-94  03:05:07     0.      unknown              Yes  No   0   0    ??  26.8
apr1058  940415_212331    0 15-APR-94  21:23:31     0.      unknown      ldsosld Yes  No   0   0    ??  25.6
apr1059  940416_003810    0 16-APR-94  00:38:10     0.      unknown     closelog  No  No   0   0    ??  25.6
apr1060  940416_005624    0 16-APR-94  00:56:24     0.      unknown               No  No   0   0    ??  40.7
apr1061  940416_200104    0 16-APR-94  20:01:04     0.      unknown      ldmctap  No  No   0   0  23.3  25.6
apr1062  940416_213358    0 16-APR-94  21:33:58     0.      unknown     lobsmode  No  No   0   0  23.3  25.6
apr1063  940417_021220    0 17-APR-94  02:12:20     0.      unknown       lcmcom  No  No   0   0  23.3  25.8
apr1064  940417_030620    0 17-APR-94  03:06:20     0.      unknown        ipcom  No  No   0   0  23.3  27.3
apr1065  940417_033526    0 17-APR-94  03:35:26     0.        ipcom       mdihoc  No  No   0   0  23.3  26.4
apr1066  940417_070211    0 17-APR-94  07:02:11     0.      unknown               No  No   0   0    ??  25.6
apr1066a 940419_015536    0 19-APR-94  01:55:36     0.         sfft              Yes  No   0   0  23.3  25.6
apr1066b 940419_184956    0 19-APR-94  18:49:56     0.      unknown              Yes  No   0   0  23.3  25.6
apr1066c 940419_191048    0 19-APR-94  19:10:48     0.      unknown              Yes  No   0   0  23.3  25.7
apr1066d 940419_191700    0 19-APR-94  19:17:00     0.      unknown              Yes  No   0   0  23.3  25.7
apr1066e 940419_192358    0 19-APR-94  19:23:58     0.      unknown     closelog Yes  No   0   0  23.3  25.9
apr1066f 940419_212453    0 19-APR-94  21:24:53     0.      unknown       lcmcom Yes  No   0   0  20.0  26.5

apr1066g 940419_221546    0 19-APR-94  22:15:46     0.      unknown       lsqcom Yes  No   0   0  23.3  28.2
apr1066h 940419_231652    0 19-APR-94  23:16:52     0.      unknown              Yes  No   0   0  23.3  26.2
apr1066i 940419_234346    0 19-APR-94  23:43:46     0.      unknown     closelog Yes  No   0   0  23.3  25.8
apr1066j 940420_004906    0 20-APR-94  00:49:06     0.      unknown              Yes  No   0   0    ??  25.6
apr1066k 940420_013337    0 20-APR-94  01:33:37     0.      unknown     iptbltst Yes  No   0   0  23.3  25.6
apr1066l 940420_185642    0 20-APR-94  18:56:42     0.      unknown       lcmcom Yes  No   0   0  20.0  25.6
apr1066m 940420_193742    0 20-APR-94  19:37:42     0.      unknown       lsqcom Yes  No   0   0  23.3  26.8
apr1066n 940420_215142    0 20-APR-94  21:51:42     0.      unknown              Yes  No   0   0    ??  26.8
apr1066o 940421_180426    0 21-APR-94  18:04:26     0.      unknown      openlog Yes  No   0   0  23.3  34.2
apr1066p 940421_182421    0 21-APR-94  18:24:21     0.      unknown     tbltst22 Yes  No   0   0  23.3  35.2
apr1066q 940421_222841    0 21-APR-94  22:28:41     0.      unknown     ldcrpdmp Yes  No   0   0  16.0  35.0
apr1067  940422_003739    0 22-APR-94  00:37:39     0. ldopser05 OB       mdicom Yes  No   0   0   0.0  35.1
apr1067a 940422_024117    0 22-APR-94  02:41:17     0.      unknown       lcmcom Yes  No   0   0  -2.9  34.8
apr1067b 940422_033011    0 22-APR-94  03:30:11     0.      unknown        ipcom Yes  No   0   0   4.0  34.8
apr1067c 940422_035054    0 22-APR-94  03:50:54     0.      unknown     iptbltst Yes  No   0   0   0.0  35.2
apr1067d 940422_060828    0 22-APR-94  06:08:28     0.      unknown              Yes  No   0   0    ??  25.6
apr1068  940422_150100    0 22-APR-94  15:01:00     0.     rundepsq     closelog Yes  No   0   0 -11.4  35.3
apr1069  940422_170509    0 22-APR-94  17:05:09     0.     rundepsq     rundepsq Yes  No   0   0 -11.4  35.2
apr1070  940422_172414    0 22-APR-94  17:24:14     0.     rundepsq              Yes  No   0   0 -11.4  35.2
apr1070a 940422_182514    0 22-APR-94  18:25:14     0.      unknown              Yes  No   0   0 -11.4  35.2
apr1070b 940422_185612    0 22-APR-94  18:56:12     0.      unknown              Yes  No   0   0 -11.4  35.2
apr1070c 940422_195952    0 22-APR-94  19:59:52     0.      unknown     mdifocus Yes  No   0   0 -11.4  35.1
apr1070d 940422_200839    0 22-APR-94  20:08:39     0.      unknown     mdifocus Yes  No   0   0 -11.4  35.2
apr1070e 940422_202556    0 22-APR-94  20:25:56     0.      unknown       mdiltc Yes  No   0   0 -11.4  35.2
apr1070f 940422_203742    0 22-APR-94  20:37:42     0.      unknown      mdipict Yes  No   0   0 -11.4  35.2
apr1071  940422_222720    0 22-APR-94  22:27:20     0.     rundepsq     closelog Yes  No   0   0 -11.4  35.0
apr1072  940422_230000    0 22-APR-94  23:00:00     0. sundopc, bi-      sundopc Yes  No   0   0 -11.4  35.2
apr1072a 940422_235802    0 22-APR-94  23:58:02     0.      sunmagc              Yes  No   0   0 -11.4  35.2
apr1073  940423_000953    0 23-APR-94  00:09:53     0. sunmagc, 1x1              Yes  No   0   0 -11.4  35.2
apr1073a 940423_002048    0 23-APR-94  00:20:48     0.      unknwon     rundepsq Yes  No   0   0 -11.4  35.2
apr1074  940423_002948    0 23-APR-94  00:29:48     0. rundepsq, ld       lsqcom Yes  No   0   0 -11.4  35.2
apr1075  940423_021755    0 23-APR-94  02:17:55     0. mdipict (cos      mdipict Yes  No   0   0 -17.1  35.2
apr1076  940423_023654    0 23-APR-94  02:36:54     0. mdiltc, (CCD       mdiltc Yes  No   0   0 -20.0  35.2
apr1077  940423_025845    0 23-APR-94  02:58:45     0. ccdheat, CCD      ccdheat Yes  No   0   0 -20.0  35.2
apr1078  940423_032027    0 23-APR-94  03:20:27     0. Dark Doppler              Yes  No   0   0 -20.0  35.2
apr1078a 940423_043634    0 23-APR-94  04:36:34     0.      unknown      mdipict Yes  No   0   0 -23.3  35.2
apr1079  940423_045403    0 23-APR-94  04:54:03     0. dark noise m     closelog Yes  No   0   0 -23.3  35.2
apr1080  940423_052704    0 23-APR-94  05:27:04     0. legff (flats        legff Yes  No   0   0 -23.3  35.2
apr1081  940423_065147    0 23-APR-94  06:51:47     0. legff (flats        sleep Yes  No   0   0 -23.3  35.2
apr1081a 940423_223815    0 23-APR-94  22:38:15     0.      unknown     iptbltst Yes  No   0   0  20.0  35.1
apr1081b 940424_001530    0 24-APR-94  00:15:30     0.      unknown     ldcrpdmp Yes  No   0   0  23.3  35.1
apr1082  940424_062607    0 24-APR-94  06:26:07     0. lyotscan; la     calmlasr Yes  No   0   0  -5.7  35.2
apr1083  940424_063208    0 24-APR-94  06:32:08     0. lyotscan, 12              Yes  No   0   0  -8.6  35.2
apr1084  940424_063857    0 24-APR-94  06:38:57     0. lyotscan, 3               Yes  No   0   0 -11.4  35.2
apr1085  940424_064544    0 24-APR-94  06:45:44     0.    6 o'clock              Yes  No   0   0 -11.4  35.2
apr1086  940424_065233    0 24-APR-94  06:52:33     0. lyotscan, 9               Yes  No   0   0 -14.3  35.2
apr1087  940424_070204    0 24-APR-94  07:02:04     0. lyotscan 1/2              Yes  No   0   0 -14.3  35.1
apr1088  940424_070710    0 24-APR-94  07:07:10     0. lyotscan 1/2              Yes  No   0   0 -14.3  35.2
apr1089  940424_071237    0 24-APR-94  07:12:37     0. lyotscan 1/2              Yes  No   0   0 -17.1  35.2
apr1090  940424_071802    0 24-APR-94  07:18:02     0. lyotscan 1/2              Yes  No   0   0 -17.1  35.2
apr1091  940424_072328    0 24-APR-94  07:23:28     0. lyotscan 1/2     obsmodev Yes  No   0   0 -17.1  35.2
apr1092  940424_190751    0 24-APR-94  19:07:51     0.     CCD Cold              Yes  No   0   0 -54.3  25.6
apr1093  940424_193854    0 24-APR-94  19:38:54     0.    CCD Cold2       mdiltc Yes  No   0   0 -60.0  25.6
apr1093a 940424_202445    0 24-APR-94  20:24:45     0.      unknown       mdiltc Yes  No   0   0 -68.6  25.6
apr1094  940424_203144    0 24-APR-94  20:31:44     0.       mdiltc       mdiltc Yes  No   0   0 -74.3  25.6
apr1095  940424_204659    0 24-APR-94  20:46:59     0.      mdipict      mdipict Yes  No   0   0 -77.1  25.6
apr1096  940424_211728    0 24-APR-94  21:17:28     0.    ccdwarmup       lhopon Yes  No   0   0 -30.0  25.6
apr1097  940425_054033    0 25-APR-94  05:40:33     0. lyotscan, OV              Yes  No   0   0 -20.0  25.6
apr1098  940425_054637    0 25-APR-94  05:46:37     0. lyotscan, 12              Yes  No   0   0 -20.0  25.6
apr1099  940425_055325    0 25-APR-94  05:53:25     0. lyotscan, 3               Yes  No   0   0 -20.0  25.6

apr1100  940425_060011    0 25-APR-94  06:00:11     0. lyotscan, 6               Yes  No   0   0 -23.3  25.6
apr1101  940425_060657    0 25-APR-94  06:06:57     0. lyotscan, 9               Yes  No   0   0 -23.3  25.6
apr1102  940425_061438    0 25-APR-94  06:14:38     0. lyotscan 1/2              Yes  No   0   0 -23.3  25.6
apr1103  940425_061943    0 25-APR-94  06:19:43     0. lyotscan 1/2              Yes  No   0   0 -23.3  25.6
apr1104  940425_062511    0 25-APR-94  06:25:11     0. lyotscan 1/2              Yes  No   0   0 -23.3  25.6
apr1105  940425_063037    0 25-APR-94  06:30:37     0. lyotscan 1/2              Yes  No   0   0 -20.0  25.6
apr1106  940425_063604    0 25-APR-94  06:36:04     0. lyotscan 1/2              Yes  No   0   0 -23.3  25.6
apr1107  940425_064553    0 25-APR-94  06:45:53     0. Darks, (soft     obsmlasr Yes  No   0   0 -23.3  25.6
apr1108  940425_065349    0 25-APR-94  06:53:49     0. Darks, (hard     lfdclose Yes  No   0   0 -23.3  25.6
apr1109  940425_164209    0 25-APR-94  16:42:09     0.      unknown       lcmcom Yes  No   0   0  20.0  25.6
apr1110  940425_173206    0 25-APR-94  17:32:06     0.      unknown       lsqcom Yes  No   0   0  30.0  25.6
apr1111  940425_192438    0 25-APR-94  19:24:38     0.      unknown     mdilaser Yes  No   0   0  23.3  25.6
apr1112  940425_213204    0 25-APR-94  21:32:04     0.      unknown       mdiiss Yes  No   0   0  26.7  25.6
apr1113  940426_035538    0 26-APR-94  03:55:38     0.      unknown      mdipict Yes  No   0   0  23.3  25.6
apr1114  940426_041407    0 26-APR-94  04:14:07     0.      unknown       mdiltc Yes  No   0   0  23.3  25.6
apr1115  940426_043043    0 26-APR-94  04:30:43     0.      unknown     mdifocus Yes  No   0   0  23.3  25.6
apr1116  940426_044621    0 26-APR-94  04:46:21     0.      unknown      mdipict Yes  No   0   0  23.3  25.6
apr1117  940426_050146    0 26-APR-94  05:01:46     0.      unknown      mdimrot Yes  No   0   0  23.3  25.6
apr1118  940426_060248    0 26-APR-94  06:02:48     0.      unknown      mdiprot Yes  No   0   0  23.3  25.6
apr1119  940426_061816    0 26-APR-94  06:18:16     0.      unknown       lsqcom Yes  No   0   0  23.3  25.6

Index

command history 36

data product codes 44 DPC 44 dust map 19

filenames 1 FITS fields 54

history, command 36 HK Mnemonics 45 HKPLOT 31 house keeping 45

IMG SUMMARY 9

LFFT routines 12

MDIDUST 19 MDILASER 17 MK FITS HEAD 10 mnemonics 45

printers 9

READ MDI 8 RFITS2 10

series numbers 1 SET PRINTER 9 SHOW DATA 6 SXADDPAR 10 SXPAR 10

tags, FITS 54 telemetry filenames 1 TRAP CHECK 29

UTPLOT 8

VIEW IMAGE 5